### v1

# Naturalistic Computational Cognitive Science Towards generalizable models and theories that capture the full range of natural behavior

Abstract
Artificial Intelligence increasingly pursues large, complex models that perform many tasks within increasingly realistic domains. How, if at all, should these developments in AI influence cognitive science? We argue that progress in AI offers timely opportunities for cognitive science to embrace experiments with increasingly naturalistic stimuli, tasks, and behaviors; and computational models that can accommodate these changes. We first review a growing body of research spanning neuroscience, cognitive science, and AI that suggests that incorporating a broader range of naturalistic experimental paradigms (and models that accommodate them) may be necessary to resolve some aspects of natural intelligence and ensure that our theories generalize. We then suggest that integrating recent progress in AI and cognitive science will enable us to engage with more naturalistic phenomena without giving up experimental control or the pursuit of theoretically grounded understanding. We offer practical guidance on how methodological practices can contribute to cumulative progress in naturalistic computational cognitive science, and illustrate a path towards building computational models that solve the real problems of natural cognition—together with a reductive understanding of the processes and principles by which they do so.
1 Introduction
Cognitive scientists build models to make our theories concrete — which offers testable predictions, eliminates ambiguities (Guest and Martin,, 2021), and can reveal unexpected properties of cognition and behavior (McClelland,, 2009). Cognitive models can range from simplified models of a high-level process (e.g. Frank and Goodman,, 2012; Cohen et al.,, 1990), to task-performing models of a particular domain (Newell,, 1973), or beyond (Newell,, 1994). Many cognitive paradigms—connectionism (McClelland et al.,, 1986), bayesian inference (Tenenbaum and Griffiths,, 2001), and cognitive architectures (e.g. Ritter et al.,, 2019; Laird,, 2019)—have sought generalizable models that can explain a broad range of behavior from a simpler set of principles. Ultimately, these approaches could aspire to build towards ”unified theories of cognition” that can explain and predict natural behavior across many tasks and domains (Newell,, 1994). However, we are not there yet. In recent years there have been substantial debates about the generalizability of our theories (Yarkoni,, 2022; Eckstein et al.,, 2021), whether our models and theories are adequately constrained (e.g. Jones and Love,, 2011; Rahnev and Denison,, 2018), and calls to accommodate a broader scope of naturalistic phenomena (Wise et al.,, 2023; Cisek and Green,, 2024). Our goal in this paper is to motivate and outline a path that we believe will begin to address many of these challenges.
The path we propose is motivated by recent practical and conceptual developments in AI. In the last ten years AI has progressed substantially in building models that can perform a wide range of tasks in various domains. We now have vision models that can classify and segment real-world images in objects without external supervision (He et al.,, 2020; Caron et al.,, 2021), we have reinforcement learning models that can learn new tasks in a domain with human-like efficiency (Team et al.,, 2023), and we have language models which can produce human-like language and solve many language-specified tasks (Radford et al.,, 2019; Wei et al.,, 2022). Surprisingly, such models produce internal representations that capture many features of brain representations (Yamins et al.,, 2014; Schrimpf et al.,, 2021). These findings have led to discussion of how such complex models can be explanatory (Cao and Yamins, 2024a, ), and how they should alter our theories (Hasson et al.,, 2020; Perconti and Plebe,, 2020; Piantadosi,, 2023). Yet they have also led to debate over how we test our models and interpret the results (Bowers et al.,, 2023; DiCarlo et al.,, 2023; Dentella et al.,, 2024). Thus, it remains an open question if and how the progress in AI can contribute to addressing the challenges of cognitive science.
Figure 1: Naturalistic computational cognitive science: the what, why, and the how. The first section (§2) provides an overview of “naturalistic computational cognitive science”. Afterwards (§34), we motivate the utility of naturalistic experimental paradigms and learning-based approached for developing a more complete understanding of cognition. The rest of the paper (§56) focuses on how to achieve these goals; how to develop models for naturalistic settings, and how to use naturalistic experiments and models as part of explanatory cognitive theories. (Figures reproduced from Saxe et al.,, 2019; Geirhos et al.,, 2018; Doshi and Konkle,, 2023.)
In this work, we attempt to weave a thread of arguments that synthesizes these literatures, spanning from cognitive science’s motivation and theory development to the practicalities of model engineering and reproducibility. We articulate a research strategy that we believe will be helpful in enabling theory-driven cognitive science to achieve deeper understanding of the full range of natural intelligence. We present an overview of this framework and our arguments in Figure 1.
Our paper proceeds as follows. First, in §2, we unpack what we mean by “naturalistic” computational cognitive science, by providing a pragmatic outline of the features that we believe deserve more emphasis in our experimental paradigms and models. In particular, we argue that researchers should expand their task paradigms to to more closely approximate the breadth of the settings in which their theoretical constructs would be expected to generalize by both adding relevant variables that may interact with those in question, and by incorporating more variability within existing parameters. Likewise, researchers should expand their models to more faithfully approximate the computational task that the natural system is solving.
In the next two sections, we motivate these goals, by arguing that increasing naturalism is necessary for developing generalizable understanding. First, in §3, we review strands of evidence spanning neuroscience, cognitive science, and AI, showing that naturalistic experimental paradigms can lead to different behavior, engage brain systems differently, and expose expose computational challenges that engage mechanisms differently. As an example, Francis et al., (2017) found that moral judgments on a trolley problem were dramatically different in a virtual reality presentation compared to text vignettes. Other works that we review have found that processes ranging from visual processing (Amme et al.,, 2024), learning (Rosenberg et al.,, 2021; Collins,, 2024), memory (Helbing et al.,, 2020) and more are changed when naturalistic elements are introduced. The works that we review in this section thus suggest that experiments that span a broader range of naturalistic settings are essential for drawing correct inferences about the system as a whole. Next, in §4, we highlight the benefits of models that learn from naturalistic data—which can perform a wide range of tasks, and yield qualitatively different patterns of generalization than models trained in simplified settings. In turn, these differences can change the inferences we make from our models, and yield new insights about the origins of cognitive and neural phenomena. Together, these two sections of our paper argue that pursuing naturalistic experiments, and models trained on them, will be necessary for achieving complete understanding of cognition.
In the remainder of the paper, we turn to the practice of naturalistic computational cognitive science, and its contributions to building generalizable cognitive theories. First, in §5, we review key lessons computational cognitive scientists can adopt from AI in developing generalizable models that can operate in increasingly naturalistic data conditions: the benefits of frictionlessly reproducing prior research artifacts, the importance of generalizability-focused model development, and the value of benchmarking for concentrating research effort on important questions. Finally, in §6, we outline how one can leverage these potentially complicated and opaque models to contribute to generalizable theories of cognition—uniting task-performing predictive models with explanatory reductive understanding, using established techniques like rational analysis (Anderson,, 1990) and new approaches to interpreting complex models (Geiger et al.,, 2021). We close in §7 by discussing the broader context and implications of naturalistic computational cognitive science.
Figure 2: Overview of naturalistic computational cognitive science. Given a novel computational model, researchers generate predictions to test their model in a simplified setting. If this model successfully predicts human behavior, researchers see if their model continues to generalize human predictions in increasingly naturalisic conditions.
2 What is “naturalistic” computational cognitive science?
Naturalistic computational cognitive science is a research strategy for theory-driven cognitive science that aims to predict human behavior across increasingly naturalistic stimuli and tasks. This approach gradually increases the ecological validity of experimental designs while ensuring models can accommodate both simplified and more natural conditions. As demonstrated in machine learning (§5), this strategy has enabled models that can work with real-world data, albeit imperfectly.
As in other approaches within computational cognitive science, this approach starts with formulating models and designing controlled experiments to isolate core phenomena. Like prior work, we place an emphasis on model generalization, i.e. predicting human behavior on data the model hasn’t been exposed to—including data from new experimental conditions or tasks (Busemeyer and Wang,, 2000). Therefore,
  1. 1.
    Model parameters are estimated with a “training” distribution before being evaluated.
  2. 2.
    Models should predict human behavior on held-out examples that neither the model nor humans have encountered.
One key facet of naturalistic computational cognitive science is that we place emphasis on having models—in particular, the same model—predict human behavior across increasingly naturalistic conditions that cover a broader space of settings in which the theoretical construct would be expected to generalize. We see two main paths toward increasing naturalism in experimental paradigms:
  1. 1.
    Changing existing experiment parameters to better cover the scope of natural distributions. For example, in an object classification task, we might expand the category set to include the broad range of objects people encounter in daily their life.
  2. 2.
    Adding ecologically motivated parameters that might interact with the theoretical constructs in question. For example, in a linguistic judgment task, we might test judgments across both spoken and written utterances.
We provides examples in Figure 3.
Figure 3: Examples of increasingly naturalistic settings that we can now study in a theory-driven manner. All of these are settings where tasks and stimuli can now be parametrically generated—i.e., thanks to “generative AI”, we can now automate the generation of photorealistic synthetic data and virtual worlds; thanks to virtual and augmented reality, we can now scan and parametrically manipulate real environments. Thanks to progress in AI, we can now build task-performing models that can operate in these stimuli and task settings.
Our goal is to encourage cognitive scientists to embrace experimental paradigms and models that capture a broader spectrum of the variability in inputs, interactions, and tasks present in the natural environment of humans or animals—including, for humans, the rich cultural constructs that form so much of our modern experience. That is, we identify a paradigm as more naturalistic if it incorporates a broader scope of the range of natural variability over which the theoretical constructs would be expected to generalize. Correspondingly, we identify a model as more naturalistic if it is capable of generalizing to make predictions across a broader range of contexts. Note that increasing naturalism does not always mean faithfully approximating all the joint statistics of the natural distribution—increasing the scope of variation along naturalistic dimensions can also enable moving outside the natural data distribution for controlled experiments (see §6.1).
Generally, identifying the parameters for increasing naturalism in a particular experiment is not well defined or obvious—it is task- and theory-specific. To illustrate “naturalistic” more explicitly, we will therefore give some examples of dimensions along which naturalism can be increased.
Task paradigm
  • Incorporating multiple paradigms across which a hypothesis would be expected to hold, rather than testing on a single task; for example testing multi-step RL tasks as well as bandit tasks.
  • Incorporating a broader set of stimuli, for example augmenting synthetic images with real images, or images with videos.
  • Having task stimuli generated by many varying latent factors, not just the ostensible variables of theoretical interest.
Environment
  • Large state spaces that reflect the complexity of real-world environments (Wise et al.,, 2023).
  • Including the presence of other social agents that continue to learn with the agent (Hu et al.,, 1998) and that the organism must interact with (Vélez and Gweon,, 2021).
  • A continually changing environment (Abel et al.,, 2024)
Model architecture
  • Architectures that can operate over sensory inputs like natural images (Yamins et al.,, 2014), speech (Kell et al.,, 2018), or natural language (Schrimpf et al.,, 2021) rather than simplified stimuli (e.g. low-dimensional or discrete inputs that contain only task-relevant features).
  • Architectures with naturalistic action spaces—for example, modeling low-level motor control and embodiment rather than abstract, symbolic actions (Merel et al.,, 2019).
Learning algorithm
  • Unsupervised (Higgins et al.,, 2021), self-supervised (Konkle and Alvarez,, 2022), or intrinsic (Chentanez et al.,, 2004; Oudeyer et al.,, 2013) learning algorithms.
  • Social (Henrich,, 2016) or cultural (Cook et al.,, 2024) learning algorithms.
  • Prospective learning algorithms that aim to model how the tasks evolve in continually changing environments (Seligman et al.,, 2013; De Silva et al.,, 2023).
Of course, it is easy to say that increasing the naturalism of our experiments and models would be useful, but how can we actually achieve it? A key goal for the later sections of our paper (§5-6) will be to give practical perspectives on how to achieve these goals, and how to develop theories that generalize across a broader range of naturalistic behavior. We briefly sketch these perspectives here.
How do we scale up to naturalistic experiments? We believe that recent developments in computer science and engineering make increasing naturalism in our experimental paradigms much more accessible to researchers, as we can use natural language to programmatically generate stimuli such as 3D scenes (Zhou et al.,, 2024) and videos of objects (Villegas et al.,, 2022), and leverage haptic feedback in virtual reality to simulate textures (Lu et al., 2022b, ). While the specific parameters will vary by domain, the core principle remains: gradually increasing ecological validity while maintaining experimental control.
How do we glean understanding from potentially opaque learning based models? (§6) We believe that task performing models offer powerful new tools for building cognitive theories. In a sense, these tools amount to treating the cognitive model as a proxy object for study—but one that admits much more opportunity for experimentation and understanding. For example, we can use “neuroscience”-like causal methods to probe a model’s computations by intervening on them (e.g. Geiger et al.,, 2021). We can also interpret its behavior as an adaptation to the naturalistic properties of its training data (Chan et al.,, 2022; Prystawski et al.,, 2024, e.g.)—as a kind of rational analysis (Anderson,, 1990; Lewis et al.,, 2014) of behavior and computations being a rational response to an environment. When applying these approaches to computational models, we have the ability to explicitly test our assumptions in ways we cannot with human subjects (e.g., how would the system behave if it were trained on only ungrammatical sentences). We can also attempt to distill our insights further, into analytic abstractions that illustrate the minimal instantiation of a phenomenon. By combining these routes, we can build rigorous links from an opaque task-performing model to conceptual understanding of cognition.
In summary naturalistic computational cognitive science seeks to (1) explain real-world intelligent behavior by developing models that operate over the scope of naturalistic inputs, produce naturalistic outputs, and are optimized according to the actual constraints and affordances of an person’s environment, and (2) develop cognitive theories that unite these models with reductive understanding. In the coming sections, we will first motivate more deeply how naturalistic experimental settings can help us develop more complete accounts of intelligence, before returning in more detail to the questions of how these goals can be achieved.
3 The benefits of naturalistic experimental settings
Figure 4: Overview of the benefits of increasingly naturalisic experimental conditions.
In this section, we outline the benefits of naturalistic experimental settings for cognitive science. Specifically, we argue that naturalistic experimental settings can engage mechanisms that are qualitatively different than those engaged in simpler settings — and thereby change the scientific inferences we draw. We review examples where naturalistic experimental settings have changed our understanding of a system, and have helped to disentangle competing models that simpler settings could not.
As an orienting conceptual example, imagine that we are doctors studying heart function, and we chose to only study it while the participants are lying down and resting. We could certainly learn interesting things thereby; the resting setting offers high test-retest reliability, and a highly-controlled environment to study the processes like how respiratory cycles affect heart rate. However, it offers this control and reliability precisely by removing the many factors of natural variation that interact with heart rate, and indeed for which heart function evolved — such as rapid adaptation to movement, stress, etc. Studying heart function across a broader range of naturalistic settings would be necessary to understand why the system is the way it is, and how all the physiological (and psychological) processes involved interact. For example, if we experimentally manipulated breathing, but did so across a range of different naturalistic activities (sitting, speaking, running, etc.) we would more effectively determine how breathing affects heart function, and how that effect depends on the state of the overall system. That is, by studying phenomena across a range of naturalistic settings we elaborate our understanding of the system, and can build more complete theories of its function.
3.1 Behavior can be different in naturalistic experiments
The starting point for many cognitive analyses is behaviors on a task designed to isolate some cognitive capability. However, there are many cases in which task-relevant behavior can be altered by seemingly-orthogonal features of the task—thus raising the risk that our theories will be overfit to a restricted task setting unless we also consider a broading range of variation. In this section, we review some examples illustrating changes in behavior or capabilities in naturalistic settings.
Example 1: moral behavior can change when moving from textual vignettes to virtual reality. Recent research has shown that behavioral choices can change when moving from text to virtual reality as it can mark a shift from moral judgment (i.e. deciding whether things are moral) to moral action (Francis et al.,, 2017). When presented with the classic trolley dilemma in text form, only 40% of participants say they would pull a switch to divert a trolley that would kill one person to save five others. However, when the same scenario is presented in virtual reality where participants must simulate the action, 55% choose to pull the switch. This difference becomes even more pronounced in the “footbridge” variant of the dilemma. In the text version, only 10% of participants say they would push someone off a bridge to save five others. Yet when able to simulate this action in virtual reality, 65% of participants choose to push the person. These findings demonstrate that people’s stated moral preferences in hypothetical scenarios can differ substantially from their actions in more naturalistic settings.
Example 2: memory performance differs between naturalistic search and explicit memorization. When participants are explicitly instructed to memorize objects in a 3D home environment, their subsequent recall accuracy was significantly lower than when they incidentally encounter the same objects during visual search tasks (Helbing et al.,, 2020). This effect extends to spatial memory as well. Not only do participants better remember the identity of objects encountered during search, they also show more accurate memory for object locations—placing objects closer to their original positions when they were initially seen during search versus explicit memorization. These findings demonstrate that tasks focused explicitly on memory may not capture all aspects of how memory naturally operates during everyday behavior. Additionally, it suggests that computational models of memory may need to be revised to account for these differences in naturalistic settings. This is supported by research showing that people display detailed memory of object and spatial information in real-world scenes (Bainbridge et al.,, 2019).
Example 3: mice can learn $1000\times$1000 × faster during natural behavior, compared to two-alternative forced choice tasks. Recent research has shown that learning rates can dramatically differ between simplified laboratory tasks and more naturalistic settings (Rosenberg et al.,, 2021). When mice are tested in traditional two-alternative forced choice (2AFC) tasks, they typically require 10,000 trials over 3-6 weeks of training to reach asymptotic performance of only about 67% accuracy. However, when allowed to freely explore a labyrinth environment (similar to a natural burrow systems (Small,, 1901)), mice demonstrate remarkably accelerated learning (Rosenberg et al.,, 2021). In this setting, mice learned to navigate a complex maze with 63 T-junctions in just a few hours, discovering and optimizing their path to a water reward after only about 10 reward experiences. The learning rate in the naturalistic setting was approximately $1000\times$1000 × times faster than in traditional 2AFC paradigms. This dramatic difference in learning speed demonstrates how traditional simplified experimental paradigms may fundamentally underestimate an organism’s learning capabilities, and suggests that natural learning mechanisms may operate differently when they are engaged through more naturalistic behavior.
3.2 Neural systems can operate differently under naturalistic conditions
In addition to behavioral differences like those reviewed above, naturalistic conditions can result in changes to neural coding and computation. Thus, in order to arrive at a complete understanding of neural function, we need to consider the responses of the system across a range of simpler and more naturalistic settings. In this section, we highlight some examples of different computations or computational roles that arise when moving from standard paradigms to more naturalistic ones.
Example 1: the model of superior colliculus shifted from a model of eye movement to one of integrating sensory inputs for body control. Cisek, (1999) suggests that the focus on computational behavior as an input-output mapping neglects the fundamental fact that natural intelligence evolved not to produce single responses, but for closed-loop control in an environment. This setting yields a rather different interpretation of the system’s representations and processes. More generally, many researchers have argued that cognition cannot be understood completely outside its embodiment and the environment in which cognitive processes are instantiated (e.g., Newen et al.,, 2018). For example, Cisek and Green, (2024) argues that as we increased naturalism from head-fixed to head-free to body-free settings when studying monkeys, we expanded our theory of superior colliculus from controlling saccadic eye movement to generally integrating multimodal cues to guide bodily action-selection (Cisek and Green,, 2024). That is, as we increased the naturalism of the experimental conditions that we used to study monkeys, we arrived at a more complete model of superior colliculus.
Example 2: visual processing systems operate differently during natural viewing compared to passive viewing paradigms. Traditional research has shown that one of the earliest visual responses in the brain (known as P100/M100) occurs about 100ms after a person’s eyes land and fixate on a new location (Vinje and Gallant,, 2000). However, this finding comes primarily from experiments where participants passively view images while keeping their eyes still. In contrast, natural vision involves actively moving our eyes multiple times per second to sample information from different parts of a scene. When the researchers examined brain responses during this more natural active viewing, they found that the P100/M100 response actually begins when the eye movement (saccade) starts, not when it ends at the new fixation point (Amme et al.,, 2024). Furthermore, they discovered that the neural patterns during active viewing were opposite (anti-correlated) to those seen during passive viewing, suggesting fundamentally different processing mechanisms. These findings reveal that simplified experimental paradigms, while valuable, may not capture all aspects of how visual processing actually operates during natural behavior.
Example 3: model-based learning systems are used for learning moral judgments, but model-free learning are used for learning to avoid harming others. Despite moral reasoning being strongly associated with model-based learning systems in the prefrontal cortex, neural evidence shows a surprising shift when people actually learn to avoid harming others (Lockwood et al.,, 2020). Moral reasoning commonly engages model-based systems in the lateral prefrontal cortex (LFPC) (Spitzer et al.,, 2007; Carlson and Crockett,, 2018) with LPFC disruption leading to reduced norm compliance and enforcement (Knoch et al.,, 2006; Ruff et al.,, 2013). However, when participants engage in tasks requiring them to learn moral behavior (rather than just reason about it), researchers instead observed neural activity consistent with model-free learning. Similar to our moral judgment vs. moral action example from §3.1, this shows that when an experiment becomes more naturalistic (not just reasoning about morality but learning moral behavior), neural systems can operate in unexpected ways.
3.3 Naturalistic experimental paradigms can expose computational challenges that engage mechanisms differently
Finally, naturalistic paradigms can likewise introduce challenges that are important to understanding functioning of the system as a whole. In this section, we highlight examples where the challenges posed by increasing naturalism can engage computational mechanisms differently—which can be useful for disentangling underlying mechanisms, and is important for achieving generalizable understanding.
Example 1: Different learning mechanisms only yield distinct predictions under working memory load. Overly-simplified tasks can yield aliasing of different solutions, where many different computational approaches yield the same behavior. Collins, (2024) presents an example, by showing that what seems superficially to be implemented as a standard reinforcement learning algorithm, on closer examination may instead be implemented through a combination of working memory and simpler value-free associative learning. [^1] Whether this alternative counts as an alternative implementation of RL or not is orthogonal to our argument; the important point is that the two models yield distinct predictions in some regimes but not others. However, the difference between these two implementations cannot be identified in standard task settings, as Collins notes: “Even simple tasks designed to elicit a target process (such as bandit tasks for RL) recruit multiple other processes, but those processes may be unidentifiable in such tasks. Disentangling multiple processes requires considering more complex tasks to elicit differentiable behavior.” The more complex tasks in question simply increase the stimulus set size within learning blocks to a handful of objects, rather than restricting to two or three—i.e., more broadly covering the range of set sizes more clearly disentangles the learning processes. Thus, by restricting to minimal paradigms, we enable a system to use many solutions and prevent ourselves from discriminating between them. By contrast, if we impose a broad range of evaluations on the system (cf. Nau et al.,, 2024)—and especially naturalistic evaluations that increase demands along different axes of task difficulty—the increased constraints actually make it easier to map the model on to natural intelligence, as is highlighted in the contravariance principle of Cao and Yamins, 2024b .
Example 2: fear conditioning. Traditional fear conditioning research has provided fundamental insights into how humans learn associations between stimuli and aversive outcomes (Sehlmeyer et al.,, 2009; Maren,, 2001; Britton et al.,, 2014). Through carefully controlled laboratory studies involving paired stimuli and outcomes, we have developed a rich mechanistic understanding of basic fear learning processes. These paradigms have proven especially valuable for studying certain types of naturally occurring fear learning, such as immediate aversive responses to foods (Riley and Tuck,, 1985) or traumatic events like combat exposure (Engelhard et al.,, 2008). However, real-world fear learning often involves additional temporal dynamics that complement these well-studied processes. For example, a student may develop fear of exams after receiving a single poor grade weeks after taking the test, without requiring repeated negative experiences (Mobbs et al.,, 2021). This type of naturalistic learning involves delayed delayed and sparse feedback for choices, i.e. it relies on a computational mechanism for credit assignment (Sutton and Barto,, 2018). Credit assignment describes how an agent credits choices (e.g. studying, socializing, diet, etc.) as responsible for future events (e.g. a poor grade). This is not to say humans have optimal credit assignment but that they have some mechanism for it and we are empowered to study it with appropriately naturalistic experimental settings. Clinical researchers have additionally pointed out that increased ecological validity can increase the applicability of our experimental results as the gap between lab-based fear conditioning and real-world fear has been a barrier to successful treatment of pathological fears (Beckers et al.,, 2013; Scheveneels et al.,, 2016; Krypotos et al.,, 2018).
Example 3: seemingly-similar tasks that engage brain mechanisms differently can be reconciled by naturalistic models. There has been a long debate about whether perceptual processes like “oddity” tasks (identifying an object that is different from the rest of the objects in a set) recruit the Medial Temporal Lobe (MTL). Specifically, some researchers (e.g Murray et al.,, 2007) have suggested that MTL is recruited to discriminate among “complex” stimulus sets that visual cortex cannot discriminate on its own (and thus MTL lesions impair performance specifically on those stimuli); other researchers have found no such interaction and suggested it stems from methodological issues (Buffalo et al.,, 1998; Suzuki,, 2009). Recently, stimulus-computable models have shed new light on this debate. Specifically, Bonnen et al., (2021); Bonnen and Eldridge, (2023) used pretrained convolutional vision models to directly process all the raw stimuli used in prior experiments, and evaluated whether the “oddity” stimulus could be distinguished from the rest of the set in the representation space of these models. By doing so, Bonnen et al. showed that some of the purportedly “complex” stimuli from some of the prior experiments could nevertheless be distinguished with the vision model alone, while the complex stimuli from other experiments could not. Furthermore, the purportedly “complex” stimuli that in fact could be discriminated by the pretrained vision network alone did not require the MTL, while the complex stimuli that could not be distinguished by the vision model were precisely the ones that required MTL. Thus, the way that prior works operationalized “complexity” was underspecified; differences in operationalization that got abstracted away in conceptual descriptions of the experiments likely drove the seemingly-conflicting findings. This example illustrates the value of building models that directly perform the same naturalistic tasks as the subjects, and testing these models on benchmarks that incorporate stimuli from many experimental paradigms. This naturalistic approach can elucidate important differences among experimental paradigms, and clarify conceptual understanding.
4 The (surprising) benefits of learning with naturalistic data
Figure 5: Overview of the benefits of learning with naturalistic data. (Figures reproduced from Doshi and Konkle,, 2023; Manning et al.,, 2020; Team et al.,, 2023.)
The standard lesson in science is that we must simplify an experimental setting so that we can better arrive at a causal conclusion. However, we have already seen examples illustrating that more is different (Anderson,, 1972)—certain properties of intelligence may only emerge in more “complicated” naturalistic settings. In this section, we detail examples where naturalistic data itself seems to play an importance role in producing the empirical phenomena of intelligence that we wish to study.
Many of the findings we will discuss here will be modern results from AI. One common trend is that there is a positive interaction between learning-based systems and naturalistic data—learning-based systems can accommodate naturalistic data, and reciprocally, learning from naturalistic data results in qualitatively better results than learning in simpler settings. While we do not yet know the implications for natural intelligence, they open interesting questions and challenge prior assumptions. In particular, we argue that cognitive science should consider the limitations placed on our models when we fail to consider naturalistic data. (Note that while we draw many case studies from particular areas of deep learning, we believe that some of these factors could likewise contribute to generalization under other learning paradigms, such as program synthesis, or even evolution—and thus should be of interest even to those who do not work with deep learning models.)
4.1 Learning enables models that can operate on natural data
Perhaps the first surprising benefit of (deep) learning is that it enables a system to operate over natural data. Until AlexNet (Krizhevsky et al.,, 2012), the prevailing wisdom was that one had to design useful representations of natural data for models. However, AlexNet and subsequent work showed that a model could learn useful representations for naturalistic data simply by being trained to make predictions about this data. Surprisingly, these models learn representations that have notable correspondence to the internal representations in visual cortex (Yamins et al.,, 2014; Khaligh-Razavi and Kriegeskorte,, 2014). Here, we review some of this work and how learning-based methods can enable more flexible cognitive models that operate on natural data.
Learning-based vision models can operate on natural data and predict neural data (surprisingly) well. There is a rich history of learning-based methods revealing how neural networks can develop internal representations from pixel-based images that strongly correspond to neural data. A seminal example comes from Olshausen and Field, (1996), who showed that V1-like Gabor filters could emerge from a convolutional neural network optimized to transmit input while minimizing overall activation (essentially, a sparse autoencoder). Prior to this, V1 activations were modeled using analytically defined Gabor filters. This breakthrough demonstrated that V1-like receptive fields could arise naturally from optimization to environmental statistics of edges at varying frequencies and orientations. Building on this foundation, Lee et al., (2007) demonstrated that V2-like convolutional filters could emerge from unsupervised optimization of sparse deep belief networks. A significant advance came when Yamins and DiCarlo, (2016) showed that task-optimized deep convolutional neural networks trained for object classification developed intermediate layers that matched neural representations in V4 and IT better than any previous models. Recent research has revealed an even more striking finding: these task-optimized networks represent images in a latent space whose geometry has remarkable correspondence to brain topography (Doshi and Konkle,, 2023). When fitting a 2D sheet to image representations such that neighboring points correspond to nearby points in the original high-dimensional space, the resulting sheet shows striking similarities to actual brain topography. Thus, learning-based approaches, particularly neural networks, not only enable effective visual processing but also provide surprisingly accurate predictions of brain organization and function.
Integrating learning-based models with program synthesis enables flexible Bayesian models of human concept learning. Recent work has shown how large language models can be combined with program synthesis to create more generalizable Bayesian models of human reasoning (Ellis,, 2024). Hand-designed symbolic hypothesis spaces don’t always transfer across domains.  Ellis, (2024) showed that one can circumvent these challenges by leveraging large language models to enable use of natural language as a flexible hypothesis space with a data-driven prior tuned to human judgments. By translating natural language hypotheses to executable programs for likelihood computation, such models can capture human learning dynamics in online concept acquisition tasks while continuing to predict human behavior robustly in new domains. This overcomes traditional limitations of pure symbolic program synthesis approaches which can fail to transfer due to incomplete symbolic vocabularies—a manifestation of the frame problem (Shanahan,, 2004), where when deciding the relevant variables for a problem, we might accidentally choose a framing that doesn’t generalize to other settings of interest. The integration of learning-based components with program synthesis thus enables Bayesian models that can flexibly adapt their hypothesis space while maintaining interpretable symbolic computation—a goal of many cognitive modellers.
4.2 Learning from naturalistic data can improve generalization
Studying how systems generalize is a key method in cognitive science and AI. Conceptual arguments often rely on empirical studies of generalization; however, these generalization properties are often studied in simple settings, with minimal models learning from minimal features (e.g. Marcus,, 1998). A fundamental assumption underlying this approach is that the simplifications do not alter the generalization problem. However, here we review studies from AI and neuroscience showing that the variability of experience according to seemingly-orthogonal variables can alter how systems learn and generalize—thus implying that studying models within more restricted settings may be misleading about the naturalistic computational problem.
Example 1: Compositional generalization of image classifiers vs. agents. Naturalistic tasks can fundamentally change what models learn and how they generalize. Hill et al., (2020) trained two models to do a vision-language grounding task, and tested their compositional generalization to held-out language instructions. Both models were trained on the same language examples, and tested on the same held-out examples. However, one model was trained as an agent that interacts with a simulated environment, while the other was a simple image-language classifier trained on screenshots from that environment. The authors found that the interacting agent exhibited perfect compositional generalization to the novel language utterances, while the image classifier was substantially worse in novel settings. The authors also explored a range of other settings, including generalization benefits of more naturalistic 3D (rather than 2D) environments, or egocentric embodiments. Their results illustrate how richer, more naturalistic settings can enhance the generalizability of the solutions a system learns. Thus, if we are interested in understanding how a system generalizes, we may need to build models that learn from appropriately naturalistic data.
Example 2: Generalization to novel syntactic structures is enhanced by variability in other structures’ fillers Generalization is of deep interest in linguistics. Recently, Misra and Mahowald, (2024) performed controlled experiments in which challenging linguistic constructions are systematically held out from the language model training data—and showed that models trained on naturalistic language can generalize to held out constructions by composing pieces of simpler constructions. Critically, the authors also found that generalization depended on the variability of the semantic fillers observed in the structures in training—thus showing how incorporating naturalistic variation in model training data can impact our theoretical inferences about generalization.
Example 3: Rats raised in enriched environments. Analogously, animals raised in more complex environments can be more skilled than those raised in simpler environments. This phenomenon has been well-documented in neuroscience, where rats raised in enriched environments — with social interaction and/or more space or toys in their home cage — show greater exploration, and improved learning and memory on novel tasks (Simpson and Kelly,, 2011). This illustrates how, for natural as well as artificial intelligence, increasing the naturalistic variation in experience can alter the system’s learning in ways that may impact our experimental and theoretical inferences.
4.3 Learning with naturalistic data can yield good performance across a range of seemingly disparate tasks
One interesting finding from the deep learning literature is that when deep learning architectures with many parameters are trained with a lot of naturalistic data and an appropriate “basic” learning objective, these architectures can develop mechanisms that go beyond the learning objective they were trained on. Crucially, these learning paradigms can allow models to transfer well (through initial performance or accelerated learning) on novel tasks beyond the training distribution.
Human learning may likewise benefit from transfer among the disparate tasks we learn. We argue that these AI findings motivate computational cognitive science research studying whether “down-stream human behavior” on a set of tasks can be recapitulated by a model which is trained on a large set of naturalistic tasks or stimuli representative of some of aspect of human experience—as in meta-learning approaches (cf. Wang,, 2021). Below, we provide examples of this kind of transfer spanning computer vision, reinforcement learning, and natural language processing.
Example 1: Computer Vision. One of the earliest successes from deep learning came in computer vision. Soon after AlexNet (Krizhevsky et al.,, 2012) achieved strong results on the ImageNet dataset (Deng et al.,, 2009), researchers showed that the features discovered by AlexNet could be repurposed to novel tasks ranging from scene recognition to medical diagnosis (Donahue et al.,, 2014; Sharif Razavian et al.,, 2014; Litjens et al.,, 2017). This was striking because these features were trained (a) via supervision on a fixed set of objects but (b) enabled transfer to novel objects or even different types of tasks. For years this pattern continued, and researchers even showed that the representations learned by these models could (by 2022, unsurprisingly) transfer from Imagenet to both 3D dexterious manipulation tasks and 3D household navigation tasks, sometimes achieving better performance than using “ground-truth” state information (Parisi et al.,, 2022; Yuan et al.,, 2022). In addition to generally useful representations, deep learning vision models trained with naturalistic data also develop mechanisms they were not trained for. For example, scene-oriented CNNs develop mechanisms for object detection (Zhou et al.,, 2014) and vision transformers develop mechanisms for segmenting objects (Caron et al.,, 2021)—both without an explicit training signal.
Example 2: Reinforcement learning. Likewise in reinforcement learning, researchers have found that reinforcement learning algorithms trained on many (e.g. billions) of tasks exhibit strong “out-of-distribution” generalization on unknown tasks (Team et al.,, 2021), with the ability to generalize to novel tasks in the same number of samples as humans (Team et al.,, 2023). Some algorithms can even generalize to collaborating with humans without any human data (Strouse et al.,, 2021). Recently, reinforcement learning algorithms for learning a modern variant of the successor representation (Gershman,, 2018; Carvalho et al.,, 2024) have shown that they develop mechanism for exploration and behavioral skills without explicit training signals (Liu et al.,, 2024).
Example 3: Large language models. Perhaps the most striking example of this general phenomenon comes from large language models. Large language models are only trained to predict the next “token” (or word) to appear following a sequence of tokens. However, the distribution of internet language effectively includes a broad mixture of many tasks (Radford et al.,, 2019). When trained on vast amounts of this naturalistic data, large language models develop mechanisms for disparate tasks such as modular arithithmic, solving word problems, etc. (Wei et al.,, 2022)—and even for adapting to new tasks from examples presented in context (Brown et al.,, 2020). Moreover, the representations learned by these models can transfer to many superficially dissimilar downstream tasks (Lu et al., 2022a, ), even ones as dissimilar as playing video games (Reid et al.,, 2022). These examples illustrate how learning from a broad naturalistic distribution can induce many abilities—and can provide important transfer to downstream tasks.
4.4 Learning from naturalistic data allows us to ask new questions about the origins of knowledge
The points outlined above have an important consequences; using models that learn from naturalistic data can change our theoretical conclusions in cognitive science. For example, if a model fails to generalize when trained on simple data, we cannot be sure if the model or the data are inadequate. By contrast, using models that learn from naturalistic data can enable us to ask more precise questions about which features—of the models or the data—are necessary to reproduce the theoretically-relevant features of cognitive and neural processes.
Example 1: language models and the learnability of language. Recent progress in language modeling from naturalistic data has suggested challenges to prior assumptions about language learnability and innateness. We briefly sketch these developments; Piantadosi, (2023) and Futrell and Mahowald, (2025) give much more thorough accounts. Classical approaches to language processing focused on simple models and ideas (e.g. Chomsky,, 2014), but in doing so imposed strong assumptions. Under certain strong assumptions, Gold, (1967) proved that even relatively simple languages cannot be learned from input alone. This proof contributed to arguments that there must be some innate universals such as recursion, and separation of syntax from semantics (Chomsky,, 1957, 1965, e.g.)—ideas that were subsequently influential in arguments about the nature of cognition more broadly (Fodor et al.,, 1975; Fodor and Pylyshyn,, 1988). These theories were based on studying particular language features in great detail [^2] Or sometimes even studying the intuitions of linguists, with which the average language user would not necessarily even agree (Spencer,, 1973).. Because of this focus, the theories did not attempt to model language processing in all its naturalistic detail.
However, recent progress in deep learning based language models (e.g. Brown et al.,, 2020) shows that many of these assumptions may need revision. By learning from large amounts of naturalistic data, these models acquire both syntax (Manning et al.,, 2020) and semantics (Li et al., 2021a, )—and account for a broad range of behavioral and neuroscientific phenomena (Schrimpf et al.,, 2021). Of course, these models often learn from inhuman quantities of language (Wilcox et al.,, 2024). However, even language models trained on a human-like amount of language data can learn complex syntactic features that have been previously considered evidence for innate gramatical knowledge (e.g. Wilcox et al.,, 2023). Likewise, while some “impossible” languages that (adult) humans struggle to learn had been considered evidence for universal grammar, Kallini et al., (2024) show that language models can more easily learn real languages than impossible ones. Finally, controlled experiments—like those from Misra and Mahowald, (2024), reviewed above—allow us to begin to understand which features of experience can contribute to generalizing to truly-novel structures, and how naturalistic variability contributes to this generalization. Thus, incorporating naturalistic data into has helped to reshape our understanding of the necessary and sufficient features for acquiring linguistic capabilities.
Example 2: naturalistic data can lead to brain-like functional specialization. Given the importance of perceiving faces in our social lives, it might seem likely that face recognition is innately encoded in the brain. Indeed, the discovery of the Fusiform Face Area (FFA)—a visual region specialized for face perception (Kanwisher and Yovel,, 2006)—would seem to lend further support to this hypothesis. However, recent computational work shows that this kind of specialization can emerge from a domain-general computer-vision model purely through training on a naturalistic distribution of object and face recognition tasks (Dobs et al.,, 2022). As the authors state: “It may be that the only inductive bias humans need to develop their face system is the already well-established early preference of infants to look at faces.” Of course, this does not necessarily imply that the FFA is not innately specified. However, it does illustrate how training computational models over naturalistic data can help us to understand the natural constraints that affect neural organization—whether those effects occur over developmental or evolutionary timescales.
5 Building generalizable models
ML research has excelled in developing generalizable models. While we focus on their utility for working on naturalistic stimuli and tasks, we believe cognitive science would broadly benefit from improved practices for developing generalizable models. In this section, we highlight three key practices of empirical ML research that we believe computational cognitive science can adopt. The first is frictionless reproducibility: the practice of developing research artifacts which can be reused, and repurposed with minimal effort (§5.1). The second is a focus on developing generalizable models that can learn successfully across many datasets (§5.2). We argue that these two offer the foundation for having groups of research collectively work on important research problems through the utilization of benchmarks (§5.3).
Before we continue, we acknowledge that modern empirical ML is a young field which, like other fields, has faced challenges in rigor and reproducibility (Melis et al.,, 2017; Lucic et al.,, 2018; Riquelme et al.,, 2018; Henderson et al.,, 2018; Recht et al.,, 2019; Agarwal et al.,, 2021). Despite these challenges, we are optimistic that some of the practical knowledge ML has developed is valuable for improving rigor and reproducibility, as we highlight below.
Figure 6: Overview of key strategies cognitive science can adopt to develop generalizable models. (Figures reproduced from (Geirhos et al.,, 2018; Bowers et al.,, 2023))
Section summary. We envision a future where cognitive scientists identify a cognitive phenomenon of interest (e.g., theory of mind, working memory, or object recognition) and create a phenomena-oriented benchmark (R8) containing experiments specifically designed to test different theoretical aspects of that phenomenon. Here, experiments are implemented with standardized evaluation protocols and interfaces (R1) that enable direct comparison between competing theoretical frameworks. All experimental data and stimuli are made available through one-line download access (R2). Models are developed and released following frictionless reproducibility standards (R1-R3) for easy adoption and modification. Researchers identify gaps in theoretical coverage and contribute new experiments to the benchmark, facilitating maintaining a separation between model and task design (R6). Thanks to the existence of a benchmark and frictionless reproducibility standards, models are evaluated comprehensively across all experimental paradigms with minimal additional human labor. This helps mitigate the challenge of disconnected, incommensurate findings that resist integration across studies (Almaatouq et al.,, 2024). Thanks to the leaderboard principle (R7), this also helps ensure that progress is defined by substantial rather than incremental advances. This approach encourages the development of models that work across both simplified and naturalistic contexts (R4) and that generalize across multiple experimental paradigms (R5). Importantly, in addition to a model that predicts behavior across many tasks, we also have a model that can perform many tasks that humans can—enabling the building of human-like models that scale and generalize.
5.1 Frictionless reproducibility: a superpower
[…]
In recent years, data-centric sciences are experiencing a profound transition to what David Donoho calls “frictionless reproducibility” (Donoho,, 2024; cf. Recht,, 2024). Donoho identifies three key pillars to frictionless reproducibility: data sharing, code sharing, and competitive challenges. We will focus on the first two in this section, and competitive challenges in §5.3. Critically, with frictionless reproducibility, the research artifacts that are produced allow future researchers to exactly re-execute the same complete workflow with minimal effort. With this, good ideas are spread, adopted, and improved with little friction. Donoho argues that ML is ostensibly the most successful at frictionless reproducibility. Below we review some supporting strategies that may aid cognitive science.
Recommendations for cognitive science
R1. Use a standardized evaluation protocol where models can be compared via a common interface. Machine learning’s foundation of standardized evaluation protocols traces back to 1959, when Highleyman created the first alphanumeric dataset and the first ever “train-test” split for comparing pattern recognition (Highleyman and Kamentsky,, 1959). This strategy continued with the famous MNIST alphanumeric dataset (LeCun et al.,, 1998). MNIST simplified data-sharing and evaluation by pre-processing data, fixing the train and test distributions, and sharing the data online. This standardization and usage of a common interface enabled fair comparisons between diverse approaches, from boosted stumps (Kégl and Busa-Fekete,, 2009) to support vector machines (Cristianini and Scholkopf,, 2002) to nearest neighbor methods (Khan,, 2017).
Standardized evaluation protocols serve a critical function: mitigating personal bias in model assessment. The evolution from MNIST through Caltech 101 (Fei-Fei et al.,, 2004) to ImageNet (Deng et al.,, 2009) demonstrates this principle, as increasingly naturalistic datasets enabled fair comparisons between diverse model families. This enabled AlexNet (Krizhevsky et al.,, 2012) to validate deep learning in an era when many researchers doubted its potential.
Thus, we recommend that cognitive scientists establish clear hold-out data for either model development or data analysis. This is especially critical for confirming exploratory results (Poldrack et al.,, 2017). Combined with standardized evaluation protocols, this will enable fair comparison between diverse theoretical frameworks ranging from program synthesis to deep neural networks to reinforcement learning and active inference.
R2. Make data accessible with a single line of code. Fully processed research data should be programmatically accessible with a single line of code, requiring no permissions. While uploading data to platforms like the Open Science Framework (OSF) is valuable, we additionally recommend uploading data to platforms like Hugging Face that offer version control, free hosting, authentication-free access, the ability to download data with one line of code, and to upload data with a few lines of code. This removes friction from reusing and updating datasets.
Model Citations Age GitHub Stars Single file examples
ResNet
He et al., (2016)
247,921 8 years 2,300 model
Contrastive Learning
He et al., (2020)
13,480 4 years 4,800 model, training
Object Discovery
Locatello et al., (2020)
790 4 years 394 model
RL Library
Huang et al., (2022)
252 2 years 5,600 model 1, model 2
RLHF Method
Rafailov et al., (2024)
1,730 1 year 2,200 train
Multi-agent RL
Rutherford et al., (2023)
24 0 years 433 model
Table 1: Comparison of ML libraries that adopted the single file philosophy.
R3. Adopt the “single file” philosophy. Ideally, key workflows (e.g. model or evaluation definititions) should be defined in a single file. This has emerged as a powerful force for reproducibility in ML research (see Table 1 for examples). While this approach contradicts traditional software engineering principles of modularity and abstraction, it offers distinct advantages for research code: (1) complete understanding of an important component (e.g. how a model works) requires reading only one file, (2) minimal abstraction layers and dependencies, and (3) easy adaptation of either complete implementations or specific components. One can simply copy this single file into another codebase. Removing friction from future reuse also benefits us—as we can see from Table 1, these methods get a large number of citations.
5.2 Developing models with an eye towards generalizability
Task-performing computational models have been a goal of cognitive science since  Newell, (1973). This sentiment still exists today (Kriegeskorte and Douglas,, 2018; Almaatouq et al.,, 2024) and is exemplified in conferences that explicitly seek to integrate cognitive science and neuroscience with AI  (Naselaris et al.,, 2018). Cognitive scientists are well aware that they can’t rely on a single paradigm for the studying an effect if they want to achieve generalizable conclusions (Holzmeister et al.,, 2024; Yarkoni,, 2022). Ideally, the models we develop generalize beyond the setting (e.g. stimulus set or task-specification) they were designed for. This requires two things. First, that the model can operate over new stimulus sets and generalize its behavior to new tasks. Second, that the behavioral (and potentially neural) predictions made by our models capture human behavior on new tasks and new stimulus sets. Clearly, the first is a requirement for the second. Here, we argue that machine learning, with its emphasis on frictionless reproducibility, has made significant strides in developing models that can generalize. We detail recommendations for cognitive science below.
Recommendations for cognitive science
R4. Develop models with simplified stimluli but ensure they work with increasingly naturalistic stimuli. The frame problem (Shanahan,, 2004) and the “simulation is doomed to succeed” critique (Grim et al.,, 2013) highlight a fundamental challenge in model development: the evaluation tasks we design often fail to capture the full complexity of the phenomenon we care about. Empirical machine learning addresses this by developing models that must succeed in both simplified and increasingly complex settings. Below we detail two examples.
Example 1: Generative Models. One of the most successful models of machine learning has been the variational autoencoder (Kingma,, 2013), a generative model that learns to infer latent variables which describe an observation. When this model was released they showed results inferring latent variables correctly on both “simplified” MNIST digits and more complex images of real-world faces. Importantly, in this more complicated setting where the latent dimensions did not have obvious correspondence, they performed a qualitative analysis showing their model had learned a face manifold capturing semantically meaningful dimensions of faces such as sentiment and facial direction.
Example 2: Reinforcement Learning. The history of RL is one of progress towards increasing naturalism: initial environments were grid-worlds (Sutton and Barto,, 2018), later 2D video games like atari (Bellemare et al.,, 2013), then virtual 3D home environments (Szot et al.,, 2021; Li et al., 2021b, ), and now an emphasis on open-world environments (Fan et al.,, 2022; Matthews et al.,, 2024; Hughes et al.,, 2024). Despite the increased emphasis on naturalistic settings, there is still emphasis on the importance of toy settings for studying specific aspects of your model (Osband et al.,, 2019; Obando Ceron et al.,, 2024). Indeed, the strategy is almost always to first develop your model for simplified settings, but one should always then see that your model continues to “work” with more naturalistic settings. Collectively, this has allowed RL researchers to both advance a theoretical understanding of how these models work (Dadashi et al.,, 2019; Grimm et al.,, 2020; Lyle et al.,, 2022), an empirical understanding of how they work (Obando Ceron et al.,, 2024; Farebrother et al.,, 2024), but also develop performant models for a wide range of “complex” settings including beating the world’s best players in games like Go (Silver et al.,, 2017) and controlling robots in the real world (Levine et al.,, 2016; Cheng et al.,, 2024).
R5. Develop models that work with many stimulus sets. The ultimate goal of ML has always been generalization (Highleyman and Kamentsky,, 1959; Hardt and Recht,, 2022; Recht,, 2024). Initially, this meant models that generalize to new data. Now, it means both models and learning procedures that generalize to new settings. Thus, the practice of evaluating models across diverse datasets has become foundational in machine learning. This has manifested with research papers where ML models are evaluated on numerous datasets to showcase their generality. Early examples include “dropout” (Srivastava et al.,, 2014), which demonstrated generalization improvements across 6 distinct image and speech datasets, and MoCo (He et al.,, 2020), which validated its contrastive learning approach on 9 computer vision datasets. Another notable example is the seminal “Deep Q-Network” paper (Mnih et al.,, 2015), which showed that a single neural network architecture with fixed hyperparameters [^3] Hyperparameters are parameters that are set for a model and algorithm before training, such as the number of layers in a neural network or the learning rate of the optimizer could master 50 different Atari games. The field has since evolved to group environments by the cognitive capabilities they test (e.g., exploration, generalization, manipulation) (Patterson et al.,, 2023), with researchers developing models on one set of tasks and validating on entirely different ones to ensure robust generalization (Machado et al.,, 2018).
We recommend cognitive scientists also test the models that support their theories more broadly. Importantly, we should develop our models with some stimulus sets or tasks, and then evaluate them (e.g. learn the parameters) on a different set of stimulus sets or tasks (Machado et al.,, 2018). This further helps prevent our computational model (and thereby theory) from “overfitting” to a particular paradigm or set of stimuli.
R6. Iterate between between (1) model-design (2) task design where the other is fixed. One crucial lesson from ML research is the importance of separating model development from task design. Patterson et al., (2023) strongly caution against simultaneously developing both the problem setting (datasets or tasks) and the solution method, as researchers may inadvertently design evaluation environments that favor their proposed solution. Historical examples illustrate this risk. When developing Q-learning algorithms, researchers modified the classic pendulum control problem by adding episode cutoffs and random state resets (Machado et al.,, 2018). While these modifications made Q-learning more tractable by implicitly improving exploration, they obscured fundamental limitations in Q-learning’s exploration capabilities compared to policy gradient methods. Maintaining a strict separation between task design and model development enforces a clear delineation between hypothesis formation and testing, which is important for reducing experimenter bias. This is especially important because researchers tend to find “positive” results (Scheel et al.,, 2021; Sarafoglou et al.,, 2022), a bias exacerbated by an increasingly competitive academic landscape (Lee,, 2014; Reithmeier et al.,, 2019; Woolston,, 2021). Much of ML’s progress can be attributed to researchers evaluating their models on independently developed datasets and tasks—a practice that not only simplifies the research process but also promotes scientific rigor.
Thus, we recommend that cognitive science also develop models on tasks/stimuli that have been developed by others. This may seem to contradict science, where researchers develop a hypothesis and create an experiment to test their hypothesis. The data from prior work was for a different hypothesis! We recommend that the field begin to explore phenomena-oriented benchmarks that aggregate stimuli and tasks designed to test specific theoretical predictions for broad phenomena. This approach would allow researchers to leverage existing experimental paradigms while maintaining scientific rigor through independent validation. We detail this more in the next section.
5.3 Benchmarks: catalysts for progress and innovation
Another challenge that cognitive science faces is fragmentation. This has been true since Newell’s seminal critique (Newell,, 1973), where his observation that piece-wise investigation of brain components would not yield a cohesive understanding remains relevant forty years later (Schrimpf et al.,, 2020; DiCarlo et al.,, 2023). Current experimental paradigms often produce disconnected, incommensurate findings that resist integration across studies, even within the same domain (Almaatouq et al.,, 2024). This fragmentation manifests not only in methodology but also in how researchers distribute their attention across different problems, potentially diluting collective progress.
The history of machine learning offers an instructive parallel. In its early days, researchers worked on disparate, self-defined problems, leading to diffuse progress (Recht,, 2024). The introduction of shared benchmarks marked a crucial shift: researchers began focusing on standardized problems that others had identified as important. We argue that this led subpopulations of researchers to converge on small subsets of problems. Putting aside whether this is good or bad, it’s been effective. Indeed, this is the third pillar of frictionless reproducibility which (Donoho,, 2024) argues has been a major catalyst for progress in ML: competitive challenges. We argue that cognitive science can similarly stand to benefit from benchmarks.
Recommendations for cognitive science
R7. Embrace the leaderboard principle. A common concern with benchmarks is that researchers will “overfit” to test sets, which theoretically should underestimate true test error (Hastie et al.,, 2009). Yet despite widespread iteration on benchmarks in machine learning—a practice that ostensibly violates statistical principles—the field has produced models with strong performance on held-out data. This apparent paradox is partially explained in a recent machine learning textbook (Hardt and Recht,, 2022) by a phenomenon known as the “leaderboard principle”. Researchers typically only publish and build upon models that demonstrate substantial improvements over prior state-of-the-art results, rather than minor variations. This selective pressure towards meaningful advances effectively constrains the degree of adaptation to test sets, as researchers focus on substantial improvements rather than exhaustively exploring the test set’s peculiarities. Still, a risk of overfitting to data remains. We argue that the remedy to this for cognitive science is the creation of phenomena-oriented benchmarks, where researchers can contribute new challenges that identify a missing component of a phenomenon of interest. We detail this below.
R8. Develop dynamic benchmarks to iteratively capture important phenomena. ML benchmarks are not collected to test specific hypotheses (Hardt and Recht,, 2022) and so may seem unscientific. However, we argue that cognitive science can embrace “scientific” benchmarks oriented toward testing specific hypotheses. We use the story of “Brain-Score” as an illustrative example of initial steps in this direction.
Brain-Score (Schrimpf et al.,, 2020) began as a benchmark with a collection of behavioral and neural data from humans and monkeys on object recognition tasks. One common critique of Brain-Score is that the behavioral analysis essentially boils down to whether models and humans found the same stimuli easy and challenging to identify. However, it did not test hypotheses about how humans identify objects, nor show that models identified objects in this same way. This was perhaps best described by Bowers et al., (2023), who noted that matching predictions of humans was not enough to show that this was a good model of human object recognition. In a scathing critique, they reference 9 experiments from the human psychophysics literature which presented stimuli generated to test specific hypothesis and showed that deep neural networks failed to capture human behavior across these. Examples include stimuli sets that tested the hypothesis that object recognition was based on local shape features (Baker et al.,, 2020), stimuli sets that tested the hypothesis that human object recognition is defined by shape more than by texture (Geirhos et al.,, 2018), and stimuli sets that tested the hypothesis that pairs of objects were judged to be similar based on shape contours (Puebla and Bowers,, 2022). In response, DiCarlo et al., (2023) added many of these stimulus sets to Brain-Score so that when future models are tested, they are tested on whether they capture human behavior across all of these stimulus sets.
We believe this story illustrates what cumulative, dynamic naturalistic computational cognitive science research could look like: where researchers can develop models and experiments that easily integrate with prior findings on a phenomenon thanks to phenomena-oriented benchmarks, frictionless reproducibility, and the development of generalization-oriented models. Of course, we must still continue to analyze and study our models to build comprehensive theories, as we discuss in the next section.
6 Building from naturalistic experiments to cognitive theories
Figure 7: Overview of how we can develop theories with potentially opaque models and unnatural manipulations of natural data. (Panel C bottom figure is reproduced from Saxe et al., 2019.)
A primary goal of cognitive science is understanding. Experiments and computational modeling are key constraints on theory building. By instantiating our theories in a model, we are forced to make our theories more precise by concretizing the ambiguous details of the mapping from a verbal hypothesis to its implementation (Guest and Martin,, 2021).
From this perspective, building computational models that perform the task in as naturalistic a way as possible, across as wide a variety of settings as possible, imposes much stronger constraints on our theories than an abstracted model at a higher-level (cf. Cao and Yamins, 2024b, ). These added constraints are important; otherwise, high-level theories that do not directly engage with the details of the implementation can be so unconstrained as to lack explanatory value (cf. Erev and Greiner,, 2015; Jones and Love,, 2011; Rahnev and Denison,, 2018; Andrews,, 2021), can be intractable for real problems (Van Rooij,, 2008), or can elide important details, as we have outlined above.
Yet moving towards naturalistic settings, and the complex models they require, alters the way that we need to approach experimentation and theory building. In this section we discuss the consequences for experiment design (§6.1-6.2), the role and interpretation of models (6.3), and theory building (§6.4).
6.1 Performing controlled experiments by parametrically manipulating naturalistic data (in unnatural ways)
Figure 8: Testing hypotheses by parametrically manipulating naturalistic data. In the natural data distribution (blue), features like shape and texture are highly correlated—cats usually have cat fur, and elephants usually have elephant skin. This confounding can make it difficult to discern how these variables independently contribute to processing in neural systems or models. One approach to deconfounding these variables is to use a narrow distribution of highly-simplified artificial stimuli (light green) that isolate the variables of interest. However, we advocate for (additionally) creating broader distributions of stimuli that preserve more of the breadth of the natural data distribution, while deconfounding or manipulating features parametrically in order to test hypotheses—as in the unnatural combinations of natural shapes and textures created by Geirhos et al., (2018).
Developing and testing models on truly naturalistic data introduces new issues that are not present in simple artificial settings. Naturalistic data distributions may be too “easy”—for example, they may only rarely include edge-cases that test key capabilities (Zhang et al.,, 2023); thus, testing average performance over a naturalistic distribution may not identify important failures. Relatedly, in naturalistic data, features may be confounded, which can prevent readily disentangling how a model is solving a task (Rust and Movshon,, 2005). For example, models trained on ImageNet (Deng et al.,, 2009) rely on features like texture more than humans do (Geirhos et al.,, 2018), but still perform similarly to humans on ImageNet. These issues limit our ability to achieve deep understanding from experimenting with only the distribution of naturalistic data (Rust and Movshon,, 2005).
However, we believe that these challenges are surmountable by parametrically manipulating rich, naturalistic data (Fig. 8)—preserving or enhancing its richness, while more carefully accounting for theoretically motivated constructs. For example, to effectively test processing of challenging syntactic structures, Futrell et al., (2021) edited natural stories to increase the density of these rare constructions, while preserving the richness and naturalism of the original content. Similarly, the work of Geirhos et al., (2018) engineered datasets that combine natural shapes and textures in unnatural ways. To do so, the authors exploited the contemporary ML technique of iterative style transfer Gatys et al., (2016)—combining the features of different images at different spatial scales—thus illustrating how progress in AI models unlocks new experimental approaches. These conflicting shape-texture datasets helped to inspire similar experiments on humans (Jagadeesh and Gardner, 2022b, ; Jagadeesh and Gardner, 2022a, ); surprisingly, those studies found that human visual cortex likewise appears to use textural rather than shape-focused representations, and shape-driven behavior is thus mediated by downstream readout processes. Thus, manipulating features to test models can lead to new insights into brain function.
A growing range of studies have taken similar approaches to systematically manipulate or ablate features of a natural stimulus, while preserving as much of the natural variety as possible—whether introducing new parameters like modality manipulation of a stimulus (e.g. Deniz et al.,, 2019) or systematically selecting stimuli from a broad distribution to drive particular responses while preserving variety (e.g. Tuckute et al.,, 2024; Hosseini et al.,, 2023). We likewise advocate for testing on tasks that augment natural data distributions with unnatural systematic manipulations in accordance with the hypothesis to be tested (cf. Jain et al.,, 2024). Experimenting with rich settings does not mean giving up theory-driven experimental manipulation; it simply means situating those controlled experiments in more complex, naturalistic settings. For example, there has been a growing research direction using games as richer settings for the experimental study of cognition (Allen et al.,, 2024). Moreover, naturalistic experiments enhance our ability to develop generalizable theories, as we argue below.
6.2 Incorporating naturalistic variation to enhance generalizability
By testing hypotheses in settings where other aspects of the task and data generating process are sampled broadly from the naturalistic distribution, we address some of the challenges of generalizability that arise from testing on a much narrower distribution than the written hypothesis implies (cf. Yarkoni,, 2022). For example, if the hypothesis is about visual processing in general, but the task paradigm relies solely on abstract shape stimuli, the conclusions may not generalize in the way the written results would suggest. Thus, testing our hypotheses across varied settings—“design[ing] with variability in mind” (Yarkoni,, 2022)—increases the likelihood of identifying robust effects that will generalize (cf. Baribault et al.,, 2018). In the language of regression, varying as many other data generating factors as possible brings us closer to estimating the “main effect” of an experimentally-manipulated construct, rather than estimating the “simple effect” of the manipulation in the single setting we have tested.
For example, it is a long-standing principle of experimental design to test a hypothesis with multiple stimuli, and statistically quantify the stimulus effect to estimate the generalizability (e.g. Clark,, 1973). However, many other experiment features—even exact task formulations such as multi-arm bandit tasks—are often preserved within a paper, or even across an entire literature. Just like testing a hypothesis with a single stimulus, this lack of variation poses challenges for generalizability (cf. Eckstein et al.,, 2022)—especially because the way that researchers operationalize a theory can substantially affect their conclusions (e.g. Holzmeister et al.,, 2024; Schweinsberg et al.,, 2021; Strickland and Suben,, 2012). While naturalistic variation cannot resolve this issue in full, we believe that incorporating a broader range of the natural variation of settings, stimuli, and tasks within which we expect those theories to hold, will help us to develop more generalizable theories.
Incorporating variation into our experiments can also help us to revise the constructs underlying our theories. For example, Eisenberg et al., (2019) study the putative construct of self-regulation by simultaneously testing a large variety of existing self-regulation measures. The authors find that these measures do not form a unitary construct. Rather, task- and survey-based measures each tap into distinct, multi-factorial constructs, with individual tasks loading primarily onto a subset of the factors in their domain. That is, two different studies that each use only a single measure for self-regulation may be tapping into entirely different constructs. By contrast, by substantially varying the tasks we use to measure a theoretical component like self-regulation, we can more fully disentangle the underlying constructs—and thereby improve the extent to which our theories generalize.
6.3 Interpreting complex models to yield explanations
If we perform experiments in naturalistic settings, and develop complex models that can predict human behavior across them, how does this support explaining the cognitive phenomena of interest? In this section we will review some paths to deriving explanations of phenomena from complex models. [^4] N.B., not all complex systems admit simple, abstract explanations. For example, where natural intelligence incorporates chaotic (Freeman,, 1995) or critical (O’Byrne and Jerbi,, 2022) dynamics, they may alter the kind of explanations we can seek (cf. Kellert,, 1993). A benefit of seeking predictive models as a route to explanation is that the models may be useful even if simple explanations do not follow.
Data properties as explanation & rational analysis. One route to explaining model behavior is as a consequence of the properties of the data from which it learns. This idea stretches back through behaviorism as well as cognitive science; for example, connectionist research has often focused on how naturalistic data could drive cognitive phenomena (McClelland and Jenkins,, 1991; Elman,, 1996; Rogers and McClelland,, 2004). Analogously, in AI, various researchers have taken inspiration from behaviors that emerge in large pretrained models, and similarly attempted to identify minimal properties of naturalistic data that give rise to those behaviors. For example, Chan et al., (2022) explore how the bursty, long-tailed nature of natural data can give rise to in-context learning, and Prystawski et al., (2024) examine how local dependencies can yield certain kinds of sequential reasoning. These works illustrate how rich properties of natural data can be distilled down to the minimal elements that give rise to a behavior—and can thus likewise offer a candidate explanation for the origin of those capabilities. Moreover, explanations of behavior in terms of the property of the data from which a system learns offer a route toward normative explanations of that behavior as a rational solution to constraints—precisely the perspective taken by rational analysis (Anderson,, 1991).
Mechanistic explanation. Understanding that a behavior arises from properties of data does not explain how that behavior is implemented. Fortunately, the internal workings of task-performing computational models can generally be inspected and intervened upon. For example, in contemporary AI research, a variety of works have analyzed the mechanisms that implement model behaviors, either in simplified settings (e.g. Nanda et al.,, 2023; Zhong et al.,, 2024), or even in large models trained on natural data (Geiger et al.,, 2021; Wu et al.,, 2024). A particularly promising approach for cognitive science is illustrated by Distributed Alignment Search (Geiger et al.,, 2024), which attempts to map postulated abstract causal models onto low-level features of the model that play corresponding causal roles—thus enabling the link between abstract models or hypotheses about the computational solution, and evidence about concrete implementational mechanisms. Morever, mechanistic interventions can even be combined with data studies, to examine the causal role of different mechanisms over the course of learning (Singh et al.,, 2024). Appropriate regularization of models can also produce representations that yield greater interpretability—e.g. Miller et al., (2024) train networks to predict animal behavior while restricting their representations, and used these to identify non-optimal features influencing the decisions—thus allowing bottom-up discovery of possible implementations. Thus, various methods for mechanistic study of models can offer powerful routes to experimentally determining sufficient implementations of cognitive processes.
Formal theories. From the way that behavior emerges from data or mechanisms, we can progress to formal theories. Saxe et al., (2019) offers an illustrative example. Prior works observed empirically that pseudo-naturalistic data give rise to various aspects of human-like semantic development and representation in neural networks (Rogers and McClelland,, 2004), such as progressive differentiation. However, these works had not explained why these parallels emerge. Saxe et al., (2019) bridged this gap by formally deriving how the data properties combine with gradient-based learning to yield these cognitive phenomena. This illustrates how preliminary explanations in terms of data can be developed into more rigorous analytic theories. Thus, building naturalistic models can be a stepping stone toward more rigorous, normative theories of intelligence as a rational solution to a particular problem setting—much like prior work in cognitive science (e.g. Anderson,, 1991; Oaksford and Chater,, 2009).
6.4 Theories that combine task-performing models with reductive explanation.
Above, we have argued for the value of embracing the richness of naturalistic tasks, even if it leads us to build more complex task-performing models that are hard to interpret, but that perhaps make more accurate predictions of behavior. In the previous section, we have outlined some of the possibilities (and challenges) of deriving understanding from these models. So what, ultimately, should we seek?
We argue that naturalistic computational cognitive science should seek theories of cognitive phenomena that consist of two components:
  1. 1.
    Task-performing (predictive) models that reproduce the phenomena across the same range of naturalistic stimuli and paradigms as the human/animal subjects.
  2. 2.
    Reductions of these task-performing models to simpler mechanisms, properties, and theories of why these models reproduce the phenomena.
These two components serve different purposes. The first helps to demonstrate that the models are real candidate models of the phenomena in question, rather than models that simply could not scale to the real problem or that are overfit to a particular instantiation of the task. We believe this component is necessary because otherwise the problem space is under-constrained, as we have argued above. Moreover, it provides a test-bed for performing experimentation that would be impossible in reality (e.g. beyond what would be feasible or ethical in the lab), which is useful for both scientific and practical reasons. The second component—where feasible—allows linking these predictions to the explanatory understanding that cognitive scientists have usually sought, and that may help us to generalize our insights beyond the current paradigms. This perspective aligns with past arguments on how deep learning can contribute to understanding in cognitive neuroscience (Saxe et al.,, 2021; Kanwisher et al.,, 2023; Cao and Yamins, 2024b, ; Doerig et al.,, 2023). However, we emphasize that combining these two components makes explicit links between the actual problem the system solves and the theorized constructs underlying that solution. It also ensures that our theories make testable predictions beyond the narrow settings of a particular task paradigm.
7 Discussion
In this paper, we have tried to outline a direction of research that we call “naturalistic computational cognitive science” that aims to build generalizable models of cognition that scale to naturalistic tasks, while still offering routes to explanatory and theoretical understanding. This perspective synthesizes a growing literature on the importance of naturalistic experiments and generalizable models, and grows out of a long-standing focus in cognitive science on explaining a broad range of phenomena. We outline these connections here.
The quest for building general models of natural intelligence. A long history of cognitive science research has sought to develop frameworks with the generality to explain a wide scope of cognitive phenomena. Researchers in connectionist (e.g. McClelland et al.,, 1986, 2003), Bayesian (e.g. Tenenbaum and Griffiths,, 2001), and cognitive architectures (e.g. Ritter et al.,, 2019; Laird,, 2019) paradigms have tried to identify underlying principles or mechanisms that explain many phenomena. However, while the modeling frameworks themselves are general, typical instantiations of these frameworks have built specialized models focused on individual tasks.
Other recent studies have built upon foundation models from AI that can perform many tasks. Some works have applied cognitive paradigms to study the behaviors of these models (e.g. Binz and Schulz,, 2023; Lampinen et al.,, 2024; Buschoff et al.,, 2025). Other recent works have finetuned these models using cognitive data, to create generalizable cognitive models that can predict behavior on new experimental paradigms in areas like vision (e.g. Muttenthaler et al., 2024b, ; Muttenthaler et al., 2024a, ; Fu et al.,, 2024), or the broader space of cognitive tasks that can be presented in language (Binz and Schulz,, 2024; Binz et al.,, 2024).
We see the arguments that we have laid out in this paper as being broadly compatible with many of these approaches and frameworks, but offering a stronger emphasis on the virtuous cycle of expanding the scope and naturalism of our experimental designs, and expanding the generalizability of our models and theories—and anchoring these in the practicalities of achieving frictionless reproducibility.
Towards experiments that involve naturalistic behavior In emphasizing more naturalistic experiments, our perspective aligns with a variety of works that have likewise highlighted the importance of considering the scope of complex naturalistic behaviors. This perspective has been advocated frequently in neuroscience (e.g. Mobbs et al.,, 2021; Cisek and Green,, 2024), where even prominent visual processing signals can be fundamentally different in active, naturalistic paradigms than in simpler classic ones (e.g. Amme et al.,, 2024). Likewise, Wise et al., (2023) advocate for the need to explore more naturalistic environments and behaviors in reinforcement learning, to grapple with the complexity of behavior in environments closer to the real world. We concur with these works, and emphasize that advances in technology both enable richer experimental paradigms, and the models that can perform them. We differ from many of these works in emphasizing the role of models that generalize across task paradigms.
Linking between learning models and the brain A large recent literature has explored the surprising alignment between the representations learned by task-optimized deep learning models and those in the brain (Khaligh-Razavi and Kriegeskorte,, 2014; Yamins et al.,, 2014; Yamins and DiCarlo,, 2016; Schrimpf et al.,, 2021; Sucholutsky et al.,, 2023). The observations of surprising alignment have led to substantial interest and debate. The reactions range from critical argument that failures to capture particular phenomena doom these models (e.g. Bowers et al.,, 2023), to suggestions that deep learning models upend the theoretical assumptions of broad areas of the field (Perconti and Plebe,, 2020; Hasson et al.,, 2020). Other authors have written frameworks that attempt to integrate deep learning models within the existing scientific paradigms (Doerig et al.,, 2023). In this vein, several papers have proposed frameworks for understanding the role of deep learning in cognitive neuroscience (e.g Richards et al.,, 2019; Storrs and Kriegeskorte,, 2019; Cichy and Kaiser,, 2019)—and how these models can contribute to deriving explanatory and theoretical understanding (Cao and Yamins, 2024a, ; Cao and Yamins, 2024b, ; Saxe et al.,, 2021). In this context, Feather et al., (2025) argue for the importance of benchmarks that place stronger constraints on models, by testing alignment of both behavior and representations. As above, our perspective aligns with many of these works, but we place greater emphasis on the complementary role of increasing naturalism of experimental paradigms, and learning from the engineering paradigms of AI, rather than simply adopting its models as artifacts.
Conclusions We have outlined a perspective that advocates for studying cognition through a broad spectrum of naturalistic experimental paradigms, building generalizable models of cognition that can perform naturalistic tasks, and deriving explanatory theories from these models and their simplifications. We have supported this perspective with examples illustrating the value of naturalistic settings and generalizable models, and concrete guidance on the practicalities of building generalizable models and using them to shape our theories. Many individual aspects of our argument align with prior works, but we hope that there is value in synthesizing these perspectives and highlighting their synergy. We hope that this perspective will inspire computational cognitive scientists to continue embracing richer naturalistic experimental paradigms and models.
Acknowledgements
Sam Gershman for guidance and feedback on this article and on how to integrate cognitive science and machine learning methodology. We thank Maria Eckstein, Tyler Bonnen, Jay McClelland, Mike Mozer, Rachel Calcott, Nick Blauch, and Shuze Liu for thoughtful comments and discussions. We thank Felix Hill for many inspirational conversations on generalization, cognition, and more.

### v2

@@
 Abstract
-Artificial Intelligence increasingly pursues large, complex models that perform many tasks within increasingly realistic domains.
-How, if at all, should these developments in AI influence cognitive science?
-We argue that progress in AI offers timely opportunities for cognitive science to embrace experiments with increasingly naturalistic stimuli, tasks, and behaviors; and computational models that can accommodate these changes.
-We first review a growing body of research spanning neuroscience, cognitive science, and AI that suggests that incorporating a broader range of naturalistic experimental paradigms (and models that accommodate them) may be necessary to resolve some aspects of natural intelligence and ensure that our theories generalize.
+How can cognitive science build generalizable theories that span the full scope of natural situations and behaviors?
+We argue that progress in Artificial Intelligence (AI) offers timely opportunities for cognitive science to embrace experiments with increasingly naturalistic stimuli, tasks, and behaviors; and computational models that can accommodate these changes.
+We first review a growing body of research spanning neuroscience, cognitive science, and AI that suggests that incorporating a broader range of naturalistic experimental paradigms, and models that accommodate them, may be necessary to resolve some aspects of natural intelligence and ensure that our theories generalize.
+First, we review cases from cognitive science and neuroscience where naturalistic paradigms elicit distinct behaviors or engage different processes.
+We then discuss recent progress in AI that shows that learning from naturalistic data yields qualitatively different patterns of behavior and generalization, and discuss how these findings impact the conclusions we draw from cognitive modeling, and can help yield new hypotheses for the roots of cognitive and neural phenomena.
 We then suggest that integrating recent progress in AI and cognitive science will enable us to engage with more naturalistic phenomena without giving up experimental control or the pursuit of theoretically grounded understanding.
 We offer practical guidance on how methodological practices can contribute to cumulative progress in naturalistic computational cognitive science, and illustrate a path towards building computational models that solve the real problems of natural cognition—together with a reductive understanding of the processes and principles by which they do so.
+Short Abstract
+How should cognitive science pursue generalizable theories?
+We argue that progress in Artificial Intelligence (AI) offers opportunities to embrace naturalistic experiments and computational models that can accommodate them.
+We first review literature suggesting that building generalizable theories may require a broader range of naturalistic experimental paradigms.
+We then suggest that progress in AI enables incorporating naturalistic paradigms and models without losing experimental control or theoretically-grounded understanding.
+We argue that we are entering a new era in which cognitive science can build computational models that solve the real problems of natural cognition, and derive a reductive understanding of those solutions.
+Keywords: Artificial intelligence, benchmarks, cognitive science, engineering, generalizability, naturalistic, task-performing models, theory
 1 Introduction
-Cognitive scientists build models to make our theories concrete — which offers testable predictions, eliminates ambiguities (Guest and Martin,, 2021), and can reveal unexpected properties of cognition and behavior (McClelland,, 2009).
+Cognitive scientists build models to make our theories concrete — which offers testable predictions, eliminates ambiguities (Guest and Martin,, 2021), and can reveal unexpected properties of cognition  (McClelland,, 2009).
 Cognitive models can range from simplified models of a high-level process (e.g.
 Frank and Goodman,, 2012; Cohen et al.,, 1990), to task-performing models of a particular domain (Newell,, 1973), or beyond (Newell,, 1994).
-Many cognitive paradigms—connectionism (McClelland et al.,, 1986), bayesian inference (Tenenbaum and Griffiths,, 2001), and cognitive architectures (e.g.
+Many cognitive paradigms—connectionism (McClelland et al.,, 1986), bayesian inference (Tenenbaum and Griffiths,, 2001; Tenenbaum et al.,, 2011), cognitive architectures (e.g.
 Ritter et al.,, 2019; Laird,, 2019)—have sought generalizable models that can explain a broad range of behavior from a simpler set of principles.
-Ultimately, these approaches could aspire to build towards ”unified theories of cognition” that can explain and predict natural behavior across many tasks and domains (Newell,, 1994).
-However, we are not there yet.
-In recent years there have been substantial debates about the generalizability of our theories (Yarkoni,, 2022; Eckstein et al.,, 2021), whether our models and theories are adequately constrained (e.g.
-Jones and Love,, 2011; Rahnev and Denison,, 2018), and calls to accommodate a broader scope of naturalistic phenomena (Wise et al.,, 2023; Cisek and Green,, 2024).
+However, despite progress, there is still substantial debate on the generalizability of our theories (Yarkoni,, 2022; Eckstein et al.,, 2021), whether our models and theories are adequately constrained (e.g.
+Jones and Love,, 2011; Rahnev and Denison,, 2018), and calls to accommodate a broader scope of naturalistic phenomena (Nastase et al.,, 2020; Wise et al.,, 2023; Cisek and Green,, 2024).
 Our goal in this paper is to motivate and outline a path that we believe will begin to address many of these challenges.
 The path we propose is motivated by recent practical and conceptual developments in AI.
-In the last ten years AI has progressed substantially in building models that can perform a wide range of tasks in various domains.
-We now have vision models that can classify and segment real-world images in objects without external supervision (He et al.,, 2020; Caron et al.,, 2021), we have reinforcement learning models that can learn new tasks in a domain with human-like efficiency (Team et al.,, 2023), and we have language models which can produce human-like language and solve many language-specified tasks (Radford et al.,, 2019; Wei et al.,, 2022).
+In the last ten years, AI has progressed substantially in building models that can perform a wide range of tasks in various domains.
+We now have vision models that can classify and segment real-world images without external supervision (He et al.,, 2020; Caron et al.,, 2021), reinforcement learning models that can learn new tasks in a domain with human-like efficiency (Team et al.,, 2023), and language models that can produce human-like language and solve many language-specified tasks (Radford et al.,, 2019; Wei et al.,, 2022).
 Surprisingly, such models produce internal representations that capture many features of brain representations (Yamins et al.,, 2014; Schrimpf et al.,, 2021).
-These findings have led to discussion of how such complex models can be explanatory (Cao and Yamins, 2024a, ), and how they should alter our theories (Hasson et al.,, 2020; Perconti and Plebe,, 2020; Piantadosi,, 2023).
+These findings have led to discussion of how such complex models can be explanatory (Cao and Yamins, 2024a, ; Kanwisher et al.,, 2023), and how they should alter our theories (Hasson et al.,, 2020; Perconti and Plebe,, 2020; Piantadosi,, 2023).
 Yet they have also led to debate over how we test our models and interpret the results (Bowers et al.,, 2023; DiCarlo et al.,, 2023; Dentella et al.,, 2024).
@@
 The first section (§2) provides an overview of “naturalistic computational cognitive science”.
-Afterwards (§34), we motivate the utility of naturalistic experimental paradigms and learning-based approached for developing a more complete understanding of cognition.
-The rest of the paper (§56) focuses on how to achieve these goals; how to develop models for naturalistic settings, and how to use naturalistic experiments and models as part of explanatory cognitive theories.
+In (§34), we motivate using naturalistic experimental paradigms and learning-based approached for cognitive research.
+The remaining sections (§56) focuses on how to achieve these goals; how to develop models for naturalistic settings, and how to use naturalistic experiments and models as part of explanatory cognitive theories.
 (Figures reproduced from Saxe et al.,, 2019; Geirhos et al.,, 2018; Doshi and Konkle,, 2023.)
 In this work, we attempt to weave a thread of arguments that synthesizes these literatures, spanning from cognitive science’s motivation and theory development to the practicalities of model engineering and reproducibility.
-We articulate a research strategy that we believe will be helpful in enabling theory-driven cognitive science to achieve deeper understanding of the full range of natural intelligence.
+We articulate a research strategy that we believe will be helpful in enabling theory-driven cognitive science to achieve a deeper understanding of the full range of natural intelligence.
 We present an overview of this framework and our arguments in Figure 1.
@@
 First, in §2, we unpack what we mean by “naturalistic” computational cognitive science, by providing a pragmatic outline of the features that we believe deserve more emphasis in our experimental paradigms and models.
-In particular, we argue that researchers should expand their task paradigms to to more closely approximate the breadth of the settings in which their theoretical constructs would be expected to generalize by both adding relevant variables that may interact with those in question, and by incorporating more variability within existing parameters.
+We argue that researchers should expand their task paradigms to more closely approximate the breadth of the settings in which their theoretical constructs should be expected to generalize—first, by adding relevant variables that may interact with the theoretical constructs in question, and second by incorporating more variability within existing parameters.
 Likewise, researchers should expand their models to more faithfully approximate the computational task that the natural system is solving.
-In the next two sections, we motivate these goals, by arguing that increasing naturalism is necessary for developing generalizable understanding.
-First, in §3, we review strands of evidence spanning neuroscience, cognitive science, and AI, showing that naturalistic experimental paradigms can lead to different behavior, engage brain systems differently, and expose expose computational challenges that engage mechanisms differently.
-As an example, Francis et al., (2017) found that moral judgments on a trolley problem were dramatically different in a virtual reality presentation compared to text vignettes.
-Other works that we review have found that processes ranging from visual processing (Amme et al.,, 2024), learning (Rosenberg et al.,, 2021; Collins,, 2024), memory (Helbing et al.,, 2020) and more are changed when naturalistic elements are introduced.
-The works that we review in this section thus suggest that experiments that span a broader range of naturalistic settings are essential for drawing correct inferences about the system as a whole.
-Next, in §4, we highlight the benefits of models that learn from naturalistic data—which can perform a wide range of tasks, and yield qualitatively different patterns of generalization than models trained in simplified settings.
-In turn, these differences can change the inferences we make from our models, and yield new insights about the origins of cognitive and neural phenomena.
-Together, these two sections of our paper argue that pursuing naturalistic experiments, and models trained on them, will be necessary for achieving complete understanding of cognition.
+In the next two sections, we motivate these goals by arguing that increasing naturalism is necessary for developing generalizable understanding.
+First, in §3, we review strands of evidence spanning neuroscience, cognitive science, and AI, showing that naturalistic experimental paradigms can yield different behavior, and alter how tasks engage computational mechanisms or neural systems.
+For example, Francis et al., (2017) found that moral judgments on a trolley problem were dramatically different in a virtual reality presentation compared to text vignettes.
+Other work has found that processes ranging from visual processing (Amme et al.,, 2024), learning (Rosenberg et al.,, 2021; Collins,, 2024), memory (Helbing et al.,, 2020; Hasson et al.,, 2015) and more are changed when naturalistic elements are introduced.
+These works suggest that experiments that span a broader range of naturalistic settings can be essential for drawing correct inferences about the system as a whole.
+Next, in §4, we highlight the benefits of models that learn from naturalistic data—which can yield qualitatively different patterns of generalization than models trained in simplified settings.
+These differences can change the inferences we make from our models, and yield new insights about the origins of cognitive and neural phenomena.
+Together, these two sections of our paper argue that pursuing naturalistic experiments, and models that accommodate them, may be necessary for achieving a complete understanding of cognition.
 In the remainder of the paper, we turn to the practice of naturalistic computational cognitive science, and its contributions to building generalizable cognitive theories.
-First, in §5, we review key lessons computational cognitive scientists can adopt from AI in developing generalizable models that can operate in increasingly naturalistic data conditions: the benefits of frictionlessly reproducing prior research artifacts, the importance of generalizability-focused model development, and the value of benchmarking for concentrating research effort on important questions.
-Finally, in §6, we outline how one can leverage these potentially complicated and opaque models to contribute to generalizable theories of cognition—uniting task-performing predictive models with explanatory reductive understanding, using established techniques like rational analysis (Anderson,, 1990) and new approaches to interpreting complex models (Geiger et al.,, 2021).
+First, in §5, we outline key lessons from AI in developing generalizable task-performing models for increasingly naturalistic data conditions: the benefits of being able to frictionlessly reproducing research artifacts, the importance of generalizability-focused model development, and the value of benchmarking for concentrating research effort on important questions.
+Finally, in §6, we outline how complicated experiments and opaque models can contribute to generalizable theories of cognition—uniting task-performing predictive models with explanatory reductive understanding, using established techniques like rational analysis (Anderson,, 1990) and new approaches to interpreting complex models (Geiger et al.,, 2021).
 We close in §7 by discussing the broader context and implications of naturalistic computational cognitive science.
+1.1 What this article is not
+We first address some potential misconceptions we anticipate about our arguments.
+First, we do not intend to argue that simple experimental paradigms are useless.
+Rather, we argue that incorporating more naturalistic settings alongside simpler ones is essential to developing more generalizable cognitive theories.
+We also do not intend to promote a particular class of models, aside from our emphasis on learning.
+We believe that our arguments are compatible with many learning-based cognitive models, e.g.
+Bayesian cognitive models (Lee and Wagenmakers,, 2013), reinforcement learning (Collins,, 2019), predictive representations (Carvalho et al.,, 2024), program synthesis models (Rule et al.,, 2024), etc.
+Finally, we do not advocate for black-box foundation models trained directly to reproduce experimental data as the goal of naturalistic computational cognitive science.
+While this can be a useful tool (e.g.
+for experimental design), we advocate for theory-driven development of predictive models that capture human behavior as an important tool in the quest to develop and validate generalizable cognitive theories.
+In summary, this paper is not intended to reject the current experimental, modeling, and theoretical approaches of cognitive science.
+Rather, we advocate for extending from simpler experiments to more naturalistic ones.
+We then advocate for building task-performing models that can predict human behavior across these paradigms, and linking these models to simpler reductive models and theories of that behavior.
+We argue that this approach will enable building more generalizable theories that explicitly link cognitive capacities to the full scope of natural behavior to which those capacities contribute.
 Figure 2: Overview of naturalistic computational cognitive science.
@@
 If this model successfully predicts human behavior, researchers see if their model continues to generalize human predictions in increasingly naturalisic conditions.
+Figure 3: Examples of increasingly naturalistic settings that we can now study in a theory-driven manner.
+All of these are settings where tasks and stimuli can now be parametrically generated—i.e., thanks to “generative AI”, we can now automate the generation of photorealistic synthetic data and virtual worlds; thanks to virtual and augmented reality, we can now scan and parametrically manipulate real environments.
+Thanks to progress in AI, we can now build task-performing models that can operate in these stimuli and task settings.
 2 What is “naturalistic” computational cognitive science?
-Naturalistic computational cognitive science is a research strategy for theory-driven cognitive science that aims to predict human behavior across increasingly naturalistic stimuli and tasks.
+Naturalistic computational cognitive science is a research strategy for theory-driven cognitive science that aims to predict and understand human behavior across increasingly naturalistic settings.
 This approach gradually increases the ecological validity of experimental designs while ensuring models can accommodate both simplified and more natural conditions.
-As demonstrated in machine learning (§5), this strategy has enabled models that can work with real-world data, albeit imperfectly.
+In machine learning (§5), analogous strategies have enabled models that can work with real-world data, albeit imperfectly.
 As in other approaches within computational cognitive science, this approach starts with formulating models and designing controlled experiments to isolate core phenomena.
@@
   1. 1.
-    Changing existing experiment parameters to better cover the scope of natural distributions.
-For example, in an object classification task, we might expand the category set to include the broad range of objects people encounter in daily their life.
+    Expanding existing experiment parameters to better cover the scope of natural distributions.
+For example, in an object classification task, we might expand the category set to include the many objects people encounter in daily life.
   2. 2.
@@
 We provides examples in Figure 3.
-Figure 3: Examples of increasingly naturalistic settings that we can now study in a theory-driven manner.
-All of these are settings where tasks and stimuli can now be parametrically generated—i.e., thanks to “generative AI”, we can now automate the generation of photorealistic synthetic data and virtual worlds; thanks to virtual and augmented reality, we can now scan and parametrically manipulate real environments.
-Thanks to progress in AI, we can now build task-performing models that can operate in these stimuli and task settings.
-Our goal is to encourage cognitive scientists to embrace experimental paradigms and models that capture a broader spectrum of the variability in inputs, interactions, and tasks present in the natural environment of humans or animals—including, for humans, the rich cultural constructs that form so much of our modern experience.
-That is, we identify a paradigm as more naturalistic if it incorporates a broader scope of the range of natural variability over which the theoretical constructs would be expected to generalize.
-Correspondingly, we identify a model as more naturalistic if it is capable of generalizing to make predictions across a broader range of contexts.
+We describe a paradigm as more naturalistic if it incorporates a broader scope of the range of natural variability over which the theoretical constructs would be expected to generalize.
+We describe a model as more naturalistic if it is capable of generalizing its predictions across a broader range of contexts.
 Note that increasing naturalism does not always mean faithfully approximating all the joint statistics of the natural distribution—increasing the scope of variation along naturalistic dimensions can also enable moving outside the natural data distribution for controlled experiments (see §6.1).
-Generally, identifying the parameters for increasing naturalism in a particular experiment is not well defined or obvious—it is task- and theory-specific.
-To illustrate “naturalistic” more explicitly, we will therefore give some examples of dimensions along which naturalism can be increased.
+Generally, identifying the parameters for increasing naturalism in a particular experiment is not trivial—it is task- and theory-specific.
+To illustrate “naturalistic” more explicitly, we therefore offer some examples of dimensions along which naturalism can be increased.
 Task paradigm
-  • Incorporating multiple paradigms across which a hypothesis would be expected to hold, rather than testing on a single task; for example testing multi-step RL tasks as well as bandit tasks.
-  • Incorporating a broader set of stimuli, for example augmenting synthetic images with real images, or images with videos.
-  • Having task stimuli generated by many varying latent factors, not just the ostensible variables of theoretical interest.
+  • Incorporating multiple paradigms across which a hypothesis would be expected to hold, rather than relying on a single task; for example testing multi-step RL tasks as well as bandit tasks.
+  • Incorporating broader stimulus sets, for example augmenting synthetic images with real images, or images with videos.
+  • Having task stimuli generated by many varying latent factors, not just the variables of ostensible theoretical interest.
 Environment
   • Large state spaces that reflect the complexity of real-world environments (Wise et al.,, 2023).
-  • Including the presence of other social agents that continue to learn with the agent (Hu et al.,, 1998) and that the organism must interact with (Vélez and Gweon,, 2021).
-  • A continually changing environment (Abel et al.,, 2024)
+  • Incorporating other social agents that learn with the agent (Hu et al.,, 1998) and that the organism must interact with (Vélez and Gweon,, 2021).
+  • A continually changing environment (Abel et al.,, 2024).
 Model architecture
-  • Architectures that can operate over sensory inputs like natural images (Yamins et al.,, 2014), speech (Kell et al.,, 2018), or natural language (Schrimpf et al.,, 2021) rather than simplified stimuli (e.g.
-low-dimensional or discrete inputs that contain only task-relevant features).
-  • Architectures with naturalistic action spaces—for example, modeling low-level motor control and embodiment rather than abstract, symbolic actions (Merel et al.,, 2019).
+  • Architectures that operate over sensory inputs like natural images (Yamins et al.,, 2014), speech (Kell et al.,, 2018), or natural language (Schrimpf et al.,, 2021) rather than simplified stimuli (e.g.
+low-dimensional or discrete inputs containing only task-relevant features).
+  • Architectures with naturalistic action spaces—for example, modeling motor control and embodiment rather than abstract, symbolic actions (Merel et al.,, 2019).
 Learning algorithm
@@
   • Social (Henrich,, 2016) or cultural (Cook et al.,, 2024) learning algorithms.
-  • Prospective learning algorithms that aim to model how the tasks evolve in continually changing environments (Seligman et al.,, 2013; De Silva et al.,, 2023).
-Of course, it is easy to say that increasing the naturalism of our experiments and models would be useful, but how can we actually achieve it?
-A key goal for the later sections of our paper (§5-6) will be to give practical perspectives on how to achieve these goals, and how to develop theories that generalize across a broader range of naturalistic behavior.
-We briefly sketch these perspectives here.
-How do we scale up to naturalistic experiments?
-We believe that recent developments in computer science and engineering make increasing naturalism in our experimental paradigms much more accessible to researchers, as we can use natural language to programmatically generate stimuli such as 3D scenes (Zhou et al.,, 2024) and videos of objects (Villegas et al.,, 2022), and leverage haptic feedback in virtual reality to simulate textures (Lu et al., 2022b, ).
-While the specific parameters will vary by domain, the core principle remains: gradually increasing ecological validity while maintaining experimental control.
-How do we glean understanding from potentially opaque learning based models?
-(§6) We believe that task performing models offer powerful new tools for building cognitive theories.
-In a sense, these tools amount to treating the cognitive model as a proxy object for study—but one that admits much more opportunity for experimentation and understanding.
-For example, we can use “neuroscience”-like causal methods to probe a model’s computations by intervening on them (e.g.
-Geiger et al.,, 2021).
-We can also interpret its behavior as an adaptation to the naturalistic properties of its training data (Chan et al.,, 2022; Prystawski et al.,, 2024, e.g.)—as a kind of rational analysis (Anderson,, 1990; Lewis et al.,, 2014) of behavior and computations being a rational response to an environment.
-When applying these approaches to computational models, we have the ability to explicitly test our assumptions in ways we cannot with human subjects (e.g., how would the system behave if it were trained on only ungrammatical sentences).
-We can also attempt to distill our insights further, into analytic abstractions that illustrate the minimal instantiation of a phenomenon.
-By combining these routes, we can build rigorous links from an opaque task-performing model to conceptual understanding of cognition.
-In summary naturalistic computational cognitive science seeks to (1) explain real-world intelligent behavior by developing models that operate over the scope of naturalistic inputs, produce naturalistic outputs, and are optimized according to the actual constraints and affordances of an person’s environment, and (2) develop cognitive theories that unite these models with reductive understanding.
-In the coming sections, we will first motivate more deeply how naturalistic experimental settings can help us develop more complete accounts of intelligence, before returning in more detail to the questions of how these goals can be achieved.
+  • Prospective algorithms that aim to model how the tasks evolve in continually changing environments (Seligman et al.,, 2013; De Silva et al.,, 2023).
+In summary, naturalistic computational cognitive science seeks to (1) explain real-world intelligent behavior by developing models that operate over naturalistic inputs, produce naturalistic outputs, and are optimized according to the actual constraints and affordances of the environment, and (2) develop cognitive theories that unite these models with reductive understanding.
+In the coming sections, we will first motivate how naturalistic experimental settings can help us develop more complete accounts of intelligence, before returning in more detail to how these goals can be achieved.
 3 The benefits of naturalistic experimental settings
@@
 In this section, we outline the benefits of naturalistic experimental settings for cognitive science.
-Specifically, we argue that naturalistic experimental settings can engage mechanisms that are qualitatively different than those engaged in simpler settings — and thereby change the scientific inferences we draw.
-We review examples where naturalistic experimental settings have changed our understanding of a system, and have helped to disentangle competing models that simpler settings could not.
+Specifically, we argue that naturalistic experimental settings can engage mechanisms that are qualitatively different than those engaged in simpler settings — and thereby alter the scientific inferences that we draw.
+We review examples where naturalistic experimental settings have changed our understanding of a system, or have helped to disentangle competing models that simpler settings could not.
 As an orienting conceptual example, imagine that we are doctors studying heart function, and we chose to only study it while the participants are lying down and resting.
-We could certainly learn interesting things thereby; the resting setting offers high test-retest reliability, and a highly-controlled environment to study the processes like how respiratory cycles affect heart rate.
-However, it offers this control and reliability precisely by removing the many factors of natural variation that interact with heart rate, and indeed for which heart function evolved — such as rapid adaptation to movement, stress, etc.
-Studying heart function across a broader range of naturalistic settings would be necessary to understand why the system is the way it is, and how all the physiological (and psychological) processes involved interact.
-For example, if we experimentally manipulated breathing, but did so across a range of different naturalistic activities (sitting, speaking, running, etc.) we would more effectively determine how breathing affects heart function, and how that effect depends on the state of the overall system.
-That is, by studying phenomena across a range of naturalistic settings we elaborate our understanding of the system, and can build more complete theories of its function.
+We could certainly learn interesting things thereby; the resting setting offers high test-retest reliability, and a highly-controlled environment to study processes like how respiration affects heart rate.
+However, this control and reliability comes precisely from removing the many factors of natural variation that affect the heart, and indeed for which heart function evolved — such as rapid adaptation to movement, stress, etc.
+Studying heart function across a broader range of naturalistic settings would be necessary to understand the system in full, and how the physiological (and psychological) processes involved interact.
+For example, if we experimentally manipulated breathing across a range of different naturalistic activities (sitting, speaking, running, etc.) we would more effectively determine its effect on heart function, and how that effect varies across contexts.
+By studying phenomena across varied naturalistic settings we elaborate our understanding of the system, and can build more complete theories of its function.
 3.1 Behavior can be different in naturalistic experiments
-The starting point for many cognitive analyses is behaviors on a task designed to isolate some cognitive capability.
-However, there are many cases in which task-relevant behavior can be altered by seemingly-orthogonal features of the task—thus raising the risk that our theories will be overfit to a restricted task setting unless we also consider a broading range of variation.
+Many cognitive analyses begin by studying a task designed to isolate some cognitive capability.
+However, in many cases task-relevant behavior can be altered by seemingly-orthogonal features of the task—thus raising the risk that our theories will be overfit to a restricted setting unless we also consider a broader range of variation.
 In this section, we review some examples illustrating changes in behavior or capabilities in naturalistic settings.
 Example 1: moral behavior can change when moving from textual vignettes to virtual reality.
-Recent research has shown that behavioral choices can change when moving from text to virtual reality as it can mark a shift from moral judgment (i.e.
-deciding whether things are moral) to moral action (Francis et al.,, 2017).
-When presented with the classic trolley dilemma in text form, only 40% of participants say they would pull a switch to divert a trolley that would kill one person to save five others.
-However, when the same scenario is presented in virtual reality where participants must simulate the action, 55% choose to pull the switch.
-This difference becomes even more pronounced in the “footbridge” variant of the dilemma.
-In the text version, only 10% of participants say they would push someone off a bridge to save five others.
-Yet when able to simulate this action in virtual reality, 65% of participants choose to push the person.
+Recent research shows that behavioral choices can change when moving from text to virtual reality, by driving a shift from moral judgment (i.e.
+deciding whether things are moral) to moral action (Francis et al.,, 2017, see Figure 4 A).
+When participants were presented the classic trolley dilemma in text form, only 40% said they would pull a switch to divert the trolley, thus killing one person to save five others.
+However, when the same scenario was presented in virtual reality, where participants must simulate the action, 55% chose to pull the switch.
+The difference was even more pronounced in the “footbridge” variant of the dilemma.
+In the text version, only 10% of participants said they would push someone off a bridge to save five others.
+However, in virtual reality, 65% of participants choose to push the person.
 These findings demonstrate that people’s stated moral preferences in hypothetical scenarios can differ substantially from their actions in more naturalistic settings.
@@
 When participants are explicitly instructed to memorize objects in a 3D home environment, their subsequent recall accuracy was significantly lower than when they incidentally encounter the same objects during visual search tasks (Helbing et al.,, 2020).
-This effect extends to spatial memory as well.
-Not only do participants better remember the identity of objects encountered during search, they also show more accurate memory for object locations—placing objects closer to their original positions when they were initially seen during search versus explicit memorization.
-These findings demonstrate that tasks focused explicitly on memory may not capture all aspects of how memory naturally operates during everyday behavior.
-Additionally, it suggests that computational models of memory may need to be revised to account for these differences in naturalistic settings.
-This is supported by research showing that people display detailed memory of object and spatial information in real-world scenes (Bainbridge et al.,, 2019).
-Example 3: mice can learn $1000\times$1000 × faster during natural behavior, compared to two-alternative forced choice tasks.
+This effect also extends to spatial memory—participants were able to place objects closer to their original positions in the search condition.
+These findings demonstrate that tasks focused explicitly on memory may not capture all aspects of how memory operates during naturalistic behavior.
+Thus, computational models of memory may need to be revised to account for these differences in naturalistic settings.
+Example 3: mice can learn $1000\times$1000 × faster during natural behavior, compared to two-alternative forced choice (2AFC) tasks.
 Recent research has shown that learning rates can dramatically differ between simplified laboratory tasks and more naturalistic settings (Rosenberg et al.,, 2021).
-When mice are tested in traditional two-alternative forced choice (2AFC) tasks, they typically require 10,000 trials over 3-6 weeks of training to reach asymptotic performance of only about 67% accuracy.
-However, when allowed to freely explore a labyrinth environment (similar to a natural burrow systems (Small,, 1901)), mice demonstrate remarkably accelerated learning (Rosenberg et al.,, 2021).
-In this setting, mice learned to navigate a complex maze with 63 T-junctions in just a few hours, discovering and optimizing their path to a water reward after only about 10 reward experiences.
-The learning rate in the naturalistic setting was approximately $1000\times$1000 × times faster than in traditional 2AFC paradigms.
-This dramatic difference in learning speed demonstrates how traditional simplified experimental paradigms may fundamentally underestimate an organism’s learning capabilities, and suggests that natural learning mechanisms may operate differently when they are engaged through more naturalistic behavior.
+When mice are tested in traditional 2AFC tasks, they typically require 10,000 training trials over 3-6 weeks to reach asymptotic performance of only about 67% accuracy.
+However, when allowed to freely explore a labyrinth (similar to natural burrow systems (Small,, 1901)), mice demonstrate remarkably accelerated learning (Rosenberg et al.,, 2021).
+In this setting, mice learned to navigate a complex maze with 63 T-junctions in just a few hours, optimizing their path to a reward after only about 10 reward experiences.
+The learning rate in the naturalistic setting was approximately $1000\times$1000 × times faster than in 2AFC paradigms.
+This dramatic difference in learning speed demonstrates how traditional simplified experimental paradigms may fundamentally underestimate an organism’s learning capabilities, and suggests that learning mechanisms may operate differently when they are engaged through more naturalistic behavior.
 3.2 Neural systems can operate differently under naturalistic conditions
-In addition to behavioral differences like those reviewed above, naturalistic conditions can result in changes to neural coding and computation.
-Thus, in order to arrive at a complete understanding of neural function, we need to consider the responses of the system across a range of simpler and more naturalistic settings.
-In this section, we highlight some examples of different computations or computational roles that arise when moving from standard paradigms to more naturalistic ones.
-Example 1: the model of superior colliculus shifted from a model of eye movement to one of integrating sensory inputs for body control.
-Cisek, (1999) suggests that the focus on computational behavior as an input-output mapping neglects the fundamental fact that natural intelligence evolved not to produce single responses, but for closed-loop control in an environment.
-This setting yields a rather different interpretation of the system’s representations and processes.
-More generally, many researchers have argued that cognition cannot be understood completely outside its embodiment and the environment in which cognitive processes are instantiated (e.g., Newen et al.,, 2018).
-For example, Cisek and Green, (2024) argues that as we increased naturalism from head-fixed to head-free to body-free settings when studying monkeys, we expanded our theory of superior colliculus from controlling saccadic eye movement to generally integrating multimodal cues to guide bodily action-selection (Cisek and Green,, 2024).
-That is, as we increased the naturalism of the experimental conditions that we used to study monkeys, we arrived at a more complete model of superior colliculus.
-Example 2: visual processing systems operate differently during natural viewing compared to passive viewing paradigms.
-Traditional research has shown that one of the earliest visual responses in the brain (known as P100/M100) occurs about 100ms after a person’s eyes land and fixate on a new location (Vinje and Gallant,, 2000).
-However, this finding comes primarily from experiments where participants passively view images while keeping their eyes still.
-In contrast, natural vision involves actively moving our eyes multiple times per second to sample information from different parts of a scene.
-When the researchers examined brain responses during this more natural active viewing, they found that the P100/M100 response actually begins when the eye movement (saccade) starts, not when it ends at the new fixation point (Amme et al.,, 2024).
-Furthermore, they discovered that the neural patterns during active viewing were opposite (anti-correlated) to those seen during passive viewing, suggesting fundamentally different processing mechanisms.
-These findings reveal that simplified experimental paradigms, while valuable, may not capture all aspects of how visual processing actually operates during natural behavior.
-Example 3: model-based learning systems are used for learning moral judgments, but model-free learning are used for learning to avoid harming others.
-Despite moral reasoning being strongly associated with model-based learning systems in the prefrontal cortex, neural evidence shows a surprising shift when people actually learn to avoid harming others (Lockwood et al.,, 2020).
+In addition to behavioral differences like those reviewed above, naturalistic conditions can alter neural coding and computation.
+In this section, we highlight some examples where different computations or roles arise when moving from standard paradigms to more naturalistic ones.
+Example 1: the role of superior colliculus in eye movements vs.
+integrating sensory inputs for body control.
+Cisek and Green, (2024) argues that as we increased naturalism from head-fixed to head-free and body-free settings for studying monkeys, we expanded our theory of superior colliculus from controlling saccadic eye movement to generally integrating multimodal cues to guide bodily action-selection (Figure 4 B).
+That is, as we increased the naturalism of the experimental conditions, we arrived at a more complete model of superior colliculus.
+This illustrates the more general point, that overly focusing on computation as an input-output mapping neglects the fundamental fact that natural intelligence evolved not for single responses, but for closed-loop control in an environment (Cisek,, 1999).
+This perspectives yields a different interpretation of the system’s representations and processes.
+More generally, many researchers argue that cognition cannot be understood completely outside its embodiment and environment (e.g., Newen et al.,, 2018).
+Example 2: visual processing differs during natural viewing compared to passive viewing paradigms.
+One early visual response in the brain (the P100/M100) occurs about 100ms after a person’s eyes fixate on a new location (Vinje and Gallant,, 2000).
+However, this finding comes primarily from experiments where participants passively view images while fixating.
+In contrast, natural vision involves frequent active eye movements to sample information throughout a scene.
+When Amme et al., (2024) examined brain responses during more natural active viewing, they found that the P100/M100 response actually begins when the eye movement (saccade) begins, not when it ends at a new fixation point.
+Furthermore, the neural patterns during active viewing were anti-correlated to those seen during passive viewing, suggesting fundamentally different processing mechanisms.
+These findings reveal that simplified experimental paradigms, while valuable, may not capture all aspects of how visual processing operates during natural behavior.
+Example 3: model-based systems are used for learning moral judgments, but model-free systems are used for learning to avoid harming others.
+Despite moral reasoning being strongly associated with model-based learning in the prefrontal cortex, neural evidence shows a surprising shift when people actually learn to avoid harming others (Lockwood et al.,, 2020).
 Moral reasoning commonly engages model-based systems in the lateral prefrontal cortex (LFPC) (Spitzer et al.,, 2007; Carlson and Crockett,, 2018) with LPFC disruption leading to reduced norm compliance and enforcement (Knoch et al.,, 2006; Ruff et al.,, 2013).
-However, when participants engage in tasks requiring them to learn moral behavior (rather than just reason about it), researchers instead observed neural activity consistent with model-free learning.
-Similar to our moral judgment vs.
-moral action example from §3.1, this shows that when an experiment becomes more naturalistic (not just reasoning about morality but learning moral behavior), neural systems can operate in unexpected ways.
+However, when participants perform tasks requiring them to learn moral behavior (rather than just reason about it), researchers instead observed neural activity consistent with model-free learning.
+As in §3.1, this illustrates that increasing naturalism of experiments (learning to take moral actions rather than simply reasoning about them), neural systems can operate in unexpected ways.
 3.3 Naturalistic experimental paradigms can expose computational challenges that engage mechanisms differently
-Finally, naturalistic paradigms can likewise introduce challenges that are important to understanding functioning of the system as a whole.
-In this section, we highlight examples where the challenges posed by increasing naturalism can engage computational mechanisms differently—which can be useful for disentangling underlying mechanisms, and is important for achieving generalizable understanding.
+Here, we highlight examples where the challenges posed by increasing naturalism can engage computational mechanisms differently—which can help to disentangle underlying mechanisms, and achieve more generalizable understanding.
 Example 1: Different learning mechanisms only yield distinct predictions under working memory load.
-Overly-simplified tasks can yield aliasing of different solutions, where many different computational approaches yield the same behavior.
-Collins, (2024) presents an example, by showing that what seems superficially to be implemented as a standard reinforcement learning algorithm, on closer examination may instead be implemented through a combination of working memory and simpler value-free associative learning.
-[^1] Whether this alternative counts as an alternative implementation of RL or not is orthogonal to our argument; the important point is that the two models yield distinct predictions in some regimes but not others.
-However, the difference between these two implementations cannot be identified in standard task settings, as Collins notes: “Even simple tasks designed to elicit a target process (such as bandit tasks for RL) recruit multiple other processes, but those processes may be unidentifiable in such tasks.
-Disentangling multiple processes requires considering more complex tasks to elicit differentiable behavior.” The more complex tasks in question simply increase the stimulus set size within learning blocks to a handful of objects, rather than restricting to two or three—i.e., more broadly covering the range of set sizes more clearly disentangles the learning processes.
+One challenge of overly-simplified tasks is potential aliasing of different solutions, where many different computational approaches yield the same behavior.
+Collins, (2024) presents one example where a behavior ostensibly implemented as a standard reinforcement learning algorithm may instead be produced from simpler value-free associative learning combined with working memory.
+[^1] Whether this alternative counts as an alternative implementation of RL is orthogonal to our argument; our point is that the models yield distinct predictions in only some experimental regimes.
+They note that “Disentangling multiple processes requires considering more complex tasks to elicit differentiable behavior.” The more complex tasks in question simply increase the set of stimuli within learning blocks to a handful of objects, rather than only two or three—i.e., expanding the range of set sizes more clearly disentangles the learning processes.
 Thus, by restricting to minimal paradigms, we enable a system to use many solutions and prevent ourselves from discriminating between them.
-By contrast, if we impose a broad range of evaluations on the system (cf.
-Nau et al.,, 2024)—and especially naturalistic evaluations that increase demands along different axes of task difficulty—the increased constraints actually make it easier to map the model on to natural intelligence, as is highlighted in the contravariance principle of Cao and Yamins, 2024b .
+By contrast, by imposing a broad range of evaluations on the system (cf.
+Nau et al.,, 2024)—especially naturalistic evaluations that increase demands along different axes of task difficulty—the increased constraints actually make it easier to map the model on to natural intelligence (more constraints on the mapping make it less ambiguous), as is highlighted in the contravariance principle of Cao and Yamins, 2024b .
 Example 2: fear conditioning.
-Traditional fear conditioning research has provided fundamental insights into how humans learn associations between stimuli and aversive outcomes (Sehlmeyer et al.,, 2009; Maren,, 2001; Britton et al.,, 2014).
-Through carefully controlled laboratory studies involving paired stimuli and outcomes, we have developed a rich mechanistic understanding of basic fear learning processes.
-These paradigms have proven especially valuable for studying certain types of naturally occurring fear learning, such as immediate aversive responses to foods (Riley and Tuck,, 1985) or traumatic events like combat exposure (Engelhard et al.,, 2008).
-However, real-world fear learning often involves additional temporal dynamics that complement these well-studied processes.
-For example, a student may develop fear of exams after receiving a single poor grade weeks after taking the test, without requiring repeated negative experiences (Mobbs et al.,, 2021).
-This type of naturalistic learning involves delayed delayed and sparse feedback for choices, i.e.
-it relies on a computational mechanism for credit assignment (Sutton and Barto,, 2018).
-Credit assignment describes how an agent credits choices (e.g.
-studying, socializing, diet, etc.) as responsible for future events (e.g.
+Traditional fear conditioning research offers fundamental insights into how humans learn associations between stimuli and aversive outcomes (Sehlmeyer et al.,, 2009; Maren,, 2001; Britton et al.,, 2014).
+Through carefully controlled laboratory studies, we have developed a rich mechanistic understanding of basic fear learning processes.
+These paradigms have proven especially valuable for studying certain types of naturally occurring fear learning, such as immediate aversive responses to foods (Riley and Tuck,, 1985) or traumatic events (Engelhard et al.,, 2008).
+However, real-world fear learning often involves additional temporal dynamics that complement these well-studied processes (Figure 4 C).
+For example, a student may develop a fear of exams after receiving a single poor grade weeks after taking a test, without requiring repeated negative experiences (Mobbs et al.,, 2021).
+This naturalistic learning involves delayed and sparse feedback.
+Thus, it relies on a computational mechanism for credit assignment (Sutton and Barto,, 2018)—a mechanism for crediting choices (e.g.
+studying or sleep) as responsible for future events (e.g.
 a poor grade).
-This is not to say humans have optimal credit assignment but that they have some mechanism for it and we are empowered to study it with appropriately naturalistic experimental settings.
-Clinical researchers have additionally pointed out that increased ecological validity can increase the applicability of our experimental results as the gap between lab-based fear conditioning and real-world fear has been a barrier to successful treatment of pathological fears (Beckers et al.,, 2013; Scheveneels et al.,, 2016; Krypotos et al.,, 2018).
+Naturalistic settings enable studying important functions such as mechanisms for credit assignment.
+Increased ecological validity can thereby increase the applicability of our findings; indeed, the gap between lab-based fear conditioning and real-world fear has been a barrier to treating pathological fears (Beckers et al.,, 2013; Scheveneels et al.,, 2016; Krypotos et al.,, 2018).
 Example 3: seemingly-similar tasks that engage brain mechanisms differently can be reconciled by naturalistic models.
-There has been a long debate about whether perceptual processes like “oddity” tasks (identifying an object that is different from the rest of the objects in a set) recruit the Medial Temporal Lobe (MTL).
-Specifically, some researchers (e.g Murray et al.,, 2007) have suggested that MTL is recruited to discriminate among “complex” stimulus sets that visual cortex cannot discriminate on its own (and thus MTL lesions impair performance specifically on those stimuli); other researchers have found no such interaction and suggested it stems from methodological issues (Buffalo et al.,, 1998; Suzuki,, 2009).
+There has been a long debate about whether perceptual processes like “oddity” tasks (identifying an object that differs from the others in a set) recruit the Medial Temporal Lobe (MTL).
+Some researchers (e.g Murray et al.,, 2007) suggest that MTL is recruited to discriminate among “complex” stimulus sets that visual cortex cannot discriminate on its own (thus MTL lesions selectively impair performance on those stimuli); other researchers have found no such interaction and suggested it stems from methodological issues (Buffalo et al.,, 1998; Suzuki,, 2009).
 Recently, stimulus-computable models have shed new light on this debate.
@@
 By doing so, Bonnen et al.
-showed that some of the purportedly “complex” stimuli from some of the prior experiments could nevertheless be distinguished with the vision model alone, while the complex stimuli from other experiments could not.
-Furthermore, the purportedly “complex” stimuli that in fact could be discriminated by the pretrained vision network alone did not require the MTL, while the complex stimuli that could not be distinguished by the vision model were precisely the ones that required MTL.
-Thus, the way that prior works operationalized “complexity” was underspecified; differences in operationalization that got abstracted away in conceptual descriptions of the experiments likely drove the seemingly-conflicting findings.
+showed that purportedly “complex” stimuli that could be discriminated by the pretrained vision network alone did not require the MTL, while the complex stimuli that could not be distinguished by the vision model were precisely the ones that required MTL.
+This shows that the way that prior works had operationalized “complexity” was underspecified; differences in operationalization that got abstracted away in conceptual descriptions of the experiments likely drove the seemingly-conflicting findings.
 This example illustrates the value of building models that directly perform the same naturalistic tasks as the subjects, and testing these models on benchmarks that incorporate stimuli from many experimental paradigms.
@@
 Figure 5: Overview of the benefits of learning with naturalistic data. (Figures reproduced from Doshi and Konkle,, 2023; Manning et al.,, 2020; Team et al.,, 2023.)
-The standard lesson in science is that we must simplify an experimental setting so that we can better arrive at a causal conclusion.
-However, we have already seen examples illustrating that more is different (Anderson,, 1972)—certain properties of intelligence may only emerge in more “complicated” naturalistic settings.
-In this section, we detail examples where naturalistic data itself seems to play an importance role in producing the empirical phenomena of intelligence that we wish to study.
-Many of the findings we will discuss here will be modern results from AI.
-One common trend is that there is a positive interaction between learning-based systems and naturalistic data—learning-based systems can accommodate naturalistic data, and reciprocally, learning from naturalistic data results in qualitatively better results than learning in simpler settings.
-While we do not yet know the implications for natural intelligence, they open interesting questions and challenge prior assumptions.
-In particular, we argue that cognitive science should consider the limitations placed on our models when we fail to consider naturalistic data.
-(Note that while we draw many case studies from particular areas of deep learning, we believe that some of these factors could likewise contribute to generalization under other learning paradigms, such as program synthesis, or even evolution—and thus should be of interest even to those who do not work with deep learning models.)
+A standard lesson in science is that we must simplify an experimental setting to better arrive at a causal conclusion.
+However, we have reviewed examples illustrating that more is different (Anderson,, 1972)—some properties of intelligence may only emerge in more “complicated” naturalistic settings.
+Here, we detail examples where naturalistic data itself plays an important role in producing the empirical phenomena of intelligence.
+Many results we discuss are drawn from modern AI research.
+One common trend is a positive interaction between learning-based systems and naturalistic data—learning-based systems can accommodate naturalistic data, and reciprocally, learning from naturalistic data results in qualitatively different generalization than learning in simpler settings.
+While we do not yet know the full implications for natural intelligence, these results open interesting questions and challenge prior assumptions.
+In particular, cognitive science should consider the limitations placed on our models when we fail to consider naturalistic data.
+(While we draw case studies from deep learning, we believe that similar factors could likewise affect generalization under other learning paradigms, such as program synthesis, or even evolution—and thus should be of interest even to those who do not work with deep learning models.)
 4.1 Learning enables models that can operate on natural data
-Perhaps the first surprising benefit of (deep) learning is that it enables a system to operate over natural data.
+One surprising benefit of learning is that it can enable a system to operate over natural data.
 Until AlexNet (Krizhevsky et al.,, 2012), the prevailing wisdom was that one had to design useful representations of natural data for models.
-However, AlexNet and subsequent work showed that a model could learn useful representations for naturalistic data simply by being trained to make predictions about this data.
+However, AlexNet and subsequent work showed that a model could learn useful representations for naturalistic data simply through learning to make predictions about those data.
 Surprisingly, these models learn representations that have notable correspondence to the internal representations in visual cortex (Yamins et al.,, 2014; Khaligh-Razavi and Kriegeskorte,, 2014).
@@
 Learning-based vision models can operate on natural data and predict neural data (surprisingly) well.
-There is a rich history of learning-based methods revealing how neural networks can develop internal representations from pixel-based images that strongly correspond to neural data.
-A seminal example comes from Olshausen and Field, (1996), who showed that V1-like Gabor filters could emerge from a convolutional neural network optimized to transmit input while minimizing overall activation (essentially, a sparse autoencoder).
-Prior to this, V1 activations were modeled using analytically defined Gabor filters.
-This breakthrough demonstrated that V1-like receptive fields could arise naturally from optimization to environmental statistics of edges at varying frequencies and orientations.
-Building on this foundation, Lee et al., (2007) demonstrated that V2-like convolutional filters could emerge from unsupervised optimization of sparse deep belief networks.
-A significant advance came when Yamins and DiCarlo, (2016) showed that task-optimized deep convolutional neural networks trained for object classification developed intermediate layers that matched neural representations in V4 and IT better than any previous models.
-Recent research has revealed an even more striking finding: these task-optimized networks represent images in a latent space whose geometry has remarkable correspondence to brain topography (Doshi and Konkle,, 2023).
-When fitting a 2D sheet to image representations such that neighboring points correspond to nearby points in the original high-dimensional space, the resulting sheet shows striking similarities to actual brain topography.
-Thus, learning-based approaches, particularly neural networks, not only enable effective visual processing but also provide surprisingly accurate predictions of brain organization and function.
+A rich history of work reveals how neural networks can learn internal representations from natural images that strongly correspond to neural data.
+A seminal example comes from Olshausen and Field, (1996), who showed that V1-like Gabor filters could emerge from a convolutional network optimized to transmit input while minimizing overall activation (essentially, a sparse autoencoder).
+This breakthrough demonstrated that V1-like receptive fields could arise from optimization for natural environmental statistics.
+Subsequently, Lee et al., (2007) demonstrated that V2-like convolutional filters could emerge from unsupervised optimization of sparse deep belief networks.
+Later, Yamins and DiCarlo, (2016) showed that deep convolutional networks trained for object classification developed intermediate layers that matched neural representations in V4 and IT better than any previous models.
+Strikingly, task-optimized networks trained to self-organize represent images in a latent space whose geometry corresponds closely to brain topography (Doshi and Konkle,, 2023, Figure 5 A).
+Thus, learning-based approaches not only enable effective visual processing but also can provide surprisingly accurate predictions of brain organization and function.
 Integrating learning-based models with program synthesis enables flexible Bayesian models of human concept learning.
-Recent work has shown how large language models can be combined with program synthesis to create more generalizable Bayesian models of human reasoning (Ellis,, 2024).
-Hand-designed symbolic hypothesis spaces don’t always transfer across domains.
-Ellis, (2024) showed that one can circumvent these challenges by leveraging large language models to enable use of natural language as a flexible hypothesis space with a data-driven prior tuned to human judgments.
+Recently, Ellis, (2024) showed how language models can be combined with program synthesis to create more generalizable Bayesian models of human reasoning.
+Language models enable using natural language as a flexible hypothesis space with a prior derived from human data.
 By translating natural language hypotheses to executable programs for likelihood computation, such models can capture human learning dynamics in online concept acquisition tasks while continuing to predict human behavior robustly in new domains.
-This overcomes traditional limitations of pure symbolic program synthesis approaches which can fail to transfer due to incomplete symbolic vocabularies—a manifestation of the frame problem (Shanahan,, 2004), where when deciding the relevant variables for a problem, we might accidentally choose a framing that doesn’t generalize to other settings of interest.
-The integration of learning-based components with program synthesis thus enables Bayesian models that can flexibly adapt their hypothesis space while maintaining interpretable symbolic computation—a goal of many cognitive modellers.
+This overcomes traditional limitations of hand-designed symbolic hypothesis spaces and pure program synthesis approaches, which can fail to transfer due to incomplete symbolic vocabularies—a manifestation of the frame problem (Shanahan,, 2004) in hypothesis space design.Integrating learning-based components with program synthesis thus enables Bayesian models that can flexibly adapt their hypothesis space while maintaining interpretable symbolic computation—a goal of many cognitive modelers.
 4.2 Learning from naturalistic data can improve generalization
-Studying how systems generalize is a key method in cognitive science and AI.
-Conceptual arguments often rely on empirical studies of generalization; however, these generalization properties are often studied in simple settings, with minimal models learning from minimal features (e.g.
-Marcus,, 1998).
-A fundamental assumption underlying this approach is that the simplifications do not alter the generalization problem.
-However, here we review studies from AI and neuroscience showing that the variability of experience according to seemingly-orthogonal variables can alter how systems learn and generalize—thus implying that studying models within more restricted settings may be misleading about the naturalistic computational problem.
+Studying how systems generalize is fundamental to cognitive science and AI.
+When studying generalization in simplified settings, a key assumption is that the simplifications do not alter the generalization problem.
+Here, we review studies from AI and neuroscience showing that variability in experience (even aloong seemingly-orthogonal axes) can alter learning and generalization—thus implying that studying models within more restricted settings may mislead us about the naturalistic computational problem.
 Example 1: Compositional generalization of image classifiers vs.
@@
 Naturalistic tasks can fundamentally change what models learn and how they generalize.
-Hill et al., (2020) trained two models to do a vision-language grounding task, and tested their compositional generalization to held-out language instructions.
+Hill et al., (2020) trained two models on a vision-language grounding task, and tested their compositional generalization to held-out instructions.
 Both models were trained on the same language examples, and tested on the same held-out examples.
-However, one model was trained as an agent that interacts with a simulated environment, while the other was a simple image-language classifier trained on screenshots from that environment.
-The authors found that the interacting agent exhibited perfect compositional generalization to the novel language utterances, while the image classifier was substantially worse in novel settings.
-The authors also explored a range of other settings, including generalization benefits of more naturalistic 3D (rather than 2D) environments, or egocentric embodiments.
-Their results illustrate how richer, more naturalistic settings can enhance the generalizability of the solutions a system learns.
-Thus, if we are interested in understanding how a system generalizes, we may need to build models that learn from appropriately naturalistic data.
-Example 2: Generalization to novel syntactic structures is enhanced by variability in other structures’ fillers Generalization is of deep interest in linguistics.
-Recently, Misra and Mahowald, (2024) performed controlled experiments in which challenging linguistic constructions are systematically held out from the language model training data—and showed that models trained on naturalistic language can generalize to held out constructions by composing pieces of simpler constructions.
-Critically, the authors also found that generalization depended on the variability of the semantic fillers observed in the structures in training—thus showing how incorporating naturalistic variation in model training data can impact our theoretical inferences about generalization.
+However, one model was trained as an agent interacting with a simulated environment, while the other was simply trained trained to classify screenshots from that environment based on language instructions.
+The authors found that the interacting agent exhibited perfect compositional generalization to novel language utterances, while the image classifier generalized substantially worse.
+The authors also explored other factors, including generalization benefits of more naturalistic 3D (rather than 2D) environments, or egocentric embodiments.
+Their results illustrate how richer naturalistic settings can enhance the generalizability of learned solution.
+Thus, if we are interested in understanding generalization, we may need to build models that learn from appropriately naturalistic data.
+Example 2: Syntactic generalization is enhanced by variability in other structures’ fillers Structural generalization is of deep interest in linguistics.
+Recently, Misra and Mahowald, (2024) performed controlled experiments in which challenging linguistic constructions are systematically held out from language model training data, and showed that models trained on naturalistic language generalize to held out constructions by composing pieces of simpler constructions.
+Critically, this generalization depended on sufficient variability in the semantic fillers observed in the structures seen in training—thus, incorporating naturalistic variation in model training data can impact our theoretical inferences about generalization.
 Example 3: Rats raised in enriched environments.
-Analogously, animals raised in more complex environments can be more skilled than those raised in simpler environments.
-This phenomenon has been well-documented in neuroscience, where rats raised in enriched environments — with social interaction and/or more space or toys in their home cage — show greater exploration, and improved learning and memory on novel tasks (Simpson and Kelly,, 2011).
-This illustrates how, for natural as well as artificial intelligence, increasing the naturalistic variation in experience can alter the system’s learning in ways that may impact our experimental and theoretical inferences.
+Analogously, neuroscientists have found that animals raised in more complex environments can be more skilled than those raised in simpler environments.
+Rats raised in enriched environments — with social interaction and/or more space or toys — show greater exploration, and improved learning and memory on novel tasks (Simpson and Kelly,, 2011).
+Thus, for natural as well as artificial intelligence, greater naturalistic variation in experience can alter the system’s learning in ways that impact our experimental and theoretical inferences.
 4.3 Learning with naturalistic data can yield good performance across a range of seemingly disparate tasks
-One interesting finding from the deep learning literature is that when deep learning architectures with many parameters are trained with a lot of naturalistic data and an appropriate “basic” learning objective, these architectures can develop mechanisms that go beyond the learning objective they were trained on.
+One interesting finding from AI is that when highly-parameterized deep learning architectures are trained with bountiful naturalistic data and appropriate “basic” learning objectives, these architectures can develop mechanisms that go beyond the training objective.
 Crucially, these learning paradigms can allow models to transfer well (through initial performance or accelerated learning) on novel tasks beyond the training distribution.
 Human learning may likewise benefit from transfer among the disparate tasks we learn.
-We argue that these AI findings motivate computational cognitive science research studying whether “down-stream human behavior” on a set of tasks can be recapitulated by a model which is trained on a large set of naturalistic tasks or stimuli representative of some of aspect of human experience—as in meta-learning approaches (cf.
-Wang,, 2021).
-Below, we provide examples of this kind of transfer spanning computer vision, reinforcement learning, and natural language processing.
+We argue that these AI findings should motivate cognitive science research studying whether “downstream human behavior” on a set of tasks can be recapitulated by a model which is trained on varied naturalistic tasks or stimuli representative of (some aspects of) human experience.
+Below, we provide examples illustrating this kind of transfer.
 Example 1: Computer Vision.
-One of the earliest successes from deep learning came in computer vision.
-Soon after AlexNet (Krizhevsky et al.,, 2012) achieved strong results on the ImageNet dataset (Deng et al.,, 2009), researchers showed that the features discovered by AlexNet could be repurposed to novel tasks ranging from scene recognition to medical diagnosis (Donahue et al.,, 2014; Sharif Razavian et al.,, 2014; Litjens et al.,, 2017).
-This was striking because these features were trained (a) via supervision on a fixed set of objects but (b) enabled transfer to novel objects or even different types of tasks.
-For years this pattern continued, and researchers even showed that the representations learned by these models could (by 2022, unsurprisingly) transfer from Imagenet to both 3D dexterious manipulation tasks and 3D household navigation tasks, sometimes achieving better performance than using “ground-truth” state information (Parisi et al.,, 2022; Yuan et al.,, 2022).
-In addition to generally useful representations, deep learning vision models trained with naturalistic data also develop mechanisms they were not trained for.
-For example, scene-oriented CNNs develop mechanisms for object detection (Zhou et al.,, 2014) and vision transformers develop mechanisms for segmenting objects (Caron et al.,, 2021)—both without an explicit training signal.
+One early success of deep learning came in computer vision.
+Soon after AlexNet (Krizhevsky et al.,, 2012) achieved strong results on the ImageNet dataset (Deng et al.,, 2009), researchers showed that AlexNet’s feature could be repurposed to novel tasks like scene recognition or medical diagnosis (Donahue et al.,, 2014; Sharif Razavian et al.,, 2014; Litjens et al.,, 2017).
+This result was striking because features trained via supervision on fixed object classes nevertheless enabled transfer to novel objects or even different types of tasks.
+This pattern has continuend, and more recently researchers showed that the representations learned by these models could transfer from Imagenet 3D household navigation and dexterous tasks, sometimes achieving better performance than using “ground-truth” state (Parisi et al.,, 2022; Yuan et al.,, 2022).
+Beyond their useful representations, vision models trained with naturalistic data also develop mechanisms they were not trained for.
+For example, scene-oriented CNNs develop mechanisms for object detection (Zhou et al.,, 2014) and vision transformers develop mechanisms for segmentation (Caron et al.,, 2021)—both without an explicit training signal.
 Example 2: Reinforcement learning.
-Likewise in reinforcement learning, researchers have found that reinforcement learning algorithms trained on many (e.g.
-billions) of tasks exhibit strong “out-of-distribution” generalization on unknown tasks (Team et al.,, 2021), with the ability to generalize to novel tasks in the same number of samples as humans (Team et al.,, 2023).
+Likewise in reinforcement learning, researchers have found that agents trained on many (billions) of tasks exhibit strong “out-of-distribution” generalization on unknown tasks (Team et al.,, 2021), learning novel tasks as efficiently as humans (Team et al.,, 2023).
 Some algorithms can even generalize to collaborating with humans without any human data (Strouse et al.,, 2021).
-Recently, reinforcement learning algorithms for learning a modern variant of the successor representation (Gershman,, 2018; Carvalho et al.,, 2024) have shown that they develop mechanism for exploration and behavioral skills without explicit training signals (Liu et al.,, 2024).
+Recently, reinforcement learning algorithms for learning a modern variant of the successor representation (Gershman,, 2018; Carvalho et al.,, 2024) have been shown to develop skills for exploration and behavior without explicit training signals (Liu et al.,, 2024).
 Example 3: Large language models.
 Perhaps the most striking example of this general phenomenon comes from large language models.
-Large language models are only trained to predict the next “token” (or word) to appear following a sequence of tokens.
+These models are only trained to predict the next “token” (or word) to appear in a sequence.
 However, the distribution of internet language effectively includes a broad mixture of many tasks (Radford et al.,, 2019).
-When trained on vast amounts of this naturalistic data, large language models develop mechanisms for disparate tasks such as modular arithithmic, solving word problems, etc.
-(Wei et al.,, 2022)—and even for adapting to new tasks from examples presented in context (Brown et al.,, 2020).
-Moreover, the representations learned by these models can transfer to many superficially dissimilar downstream tasks (Lu et al., 2022a, ), even ones as dissimilar as playing video games (Reid et al.,, 2022).
-These examples illustrate how learning from a broad naturalistic distribution can induce many abilities—and can provide important transfer to downstream tasks.
+When trained on these vast naturalistic data, language models develop mechanisms for disparate tasks such as modular arithithmic, solving word problems, etc.
+(Wei et al.,, 2022)—and even for adapting to new tasks from examples (Brown et al.,, 2020).
+Moreover, the representations learned by these models can transfer to superficially dissimilar downstream tasks (Lu et al.,, 2022), including playing video games (Reid et al.,, 2022).
+These examples illustrate how learning from a broad naturalistic distribution can induce many abilities—and can transfer to downstream tasks.
 4.4 Learning from naturalistic data allows us to ask new questions about the origins of knowledge
-The points outlined above have an important consequences; using models that learn from naturalistic data can change our theoretical conclusions in cognitive science.
-For example, if a model fails to generalize when trained on simple data, we cannot be sure if the model or the data are inadequate.
-By contrast, using models that learn from naturalistic data can enable us to ask more precise questions about which features—of the models or the data—are necessary to reproduce the theoretically-relevant features of cognitive and neural processes.
-Example 1: language models and the learnability of language.
-Recent progress in language modeling from naturalistic data has suggested challenges to prior assumptions about language learnability and innateness.
-We briefly sketch these developments; Piantadosi, (2023) and Futrell and Mahowald, (2025) give much more thorough accounts.
+The points outlined above have important consequences: using models that learn from naturalistic data can change our theoretical conclusions in cognitive science.
+For example, if a model fails to generalize from simple data, we cannot be sure whether the model or the data are inadequate.
+By contrast, using models that learn from naturalistic data can enable us to ask more precise questions about which features—of models or data—are necessary to reproduce the theoretically-relevant features of cognitive and neural processes.
+Example 1: language models and learnability.
+Recent progress in language modeling has suggested challenges to prior assumptions about language learnability and innateness.
+We briefly sketch these issues; see Piantadosi, (2023) and Futrell and Mahowald, (2025) for thorough accounts.
 Classical approaches to language processing focused on simple models and ideas (e.g.
@@
 Under certain strong assumptions, Gold, (1967) proved that even relatively simple languages cannot be learned from input alone.
-This proof contributed to arguments that there must be some innate universals such as recursion, and separation of syntax from semantics (Chomsky,, 1957, 1965, e.g.)—ideas that were subsequently influential in arguments about the nature of cognition more broadly (Fodor et al.,, 1975; Fodor and Pylyshyn,, 1988).
-These theories were based on studying particular language features in great detail [^2] Or sometimes even studying the intuitions of linguists, with which the average language user would not necessarily even agree (Spencer,, 1973)..
-Because of this focus, the theories did not attempt to model language processing in all its naturalistic detail.
-However, recent progress in deep learning based language models (e.g.
-Brown et al.,, 2020) shows that many of these assumptions may need revision.
-By learning from large amounts of naturalistic data, these models acquire both syntax (Manning et al.,, 2020) and semantics (Li et al., 2021a, )—and account for a broad range of behavioral and neuroscientific phenomena (Schrimpf et al.,, 2021).
-Of course, these models often learn from inhuman quantities of language (Wilcox et al.,, 2024).
-However, even language models trained on a human-like amount of language data can learn complex syntactic features that have been previously considered evidence for innate gramatical knowledge (e.g.
+This proof contributed to arguments that for innate universals such as recursion, and separation of syntax from semantics (Chomsky,, 1957, 1965, e.g.)—ideas that were subsequently influential in arguments about the nature of cognition more broadly (Fodor et al.,, 1975; Fodor and Pylyshyn,, 1988; Quilty-Dunn et al.,, 2023).
+However, these theories did not attempt to model language processing in all its naturalistic detail.
+By contrast, when deep-learning models learn from large amounts of naturalistic data, these models acquire aspects of both syntax (Manning et al.,, 2020) and semantics (Li et al., 2021a, )—and account for many behavioral and neuroscientific phenomena (Schrimpf et al.,, 2021).
+Although these models often learn from inhuman quantities of language (Wilcox et al.,, 2024), even models trained on a human-like quantity can learn complex syntactic features that were previously considered evidence for innate knowledge (e.g.
 Wilcox et al.,, 2023).
 Likewise, while some “impossible” languages that (adult) humans struggle to learn had been considered evidence for universal grammar, Kallini et al., (2024) show that language models can more easily learn real languages than impossible ones.
-Finally, controlled experiments—like those from Misra and Mahowald, (2024), reviewed above—allow us to begin to understand which features of experience can contribute to generalizing to truly-novel structures, and how naturalistic variability contributes to this generalization.
-Thus, incorporating naturalistic data into has helped to reshape our understanding of the necessary and sufficient features for acquiring linguistic capabilities.
+Finally, controlled experiments—like those from Misra and Mahowald, (2024), reviewed above—begin to illustrate which features of experience contribute to syntactic generalization: particularly the importance of naturalistic variation.
+Thus, studying learnability from naturalistic data has helped to reshape our understanding of the necessary and sufficient features for acquiring linguistic capabilities.
 Example 2: naturalistic data can lead to brain-like functional specialization.
-Given the importance of perceiving faces in our social lives, it might seem likely that face recognition is innately encoded in the brain.
-Indeed, the discovery of the Fusiform Face Area (FFA)—a visual region specialized for face perception (Kanwisher and Yovel,, 2006)—would seem to lend further support to this hypothesis.
+Given the social importance of face perception, it might seem likely that face recognition is innately encoded in the brain.
+Indeed, the discovery of the Fusiform Face Area (FFA)—a visual region specialized for face perception (Kanwisher and Yovel,, 2006)—supports this hypothesis.
 However, recent computational work shows that this kind of specialization can emerge from a domain-general computer-vision model purely through training on a naturalistic distribution of object and face recognition tasks (Dobs et al.,, 2022).
 As the authors state: “It may be that the only inductive bias humans need to develop their face system is the already well-established early preference of infants to look at faces.” Of course, this does not necessarily imply that the FFA is not innately specified.
-However, it does illustrate how training computational models over naturalistic data can help us to understand the natural constraints that affect neural organization—whether those effects occur over developmental or evolutionary timescales.
+However, it does illustrate how training computational models over naturalistic data can help us generate hypotheses about the causes of our neural organization.
 5 Building generalizable models
-ML research has excelled in developing generalizable models.
-While we focus on their utility for working on naturalistic stimuli and tasks, we believe cognitive science would broadly benefit from improved practices for developing generalizable models.
-In this section, we highlight three key practices of empirical ML research that we believe computational cognitive science can adopt.
-The first is frictionless reproducibility: the practice of developing research artifacts which can be reused, and repurposed with minimal effort (§5.1).
+Figure 6: Overview of key strategies cognitive science can adopt to develop generalizable models. (Figures reproduced from Geirhos et al.,, 2018; Bowers et al.,, 2023)
+ML research focuses on developing generalizable models.
+While we focus on their utility for working with naturalistic tasks, we believe cognitive science can broadly benefit from improved practices for developing generalizable models.
+Here, we highlight three key practices of empirical ML research that we believe computational cognitive science can adopt.
+The first is frictionless reproducibility: developing research artifacts that can be reused and repurposed with minimal effort (§5.1).
 The second is a focus on developing generalizable models that can learn successfully across many datasets (§5.2).
-We argue that these two offer the foundation for having groups of research collectively work on important research problems through the utilization of benchmarks (§5.3).
+We argue that these two offer the foundation for the third: collectively collaborating on important research problems through dynamic benchmarks (§5.3).
 Before we continue, we acknowledge that modern empirical ML is a young field which, like other fields, has faced challenges in rigor and reproducibility (Melis et al.,, 2017; Lucic et al.,, 2018; Riquelme et al.,, 2018; Henderson et al.,, 2018; Recht et al.,, 2019; Agarwal et al.,, 2021).
-Despite these challenges, we are optimistic that some of the practical knowledge ML has developed is valuable for improving rigor and reproducibility, as we highlight below.
-Figure 6: Overview of key strategies cognitive science can adopt to develop generalizable models. (Figures reproduced from (Geirhos et al.,, 2018; Bowers et al.,, 2023))
-Section summary.
-We envision a future where cognitive scientists identify a cognitive phenomenon of interest (e.g., theory of mind, working memory, or object recognition) and create a phenomena-oriented benchmark (R8) containing experiments specifically designed to test different theoretical aspects of that phenomenon.
-Here, experiments are implemented with standardized evaluation protocols and interfaces (R1) that enable direct comparison between competing theoretical frameworks.
-All experimental data and stimuli are made available through one-line download access (R2).
-Models are developed and released following frictionless reproducibility standards (R1-R3) for easy adoption and modification.
-Researchers identify gaps in theoretical coverage and contribute new experiments to the benchmark, facilitating maintaining a separation between model and task design (R6).
-Thanks to the existence of a benchmark and frictionless reproducibility standards, models are evaluated comprehensively across all experimental paradigms with minimal additional human labor.
-This helps mitigate the challenge of disconnected, incommensurate findings that resist integration across studies (Almaatouq et al.,, 2024).
-Thanks to the leaderboard principle (R7), this also helps ensure that progress is defined by substantial rather than incremental advances.
-This approach encourages the development of models that work across both simplified and naturalistic contexts (R4) and that generalize across multiple experimental paradigms (R5).
-Importantly, in addition to a model that predicts behavior across many tasks, we also have a model that can perform many tasks that humans can—enabling the building of human-like models that scale and generalize.
+Despite these challenges, we believe that some of the practices of ML are valuable for improving rigor and reproducibility, as we highlight below.
 5.1 Frictionless reproducibility: a superpower
@@
 Donoho identifies three key pillars to frictionless reproducibility: data sharing, code sharing, and competitive challenges.
-We will focus on the first two in this section, and competitive challenges in §5.3.
-Critically, with frictionless reproducibility, the research artifacts that are produced allow future researchers to exactly re-execute the same complete workflow with minimal effort.
-With this, good ideas are spread, adopted, and improved with little friction.
+We focus on the first two in this section, and competitive challenges in §5.3.
+Critically, with frictionless reproducibility, the research artifacts enable future researchers to exactly re-execute the complete workflow with minimal effort.
+Thus, good ideas are spread, adopted, and improved with little friction.
 Donoho argues that ML is ostensibly the most successful at frictionless reproducibility.
-Below we review some supporting strategies that may aid cognitive science.
+Below we review some strategies that may aid cognitive science.
 Recommendations for cognitive science
 R1.
-Use a standardized evaluation protocol where models can be compared via a common interface.
-Machine learning’s foundation of standardized evaluation protocols traces back to 1959, when Highleyman created the first alphanumeric dataset and the first ever “train-test” split for comparing pattern recognition (Highleyman and Kamentsky,, 1959).
-This strategy continued with the famous MNIST alphanumeric dataset (LeCun et al.,, 1998).
-MNIST simplified data-sharing and evaluation by pre-processing data, fixing the train and test distributions, and sharing the data online.
-This standardization and usage of a common interface enabled fair comparisons between diverse approaches, from boosted stumps (Kégl and Busa-Fekete,, 2009) to support vector machines (Cristianini and Scholkopf,, 2002) to nearest neighbor methods (Khan,, 2017).
+Use standardized evaluation protocols where models are compared via a common interface.
+Machine learning’s foundation of standardized evaluation protocols traces back to 1959, when Highleyman and Kamentsky, (1959) created the first alphanumeric pattern-recognition dataset and the first ever “train-test” split.
+This strategy continued with the famous MNIST dataset (LeCun et al.,, 1998).
+MNIST simplified data-sharing and evaluation by pre-processing data, fixing train and test distributions, and sharing the data online.
+This standardization enabled fair comparisons between diverse approaches, from boosted stumps (Kégl and Busa-Fekete,, 2009) to support vector machines (Cristianini and Scholkopf,, 2002) or nearest neighbors (Khan,, 2017).
 Standardized evaluation protocols serve a critical function: mitigating personal bias in model assessment.
-The evolution from MNIST through Caltech 101 (Fei-Fei et al.,, 2004) to ImageNet (Deng et al.,, 2009) demonstrates this principle, as increasingly naturalistic datasets enabled fair comparisons between diverse model families.
+The evolution to Caltech 101 (Fei-Fei et al.,, 2004) and ImageNet (Deng et al.,, 2009) demonstrates this principle, as increasingly naturalistic datasets enabled fair comparisons between diverse model families.
 This enabled AlexNet (Krizhevsky et al.,, 2012) to validate deep learning in an era when many researchers doubted its potential.
-Thus, we recommend that cognitive scientists establish clear hold-out data for either model development or data analysis.
+Thus, we recommend that cognitive scientists establish clear hold-out data for both model development and data analysis.
 This is especially critical for confirming exploratory results (Poldrack et al.,, 2017).
-Combined with standardized evaluation protocols, this will enable fair comparison between diverse theoretical frameworks ranging from program synthesis to deep neural networks to reinforcement learning and active inference.
+Combined with standardized evaluation protocols, this strategy enables fair comparison between diverse theoretical frameworks ranging from program synthesis to deep neural networks to reinforcement learning.
 R2.
@@
 Fully processed research data should be programmatically accessible with a single line of code, requiring no permissions.
-While uploading data to platforms like the Open Science Framework (OSF) is valuable, we additionally recommend uploading data to platforms like Hugging Face that offer version control, free hosting, authentication-free access, the ability to download data with one line of code, and to upload data with a few lines of code.
+While uploading data to platforms like the Open Science Framework (OSF) is valuable, we additionally recommend platforms like Hugging Face that offer version control, free hosting, authentication-free access, access with one line of code, and uploading data with a few lines of code.
 This removes friction from reusing and updating datasets.
@@
 Ideally, key workflows (e.g.
-model or evaluation definititions) should be defined in a single file.
-This has emerged as a powerful force for reproducibility in ML research (see Table 1 for examples).
+model or evaluation definitions) should be defined in a single file.
+This has emerged as a powerful force for reproducibility in ML.
 While this approach contradicts traditional software engineering principles of modularity and abstraction, it offers distinct advantages for research code: (1) complete understanding of an important component (e.g.
-how a model works) requires reading only one file, (2) minimal abstraction layers and dependencies, and (3) easy adaptation of either complete implementations or specific components.
-One can simply copy this single file into another codebase.
-Removing friction from future reuse also benefits us—as we can see from Table 1, these methods get a large number of citations.
+a modell) requires reading only one file, (2) minimal abstraction layers and dependencies, and (3) easy adaptation of prior methods or their components, by simply copying this single file into another codebase.
+Removing friction from future reuse also benefits us—as Table 1 shows, single-file methods get many citations.
 5.2 Developing models with an eye towards generalizability
-Task-performing computational models have been a goal of cognitive science since  Newell, (1973).
-This sentiment still exists today (Kriegeskorte and Douglas,, 2018; Almaatouq et al.,, 2024) and is exemplified in conferences that explicitly seek to integrate cognitive science and neuroscience with AI  (Naselaris et al.,, 2018).
-Cognitive scientists are well aware that they can’t rely on a single paradigm for the studying an effect if they want to achieve generalizable conclusions (Holzmeister et al.,, 2024; Yarkoni,, 2022).
-Ideally, the models we develop generalize beyond the setting (e.g.
-stimulus set or task-specification) they were designed for.
+Task-performing computational models have been a goal of cognitive science since at least  Newell, (1973).
+This sentiment persists today (Kriegeskorte and Douglas,, 2018; Almaatouq et al.,, 2024) and is exemplified in conferences that explicitly seek to integrate cognitive science and neuroscience with AI  (Naselaris et al.,, 2018).
+Cognitive scientists are well aware that relying on a single paradigm to study an effect may reduce the generalizability of conclusions (Holzmeister et al.,, 2024; Yarkoni,, 2022).
+Ideally, the models we develop generalize beyond the setting they were designed for.
 This requires two things.
-First, that the model can operate over new stimulus sets and generalize its behavior to new tasks.
-Second, that the behavioral (and potentially neural) predictions made by our models capture human behavior on new tasks and new stimulus sets.
+First, that the model is generalizable, i.e.
+can learn from, operate over, and perform new tasks.
+Second, that the behavioral (and potentially neural) predictions made by our models capture data on these new tasks.
 Clearly, the first is a requirement for the second.
-Here, we argue that machine learning, with its emphasis on frictionless reproducibility, has made significant strides in developing models that can generalize.
-We detail recommendations for cognitive science below.
+Here, we detail suggestions from ML for how cognitive science can develop more generalizable models.
 Recommendations for cognitive science
 R4.
-Develop models with simplified stimluli but ensure they work with increasingly naturalistic stimuli.
+Develop models with simplified tasks, but ensure they work with increasingly naturalistic tasks.
 The frame problem (Shanahan,, 2004) and the “simulation is doomed to succeed” critique (Grim et al.,, 2013) highlight a fundamental challenge in model development: the evaluation tasks we design often fail to capture the full complexity of the phenomenon we care about.
-Empirical machine learning addresses this by developing models that must succeed in both simplified and increasingly complex settings.
-Below we detail two examples.
-Example 1: Generative Models.
-One of the most successful models of machine learning has been the variational autoencoder (Kingma,, 2013), a generative model that learns to infer latent variables which describe an observation.
-When this model was released they showed results inferring latent variables correctly on both “simplified” MNIST digits and more complex images of real-world faces.
-Importantly, in this more complicated setting where the latent dimensions did not have obvious correspondence, they performed a qualitative analysis showing their model had learned a face manifold capturing semantically meaningful dimensions of faces such as sentiment and facial direction.
-Example 2: Reinforcement Learning.
-The history of RL is one of progress towards increasing naturalism: initial environments were grid-worlds (Sutton and Barto,, 2018), later 2D video games like atari (Bellemare et al.,, 2013), then virtual 3D home environments (Szot et al.,, 2021; Li et al., 2021b, ), and now an emphasis on open-world environments (Fan et al.,, 2022; Matthews et al.,, 2024; Hughes et al.,, 2024).
-Despite the increased emphasis on naturalistic settings, there is still emphasis on the importance of toy settings for studying specific aspects of your model (Osband et al.,, 2019; Obando Ceron et al.,, 2024).
-Indeed, the strategy is almost always to first develop your model for simplified settings, but one should always then see that your model continues to “work” with more naturalistic settings.
-Collectively, this has allowed RL researchers to both advance a theoretical understanding of how these models work (Dadashi et al.,, 2019; Grimm et al.,, 2020; Lyle et al.,, 2022), an empirical understanding of how they work (Obando Ceron et al.,, 2024; Farebrother et al.,, 2024), but also develop performant models for a wide range of “complex” settings including beating the world’s best players in games like Go (Silver et al.,, 2017) and controlling robots in the real world (Levine et al.,, 2016; Cheng et al.,, 2024).
+mchine learning addresses this by developing models that must succeed in both simplified and increasingly complex settings.
+The history of RL shows one such progression, focused on increasing naturalism: initial environments were grid-worlds (Sutton and Barto,, 2018), later 2D video games (Bellemare et al.,, 2013), then virtual 3D homes (Szot et al.,, 2021; Li et al., 2021b, ), and now open-world environments (Fan et al.,, 2022; Matthews et al.,, 2024; Hughes et al.,, 2024).
+Despite this increased emphasis on naturalistic settings, there is still emphasis on the value of simpler settings for studying specific aspects of models (Osband et al.,, 2019; Obando Ceron et al.,, 2024).
+Indeed, the strategy is generally to first develop your model in simplified settings, but then to verify that your model continues to “work” with more naturalistic settings.
+Collectively, this has allowed RL researchers to advance both theoretical (Dadashi et al.,, 2019; Grimm et al.,, 2020; Lyle et al.,, 2022) and empirical (Obando Ceron et al.,, 2024; Farebrother et al.,, 2024) understanding of how their models work, but also develop performant models for many“complex” settings, including beating world experts in Go (Silver et al.,, 2017) and controlling real robots (Levine et al.,, 2016; Cheng et al.,, 2024).
 R5.
-Develop models that work with many stimulus sets.
-The ultimate goal of ML has always been generalization (Highleyman and Kamentsky,, 1959; Hardt and Recht,, 2022; Recht,, 2024).
+Develop models that work on many task.
+The ultimate goal of ML is generalization (Highleyman and Kamentsky,, 1959; Hardt and Recht,, 2022; Recht,, 2024).
 Initially, this meant models that generalize to new data.
 Now, it means both models and learning procedures that generalize to new settings.
-Thus, the practice of evaluating models across diverse datasets has become foundational in machine learning.
-This has manifested with research papers where ML models are evaluated on numerous datasets to showcase their generality.
-Early examples include “dropout” (Srivastava et al.,, 2014), which demonstrated generalization improvements across 6 distinct image and speech datasets, and MoCo (He et al.,, 2020), which validated its contrastive learning approach on 9 computer vision datasets.
-Another notable example is the seminal “Deep Q-Network” paper (Mnih et al.,, 2015), which showed that a single neural network architecture with fixed hyperparameters [^3] Hyperparameters are parameters that are set for a model and algorithm before training, such as the number of layers in a neural network or the learning rate of the optimizer could master 50 different Atari games.
-The field has since evolved to group environments by the cognitive capabilities they test (e.g., exploration, generalization, manipulation) (Patterson et al.,, 2023), with researchers developing models on one set of tasks and validating on entirely different ones to ensure robust generalization (Machado et al.,, 2018).
+Thus, the practice of evaluating models across diverse datasets to showcase their generality has become foundational in ML.
+Early examples include “dropout” (Srivastava et al.,, 2014), which demonstrated improvements across 6 image and speech datasets, and MoCo (He et al.,, 2020), which validated its approach on 9 vision datasets.
+Another notable example is the seminal “Deep Q-Network” paper (Mnih et al.,, 2015), which showed that a single neural network architecture with fixed hyperparameters [^2] E.g., the number of layers or the learning rate.
+could master 50 different Atari games.
+The field has since evolved to organize environments by the cognitive capabilities they test (e.g., exploration, generalization, manipulation) (Patterson et al.,, 2023), with researchers developing models on some task sets and validating on entirely different ones to ensure robust generalization (Machado et al.,, 2018).
 We recommend cognitive scientists also test the models that support their theories more broadly.
-Importantly, we should develop our models with some stimulus sets or tasks, and then evaluate them (e.g.
-learn the parameters) on a different set of stimulus sets or tasks (Machado et al.,, 2018).
-This further helps prevent our computational model (and thereby theory) from “overfitting” to a particular paradigm or set of stimuli.
+Importantly, we should evaluate models on different tasks from those we use to develop them (Machado et al.,, 2018).
+This further helps prevent our model (and thereby theory) from “overfitting” to a particular paradigm.
 R6.
-Iterate between between (1) model-design (2) task design where the other is fixed.
+Iterate between (1) model-design (2) task design where the other is fixed.
 One crucial lesson from ML research is the importance of separating model development from task design.
-Patterson et al., (2023) strongly caution against simultaneously developing both the problem setting (datasets or tasks) and the solution method, as researchers may inadvertently design evaluation environments that favor their proposed solution.
+Patterson et al., (2023) strongly caution against simultaneously developing both the problem setting (datasets or tasks) and the solution method, as researchers may inadvertently design evaluations that favor their proposed solution.
 Historical examples illustrate this risk.
 When developing Q-learning algorithms, researchers modified the classic pendulum control problem by adding episode cutoffs and random state resets (Machado et al.,, 2018).
-While these modifications made Q-learning more tractable by implicitly improving exploration, they obscured fundamental limitations in Q-learning’s exploration capabilities compared to policy gradient methods.
+These modifications made Q-learning more tractable by implicitly improving exploration, but thereby obscured fundamental limitations in Q-learning’s exploration compared to policy gradient methods.
 Maintaining a strict separation between task design and model development enforces a clear delineation between hypothesis formation and testing, which is important for reducing experimenter bias.
 This is especially important because researchers tend to find “positive” results (Scheel et al.,, 2021; Sarafoglou et al.,, 2022), a bias exacerbated by an increasingly competitive academic landscape (Lee,, 2014; Reithmeier et al.,, 2019; Woolston,, 2021).
-Much of ML’s progress can be attributed to researchers evaluating their models on independently developed datasets and tasks—a practice that not only simplifies the research process but also promotes scientific rigor.
-Thus, we recommend that cognitive science also develop models on tasks/stimuli that have been developed by others.
-This may seem to contradict science, where researchers develop a hypothesis and create an experiment to test their hypothesis.
+Much of ML’s progress can be attributed to researchers evaluating their models on independently developed datasets and tasks—a practice that both simplifies the research process and promotes scientific rigor.
+Thus, we recommend that cognitive science also develop models on tasks/stimuli developed by others.
+This may seem to contradict science, where researchers craft experiments to test their specific hypotheses.
 The data from prior work was for a different hypothesis!
-We recommend that the field begin to explore phenomena-oriented benchmarks that aggregate stimuli and tasks designed to test specific theoretical predictions for broad phenomena.
+We recommend that the field develop phenomena-oriented benchmarks that aggregate stimuli and tasks designed to test specific theoretical predictions for broad phenomena.
 This approach would allow researchers to leverage existing experimental paradigms while maintaining scientific rigor through independent validation.
-We detail this more in the next section.
+We detail this further in the next section.
 5.3 Benchmarks: catalysts for progress and innovation
-Another challenge that cognitive science faces is fragmentation.
-This has been true since Newell’s seminal critique (Newell,, 1973), where his observation that piece-wise investigation of brain components would not yield a cohesive understanding remains relevant forty years later (Schrimpf et al.,, 2020; DiCarlo et al.,, 2023).
-Current experimental paradigms often produce disconnected, incommensurate findings that resist integration across studies, even within the same domain (Almaatouq et al.,, 2024).
-This fragmentation manifests not only in methodology but also in how researchers distribute their attention across different problems, potentially diluting collective progress.
+Another challenge for cognitive science is fragmentation.
+Newell’s seminal critique (Newell,, 1973)—that piece-wise investigation of brain components would not yield a cohesive understanding—remains relevant forty years later (Schrimpf et al.,, 2020; DiCarlo et al.,, 2023).
+Current experimental paradigms often produce disconnected, incommensurate findings that resist integration across studies, even within a domain (Almaatouq et al.,, 2024).
+This fragmentation manifests not only in methodology but also in how researchers distribute their attention across different problems, thereby diluting collective progress.
 The history of machine learning offers an instructive parallel.
-In its early days, researchers worked on disparate, self-defined problems, leading to diffuse progress (Recht,, 2024).
-The introduction of shared benchmarks marked a crucial shift: researchers began focusing on standardized problems that others had identified as important.
-We argue that this led subpopulations of researchers to converge on small subsets of problems.
-Putting aside whether this is good or bad, it’s been effective.
-Indeed, this is the third pillar of frictionless reproducibility which (Donoho,, 2024) argues has been a major catalyst for progress in ML: competitive challenges.
+Initially, researchers worked on disparate, self-defined problems, leading to diffuse progress (Recht,, 2024).
+The introduction of shared benchmarks drove a crucial shift: researchers began focusing on standardized problems that others had identified as important.
+Putting aside whether this is good or bad, it has been effective.
+Indeed, this is the third pillar of frictionless reproducibility which Donoho, (2024) argues has been a major catalyst for progress in ML: competitive challenges.
 We argue that cognitive science can similarly stand to benefit from benchmarks.
@@
 Embrace the leaderboard principle.
-A common concern with benchmarks is that researchers will “overfit” to test sets, which theoretically should underestimate true test error (Hastie et al.,, 2009).
-Yet despite widespread iteration on benchmarks in machine learning—a practice that ostensibly violates statistical principles—the field has produced models with strong performance on held-out data.
-This apparent paradox is partially explained in a recent machine learning textbook (Hardt and Recht,, 2022) by a phenomenon known as the “leaderboard principle”.
-Researchers typically only publish and build upon models that demonstrate substantial improvements over prior state-of-the-art results, rather than minor variations.
+A common concern with benchmarks is that researchers will “overfit” to test sets, which theoretically should underestimate the true test error (Hastie et al.,, 2009).
+Yet despite widespread iteration on benchmarks in machine learning the field has produced models with strong performance on held-out data.
+This apparent paradox is partially explained by a phenomenon known as the “leaderboard principle” (Hardt and Recht,, 2022).
+Researchers typically only publish and build upon models that demonstrate substantial improvements over prior results, rather than minor variations.
 This selective pressure towards meaningful advances effectively constrains the degree of adaptation to test sets, as researchers focus on substantial improvements rather than exhaustively exploring the test set’s peculiarities.
-Still, a risk of overfitting to data remains.
-We argue that the remedy to this for cognitive science is the creation of phenomena-oriented benchmarks, where researchers can contribute new challenges that identify a missing component of a phenomenon of interest.
+Nevertheless, a risk of overfitting to data remains.
+We argue that cognitive science can avoid this through the creation of dynamic phenomena-oriented benchmarks, where researchers can contribute new challenges that identify a missing component of a phenomenon of interest.
 We detail this below.
@@
 ML benchmarks are not collected to test specific hypotheses (Hardt and Recht,, 2022) and so may seem unscientific.
-However, we argue that cognitive science can embrace “scientific” benchmarks oriented toward testing specific hypotheses.
+However, cognitive science can embrace “scientific” benchmarks oriented toward testing specific hypotheses.
 We use the story of “Brain-Score” as an illustrative example of initial steps in this direction.
-Brain-Score (Schrimpf et al.,, 2020) began as a benchmark with a collection of behavioral and neural data from humans and monkeys on object recognition tasks.
-One common critique of Brain-Score is that the behavioral analysis essentially boils down to whether models and humans found the same stimuli easy and challenging to identify.
+Brain-Score (Schrimpf et al.,, 2020) began as a benchmark collecting behavioral and neural data from humans and monkeys on object recognition tasks.
+One common critique of Brain-Score is that the behavioral analysis essentially measures whether models and humans found the same stimuli easy or challenging to identify.
 However, it did not test hypotheses about how humans identify objects, nor show that models identified objects in this same way.
-This was perhaps best described by Bowers et al., (2023), who noted that matching predictions of humans was not enough to show that this was a good model of human object recognition.
-In a scathing critique, they reference 9 experiments from the human psychophysics literature which presented stimuli generated to test specific hypothesis and showed that deep neural networks failed to capture human behavior across these.
-Examples include stimuli sets that tested the hypothesis that object recognition was based on local shape features (Baker et al.,, 2020), stimuli sets that tested the hypothesis that human object recognition is defined by shape more than by texture (Geirhos et al.,, 2018), and stimuli sets that tested the hypothesis that pairs of objects were judged to be similar based on shape contours (Puebla and Bowers,, 2022).
-In response, DiCarlo et al., (2023) added many of these stimulus sets to Brain-Score so that when future models are tested, they are tested on whether they capture human behavior across all of these stimulus sets.
-We believe this story illustrates what cumulative, dynamic naturalistic computational cognitive science research could look like: where researchers can develop models and experiments that easily integrate with prior findings on a phenomenon thanks to phenomena-oriented benchmarks, frictionless reproducibility, and the development of generalization-oriented models.
-Of course, we must still continue to analyze and study our models to build comprehensive theories, as we discuss in the next section.
+For example, Bowers et al., (2023) noted that matching human predictions is insufficient to identify a good model of human object recognition.
+The authors reference 9 experiments from the psychophysics literature which presented stimuli generated to test specific hypotheses, and showed that deep neural networks failed to capture human behavior across these.
+Examples include stimuli testing the hypothesis that object recognition was based on local shape features (Baker et al.,, 2020), stimuli testing the hypothesis that human object recognition is defined by shape more than by texture (Geirhos et al.,, 2018), and stimuli testing the hypothesis that pairs of objects were judged to be similar based on shape contours (Puebla and Bowers,, 2022).
+In response, DiCarlo et al., (2023) added many of these stimulus sets to Brain-Score so that when future models are tested on whether they capture human behavior across all of these stimulus sets.
+This story illustrates what an exciting strand of computational cognitive science research could look like: one where researchers develop models and experiments that easily integrate with prior findings on a phenomenon thanks to phenomena-oriented benchmarks, frictionless reproducibility (§5.1), and generalizable models (§5.2).
+By combining these, we achieve a robust foundation for building theories that span the full space of natural behavior associated with a phenomenon.
+We are not the first to identify that cognitive science is fragmented and to suggest benchmarks that aggregate results from the literature as a solution.
+Indeed, researchers spanning memory, timing, confidence and more have already begun to do so (MacWhinney and Snow,, 1985; Oberauer et al.,, 2018; Aydoğan et al.,, 2024; Rahnev et al.,, 2020; Frank et al.,, 2017).
+Here, we both encourage more subdisciplines within cognitive science to follow suite but also emphasize the importance of frictionless reproducibility and building generalizable models.
+While prior work has begun the development of benchmarks, documentation for downloading and using the data is sometimes lacking; download access is not always authentication-free; and maybe most importantly, if one wants to add new data, there isn’t an easy way to re-run all previous models and ensure that findings are commensurate.
+Critically, while prior benchmarks have focused on aggregating experimental results from “simplified” experiments, we advocate for dynamic benchmarks whose development is oriented around explaining data from increasingly naturalistic experiments (i.e.
+experiments with either parametrically generated naturalistic stimuli or, eventually, stimuli taken from real-world conditions).
 6 Building from naturalistic experiments to cognitive theories
@@
 A primary goal of cognitive science is understanding.
-Experiments and computational modeling are key constraints on theory building.
 By instantiating our theories in a model, we are forced to make our theories more precise by concretizing the ambiguous details of the mapping from a verbal hypothesis to its implementation (Guest and Martin,, 2021).
@@
 These added constraints are important; otherwise, high-level theories that do not directly engage with the details of the implementation can be so unconstrained as to lack explanatory value (cf.
-Erev and Greiner,, 2015; Jones and Love,, 2011; Rahnev and Denison,, 2018; Andrews,, 2021), can be intractable for real problems (Van Rooij,, 2008), or can elide important details, as we have outlined above.
+Erev and Greiner,, 2015; Jones and Love,, 2011; Rahnev and Denison,, 2018; Andrews,, 2021), can be intractable for real problems (Van Rooij,, 2008), or can elide important details (see above).
 Yet moving towards naturalistic settings, and the complex models they require, alters the way that we need to approach experimentation and theory building.
-In this section we discuss the consequences for experiment design (§6.1-6.2), the role and interpretation of models (6.3), and theory building (§6.4).
+Here, we discuss the consequences for experiment design (§6.1-6.2), the role and interpretation of models (6.3), and theory building (§6.4).
 6.1 Performing controlled experiments by parametrically manipulating naturalistic data (in unnatural ways)
 Figure 8: Testing hypotheses by parametrically manipulating naturalistic data.
-In the natural data distribution (blue), features like shape and texture are highly correlated—cats usually have cat fur, and elephants usually have elephant skin.
-This confounding can make it difficult to discern how these variables independently contribute to processing in neural systems or models.
+In the natural data distribution (blue), shape and texture are highly correlated—cats usually have fur, and elephants usually have wrinkles.
+This confounding makes it difficult to discern how these variables independently contribute to processing in neural systems or models.
 One approach to deconfounding these variables is to use a narrow distribution of highly-simplified artificial stimuli (light green) that isolate the variables of interest.
 However, we advocate for (additionally) creating broader distributions of stimuli that preserve more of the breadth of the natural data distribution, while deconfounding or manipulating features parametrically in order to test hypotheses—as in the unnatural combinations of natural shapes and textures created by Geirhos et al., (2018).
-Developing and testing models on truly naturalistic data introduces new issues that are not present in simple artificial settings.
-Naturalistic data distributions may be too “easy”—for example, they may only rarely include edge-cases that test key capabilities (Zhang et al.,, 2023); thus, testing average performance over a naturalistic distribution may not identify important failures.
-Relatedly, in naturalistic data, features may be confounded, which can prevent readily disentangling how a model is solving a task (Rust and Movshon,, 2005).
-For example, models trained on ImageNet (Deng et al.,, 2009) rely on features like texture more than humans do (Geirhos et al.,, 2018), but still perform similarly to humans on ImageNet.
+Developing and testing models on truly naturalistic data introduces new issues that are absent in simple artificial settings.
+Naturalistic data distributions may be too “easy”—e.g.
+they may only rarely include edge-cases that test key capabilities (Zhang et al.,, 2023)—thus, testing average performance over naturalistic data may disguise important failures.
+Relatedly, in naturalistic data, features may be confounded, which can prevent readily disentangling the solutions a model uses.
+For example, models trained on ImageNet (Deng et al.,, 2009) rely on features like texture more than humans do (Geirhos et al.,, 2018), but still perform similarly on ImageNet.
 These issues limit our ability to achieve deep understanding from experimenting with only the distribution of naturalistic data (Rust and Movshon,, 2005).
 However, we believe that these challenges are surmountable by parametrically manipulating rich, naturalistic data (Fig.
-8)—preserving or enhancing its richness, while more carefully accounting for theoretically motivated constructs.
-For example, to effectively test processing of challenging syntactic structures, Futrell et al., (2021) edited natural stories to increase the density of these rare constructions, while preserving the richness and naturalism of the original content.
-Similarly, the work of Geirhos et al., (2018) engineered datasets that combine natural shapes and textures in unnatural ways.
+8)—preserving or enhancing its richness, while more carefully manipulating theoretically-motivated constructs (cf.
+Jain et al.,, 2024).
+For example, Geirhos et al., (2018) engineered datasets that combine natural shapes and textures in unnatural ways.
 To do so, the authors exploited the contemporary ML technique of iterative style transfer Gatys et al., (2016)—combining the features of different images at different spatial scales—thus illustrating how progress in AI models unlocks new experimental approaches.
@@
 Thus, manipulating features to test models can lead to new insights into brain function.
-A growing range of studies have taken similar approaches to systematically manipulate or ablate features of a natural stimulus, while preserving as much of the natural variety as possible—whether introducing new parameters like modality manipulation of a stimulus (e.g.
-Deniz et al.,, 2019) or systematically selecting stimuli from a broad distribution to drive particular responses while preserving variety (e.g.
+A growing range of studies similarly systematically manipulate natural stimuli, while preserving as much of the natural variation as possible—whether introducing new parameters like stimulus modality manipulation (e.g.
+listening vs.
+reading; Deniz et al.,, 2019), inserting challenging syntactic structures in natural stories (Futrell et al.,, 2021), or systematically selecting stimuli that drive particular responses while preserving variety (e.g.
 Tuckute et al.,, 2024; Hosseini et al.,, 2023).
-We likewise advocate for testing on tasks that augment natural data distributions with unnatural systematic manipulations in accordance with the hypothesis to be tested (cf.
-Jain et al.,, 2024).
-Experimenting with rich settings does not mean giving up theory-driven experimental manipulation; it simply means situating those controlled experiments in more complex, naturalistic settings.
-For example, there has been a growing research direction using games as richer settings for the experimental study of cognition (Allen et al.,, 2024).
-Moreover, naturalistic experiments enhance our ability to develop generalizable theories, as we argue below.
+Theory-driven, controlled experiments can be performed in more complex, naturalistic settings.
+Another direction for parametrically manipulating naturalistic data comes from using video games to study sequential decision-making and higher-level cognition (Allen et al.,, 2024).
+While games are artificial, different parameters such as the size of a world, the number of agents, the length of tasks, can be parametrically manipulated to reflect values more inline with those found in the real-world.
+As such, simulated environments used in ML offer valuable settings for cognitive experimentation.
+For example, Craftax (Matthews et al.,, 2024), a 2D Minecraft environment, offers parametrically-generated open-world environments containing complex goal hierarchies, which can be used for studying topics like exploration or planning.
+Other environments offer opportunities to study long-horizon generalization (Nikulin et al.,, 2024), coordination (Rutherford et al.,, 2024), physical reasoning (Matthews et al.,, 2025), etc.
+Because these environments are parametrically generated it is easy to create highly-variable-yet-controlled manipulations, e.g.
+many environments that systematically vary planning difficulty.
+Critically, a new tool called NiceWebRL [^3] https://github.com/KempnerInstitute/nicewebrl, enables easily running web-browser human experiments using the environemnts above and the many other ML environments written using Jax (Bradbury et al.,, 2018).
+Thus, there is a growing ecosystem of tools for building complex naturalistic experimental paradigms, and evaluating humans and models with them.
+Moreover, expanding the naturalism of our experiments enhances our ability to develop generalizable theories, as we argue next.
 6.2 Incorporating naturalistic variation to enhance generalizability
@@
 Yarkoni,, 2022).
-For example, if the hypothesis is about visual processing in general, but the task paradigm relies solely on abstract shape stimuli, the conclusions may not generalize in the way the written results would suggest.
-Thus, testing our hypotheses across varied settings—“design[ing] with variability in mind” (Yarkoni,, 2022)—increases the likelihood of identifying robust effects that will generalize (cf.
+In particular, testing our hypotheses across varied settings—“design[ing] with variability in mind” (Yarkoni,, 2022)—increases the likelihood of identifying robust effects that will generalize (cf.
 Baribault et al.,, 2018).
 In the language of regression, varying as many other data generating factors as possible brings us closer to estimating the “main effect” of an experimentally-manipulated construct, rather than estimating the “simple effect” of the manipulation in the single setting we have tested.
-For example, it is a long-standing principle of experimental design to test a hypothesis with multiple stimuli, and statistically quantify the stimulus effect to estimate the generalizability (e.g.
+For example, it is a long-standing principle of experimental design to test a hypothesis with multiple stimuli, and statistically quantify the stimulus effect to estimate generalizability (e.g.
 Clark,, 1973).
 However, many other experiment features—even exact task formulations such as multi-arm bandit tasks—are often preserved within a paper, or even across an entire literature.
-Just like testing a hypothesis with a single stimulus, this lack of variation poses challenges for generalizability (cf.
-Eckstein et al.,, 2022)—especially because the way that researchers operationalize a theory can substantially affect their conclusions (e.g.
+Just like testing a hypothesis with a single stimulus, this limited variation poses challenges for generalizability (Eckstein et al.,, 2022)—especially because the way that researchers operationalize a theory can substantially affect their conclusions (e.g.
 Holzmeister et al.,, 2024; Schweinsberg et al.,, 2021; Strickland and Suben,, 2012).
-While naturalistic variation cannot resolve this issue in full, we believe that incorporating a broader range of the natural variation of settings, stimuli, and tasks within which we expect those theories to hold, will help us to develop more generalizable theories.
-Incorporating variation into our experiments can also help us to revise the constructs underlying our theories.
-For example, Eisenberg et al., (2019) study the putative construct of self-regulation by simultaneously testing a large variety of existing self-regulation measures.
+While naturalistic variation cannot resolve this issue in full, we believe that incorporating a broader range of the settings, stimuli, and tasks within which we expect a theory to hold, will help us to develop more generalizable theories.
+Incorporating variation can also help us to revise the constructs underlying our theories.
+For example, Eisenberg et al., (2019) study the putative construct of self-regulation by simultaneously testing many existing self-regulation measures.
 The authors find that these measures do not form a unitary construct.
-Rather, task- and survey-based measures each tap into distinct, multi-factorial constructs, with individual tasks loading primarily onto a subset of the factors in their domain.
-That is, two different studies that each use only a single measure for self-regulation may be tapping into entirely different constructs.
+Rather, task- and survey-based measures each form distinct, multi-factorial constructs, with individual tasks loading primarily onto a subset of factors.
+That is, two different studies that each use only a single measure for self-regulation may be measuring entirely different constructs.
 By contrast, by substantially varying the tasks we use to measure a theoretical component like self-regulation, we can more fully disentangle the underlying constructs—and thereby improve the extent to which our theories generalize.
@@
 If we perform experiments in naturalistic settings, and develop complex models that can predict human behavior across them, how does this support explaining the cognitive phenomena of interest?
-In this section we will review some paths to deriving explanations of phenomena from complex models.
-[^4] N.B., not all complex systems admit simple, abstract explanations.
-For example, where natural intelligence incorporates chaotic (Freeman,, 1995) or critical (O’Byrne and Jerbi,, 2022) dynamics, they may alter the kind of explanations we can seek (cf.
+Here, we illustrate some paths to deriving explanations from complex models.
+[^4] N.B., not all complex systems admit abstract explanations.
+Where natural intelligence incorporates chaotic (Freeman,, 1995) or critical (O’Byrne and Jerbi,, 2022) dynamics, they may alter the explanations that we can seek (cf.
 Kellert,, 1993).
@@
 Data properties as explanation & rational analysis.
-One route to explaining model behavior is as a consequence of the properties of the data from which it learns.
-This idea stretches back through behaviorism as well as cognitive science; for example, connectionist research has often focused on how naturalistic data could drive cognitive phenomena (McClelland and Jenkins,, 1991; Elman,, 1996; Rogers and McClelland,, 2004).
-Analogously, in AI, various researchers have taken inspiration from behaviors that emerge in large pretrained models, and similarly attempted to identify minimal properties of naturalistic data that give rise to those behaviors.
+One way to explain model behavior is as a consequence of the properties of the data from which it learns.
+This idea has a rich history in cognitive science andd AI (McClelland and Jenkins,, 1991; Elman,, 1996; Rogers and McClelland,, 2004), in which researchers attempted to identify minimal properties of naturalistic data that give rise to some behavior(s).
 For example, Chan et al., (2022) explore how the bursty, long-tailed nature of natural data can give rise to in-context learning, and Prystawski et al., (2024) examine how local dependencies can yield certain kinds of sequential reasoning.
-These works illustrate how rich properties of natural data can be distilled down to the minimal elements that give rise to a behavior—and can thus likewise offer a candidate explanation for the origin of those capabilities.
-Moreover, explanations of behavior in terms of the property of the data from which a system learns offer a route toward normative explanations of that behavior as a rational solution to constraints—precisely the perspective taken by rational analysis (Anderson,, 1991).
+These works illustrate how rich properties of natural data can be distilled down to minimal elements that drive a behavior—and can thus offer a candidate explanation for the origin of those capabilities.
+Moreover, explanations of behavior via properties of the data offer a route toward normative explanations of that behavior as a rational solution to constraints—precisely the perspective taken by rational analysis (Anderson,, 1991).
 Mechanistic explanation.
-Understanding that a behavior arises from properties of data does not explain how that behavior is implemented.
+Understanding that a behavior arises from data properties does not explain how that behavior is implemented.
 Fortunately, the internal workings of task-performing computational models can generally be inspected and intervened upon.
-For example, in contemporary AI research, a variety of works have analyzed the mechanisms that implement model behaviors, either in simplified settings (e.g.
-Nanda et al.,, 2023; Zhong et al.,, 2024), or even in large models trained on natural data (Geiger et al.,, 2021; Wu et al.,, 2024).
-A particularly promising approach for cognitive science is illustrated by Distributed Alignment Search (Geiger et al.,, 2024), which attempts to map postulated abstract causal models onto low-level features of the model that play corresponding causal roles—thus enabling the link between abstract models or hypotheses about the computational solution, and evidence about concrete implementational mechanisms.
+For example, a variety of works have analyzed the mechanisms that implement language model behaviors, either in simplified settings (e.g.
+Nanda et al.,, 2023; Zhong et al.,, 2024), or in large models trained on natural data (Geiger et al.,, 2021; Wu et al.,, 2024).
+A particularly promising approach for cognitive science is illustrated by Distributed Alignment Search (Geiger et al.,, 2024), which attempts to map postulated abstract causal models onto low-level features of a complex model that play corresponding causal roles—thus enabling linking between abstract models or hypotheses about the computational solution, and evidence about concrete implementational mechanisms.
 Morever, mechanistic interventions can even be combined with data studies, to examine the causal role of different mechanisms over the course of learning (Singh et al.,, 2024).
@@
 Formal theories.
-From the way that behavior emerges from data or mechanisms, we can progress to formal theories.
+From data properties or mechanisms, we can progress to formal theories.
 Saxe et al., (2019) offers an illustrative example.
-Prior works observed empirically that pseudo-naturalistic data give rise to various aspects of human-like semantic development and representation in neural networks (Rogers and McClelland,, 2004), such as progressive differentiation.
+Prior works observed empirically that pseudo-naturalistic data elicit various aspects of human-like semantic development and representation in neural networks (Rogers and McClelland,, 2004), such as progressive differentiation.
 However, these works had not explained why these parallels emerge.
@@
 The first helps to demonstrate that the models are real candidate models of the phenomena in question, rather than models that simply could not scale to the real problem or that are overfit to a particular instantiation of the task.
-We believe this component is necessary because otherwise the problem space is under-constrained, as we have argued above.
-Moreover, it provides a test-bed for performing experimentation that would be impossible in reality (e.g.
-beyond what would be feasible or ethical in the lab), which is useful for both scientific and practical reasons.
-The second component—where feasible—allows linking these predictions to the explanatory understanding that cognitive scientists have usually sought, and that may help us to generalize our insights beyond the current paradigms.
+This component is necessary because otherwise the problem space is under-constrained.
+Moreover, it provides a test-bed for experiments that would be impossible in reality, which is useful both scientifically and practically.
+The second component allows linking these predictions to the explanatory understanding that cognitive scientists typically seek, and that may help us to generalize our insights beyond current paradigms.
 This perspective aligns with past arguments on how deep learning can contribute to understanding in cognitive neuroscience (Saxe et al.,, 2021; Kanwisher et al.,, 2023; Cao and Yamins, 2024b, ; Doerig et al.,, 2023).
@@
 7 Discussion
-In this paper, we have tried to outline a direction of research that we call “naturalistic computational cognitive science” that aims to build generalizable models of cognition that scale to naturalistic tasks, while still offering routes to explanatory and theoretical understanding.
+In this paper, we have tried to outline a direction of research that we call “naturalistic computational cognitive science,” which aims to build generalizable models of cognition that scale to naturalistic tasks, while still offering routes to explanatory and theoretical understanding.
 This perspective synthesizes a growing literature on the importance of naturalistic experiments and generalizable models, and grows out of a long-standing focus in cognitive science on explaining a broad range of phenomena.
-We outline these connections here.
+In this section, we outline these connections, and highlight the future directions suggested by our perspective.
 The quest for building general models of natural intelligence.
-A long history of cognitive science research has sought to develop frameworks with the generality to explain a wide scope of cognitive phenomena.
-Researchers in connectionist (e.g.
-McClelland et al.,, 1986, 2003), Bayesian (e.g.
-Tenenbaum and Griffiths,, 2001), and cognitive architectures (e.g.
-Ritter et al.,, 2019; Laird,, 2019) paradigms have tried to identify underlying principles or mechanisms that explain many phenomena.
-However, while the modeling frameworks themselves are general, typical instantiations of these frameworks have built specialized models focused on individual tasks.
-Other recent studies have built upon foundation models from AI that can perform many tasks.
+A long history of cognitive science research has sought frameworks with the generality to explain a wide scope of cognitive phenomena.
+Researchers in cognitive architecture (e.g.
+Ritter et al.,, 2019; Laird,, 2019), connectionist (e.g.
+McClelland et al.,, 1986, 2003), and Bayesian (e.g.
+Tenenbaum and Griffiths,, 2001; Griffiths et al.,, 2010; Tenenbaum et al.,, 2011), paradigms have tried to identify underlying principles or mechanisms that explain many phenomena.
+However, while these modeling frameworks themselves are general, typical instantiations of the frameworks have built specialized models focused on individual tasks.
+Indeed, one core critiques of many modeling paradigms is their failure to accurately capture patterns of human generalization (e.g.
+Fodor et al.,, 1975; McClelland and Patterson,, 2002; Lake et al.,, 2017).
+In pursuit of generality, some recent studies have built upon foundation models from AI that can perform many tasks.
 Some works have applied cognitive paradigms to study the behaviors of these models (e.g.
 Binz and Schulz,, 2023; Lampinen et al.,, 2024; Buschoff et al.,, 2025).
-Other recent works have finetuned these models using cognitive data, to create generalizable cognitive models that can predict behavior on new experimental paradigms in areas like vision (e.g.
+Others have finetuned these models using cognitive data, to create generalizable models that can predict behavior on new experimental paradigms in areas like vision (e.g.
 Muttenthaler et al., 2024b, ; Muttenthaler et al., 2024a, ; Fu et al.,, 2024), or the broader space of cognitive tasks that can be presented in language (Binz and Schulz,, 2024; Binz et al.,, 2024).
-We see the arguments that we have laid out in this paper as being broadly compatible with many of these approaches and frameworks, but offering a stronger emphasis on the virtuous cycle of expanding the scope and naturalism of our experimental designs, and expanding the generalizability of our models and theories—and anchoring these in the practicalities of achieving frictionless reproducibility.
-Towards experiments that involve naturalistic behavior In emphasizing more naturalistic experiments, our perspective aligns with a variety of works that have likewise highlighted the importance of considering the scope of complex naturalistic behaviors.
-This perspective has been advocated frequently in neuroscience (e.g.
-Mobbs et al.,, 2021; Cisek and Green,, 2024), where even prominent visual processing signals can be fundamentally different in active, naturalistic paradigms than in simpler classic ones (e.g.
+We see the arguments that we have laid out in this paper as being broadly compatible with many of these approaches and frameworks.
+However, we offer a stronger emphasis on the role of theory (cf.
+Frank and Goodman,, 2025), and the virtuous cycle between expanding the scope and naturalism of our experimental designs, and expanding the generalizability of our models and theories.
+Towards experiments that involve naturalistic behavior Our perspective aligns with various works that have likewise highlighted the value of considering complex naturalistic behaviors and the environments in which they occur.
+The strands of cognitive research focused on embodiment and enactivism have advocated that the environment, and organism-environment interactions, form an important component of cognition (e.g.
+Clark,, 1998; Varela et al.,, 2017).
+More generally, naturalistic experiments are frequently advocated in neuroscience (e.g.
+Mobbs et al.,, 2021; Cisek and Green,, 2024), where even prominent visual processing signals can fundamentally differ in active, naturalistic paradigms (e.g.
 Amme et al.,, 2024).
+Similarly, Gao and Ganguli, (2015); Gao et al., (2017) highlight that more complex behavioral paradigms are needed to unlock the full space of neural dynamics, and Nastase et al., (2020) argue that isolating only a few variables in our experiments impairs our ability to understand the full scope of natural behavior.
 Likewise, Wise et al., (2023) advocate for the need to explore more naturalistic environments and behaviors in reinforcement learning, to grapple with the complexity of behavior in environments closer to the real world.
 We concur with these works, and emphasize that advances in technology both enable richer experimental paradigms, and the models that can perform them.
-We differ from many of these works in emphasizing the role of models that generalize across task paradigms.
-Linking between learning models and the brain A large recent literature has explored the surprising alignment between the representations learned by task-optimized deep learning models and those in the brain (Khaligh-Razavi and Kriegeskorte,, 2014; Yamins et al.,, 2014; Yamins and DiCarlo,, 2016; Schrimpf et al.,, 2021; Sucholutsky et al.,, 2023).
-The observations of surprising alignment have led to substantial interest and debate.
-The reactions range from critical argument that failures to capture particular phenomena doom these models (e.g.
-Bowers et al.,, 2023), to suggestions that deep learning models upend the theoretical assumptions of broad areas of the field (Perconti and Plebe,, 2020; Hasson et al.,, 2020).
+We differ from many of these works in more strongly emphasizing the role of models that generalize across task paradigms.
+Linking between learning models and the brain A large recent literature has explored surprising alignments between the representations learned by task-optimized deep learning models and neural representations (Khaligh-Razavi and Kriegeskorte,, 2014; Yamins et al.,, 2014; Yamins and DiCarlo,, 2016; Schrimpf et al.,, 2021; Sucholutsky et al.,, 2023).
+These findings have stimulated substantial interest and debate, ranging from critical argument that failures to capture particular phenomena doom these models (e.g.
+Bowers et al.,, 2023), to suggestions that deep learning models upend the theoretical assumptions of the field (Perconti and Plebe,, 2020; Hasson et al.,, 2020).
 Other authors have written frameworks that attempt to integrate deep learning models within the existing scientific paradigms (Doerig et al.,, 2023).
 In this vein, several papers have proposed frameworks for understanding the role of deep learning in cognitive neuroscience (e.g Richards et al.,, 2019; Storrs and Kriegeskorte,, 2019; Cichy and Kaiser,, 2019)—and how these models can contribute to deriving explanatory and theoretical understanding (Cao and Yamins, 2024a, ; Cao and Yamins, 2024b, ; Saxe et al.,, 2021).
-In this context, Feather et al., (2025) argue for the importance of benchmarks that place stronger constraints on models, by testing alignment of both behavior and representations.
+In this context, Feather et al., (2025) argue for the importance of benchmarks that test alignment of both behavior and representations—and highlight how dataset limitations like lack of stimulus variability can limit our ability to discriminate between models.
 As above, our perspective aligns with many of these works, but we place greater emphasis on the complementary role of increasing naturalism of experimental paradigms, and learning from the engineering paradigms of AI, rather than simply adopting its models as artifacts.
-Conclusions We have outlined a perspective that advocates for studying cognition through a broad spectrum of naturalistic experimental paradigms, building generalizable models of cognition that can perform naturalistic tasks, and deriving explanatory theories from these models and their simplifications.
+7.1 Looking forward: challenges & opportunities for naturalistic computational cognitive science
+While we have advocated for the importance of naturalism, there are a number of practical and conceptual challenges to achieving these goals.
+Many of these challenges provide important opportunities for future research to address.
+First, as we noted above, identifying the parameters of naturalistic variation that may interact with a particular theoretical construct of interest can be challenging.
+When we form theories—for object perception, for theory of mind, for how people plan—which “parameters” of the experimental paradigm should we make more naturalistic?
+In the trolley problem, for example, is it the number of people that must be considered?
+Is it the way in which people might die?
+Is it the length of time to decide?
+While researchers can draw on domain expertise to identify potentially-relevant parameters, there may be unexpected influences from other factors that are not typically manipulated–e.g.
+the differences between moral thoughts and moral actions when shifting from vignettes to virtual reality (Francis et al.,, 2017).
+Thus, one important direction for naturalistic research may be to identify more general overarching naturalistic factors—such as perception, embodiment, or action—that tend to influence phenomena across many domains.
+Ultimately, identifying such overarching factors may help us to identify more general unifying principles of cognition.
+Another challenge and opportunity is precisely specifying timescales of learning.
+Current learning-based theories often conflate evolution, development, and short-term learning into a single process of “optimizing” a model for the task of interest.
+This is motivated from a rational analysis perspective, where the artifact of interest is the outcome of the optimization process, which serves as a theory for behavior.
+While this has led to fruitful understanding, it leaves open the question of what exactly is learned during each stage.
+More complete “learning”-based accounts would specify precisely what an “evolutionary” data distribution would be, and the model behaviors (evolved inductive biases) that are expected from it; a succession of “developmental” data distributions and the model behaviors that would result from them; and finally predict both the behaviors that people begin tasks with and how they will with task exposure.
+Fortunately, work in this direction has already begun.
+For example, researchers have begun to explicitly map different timescales of learning to different levels of optimization in a meta-learning process (Wang,, 2021)—and indeed, some works have highlighted how meta-learning can give rise to human-like inductive biases, e.g.
+for linguistic (McCoy et al.,, 2020; McCoy and Griffiths,, 2025) or compositional (Lake and Baroni,, 2023) generalizations.
+Thus, a key opportunity for naturalistic computational cognitive science is to develop learning-based account of cognition that begin to specify how learning from different data-distributions within different segments of learning can produce the changing cognitive phenomena we observe throughout ontogeny.
+Finally, the approaches we have advocated for introduce practical challenges.
+The types of modeling and experimental paradigms we discuss here often either push towards or pull from the frontiers of contemporary computer science research.
+This can require both greater technical expertise and computational resources.
+While we acknowledge this challenge, we see it as an opportunity for larger-scale collaborative efforts—which are increasingly adopted in computer science research on training large-scale models (e.g.
+Le Scao et al.,, 2023).
+By pooling expertise and resources, collaborations make progress at larger scales than individual research groups could.
+Likewise, as discussed above, collaborations on shared benchmarks (e.g.
+Schrimpf et al.,, 2020) can allow broader and deeper assessments than any individual researcher could.
+Furthermore, the growing overlap between cognitive modeling and artificial intelligence brings opportunities for mutually-beneficial collaborations between the two fields.
+Cognitive scientists have unique expertise in developing evaluations for scientifically inferring what cognitive algorithms may underly behavioral phenomena; on the other hand AI researchers have unique expertise in training models that can perform tasks of interest in naturalistic settings.
+We hope that this paper will help to encourage deeper collaborations between these fields on their shared questions about the nature and origins of intelligence.
+7.2 Conclusions
+We have outlined a perspective that advocates for studying cognition through a broad spectrum of naturalistic experimental paradigms, building generalizable models of cognition that can perform naturalistic tasks, and deriving explanatory theories from these models and their simplifications.
 We have supported this perspective with examples illustrating the value of naturalistic settings and generalizable models, and concrete guidance on the practicalities of building generalizable models and using them to shape our theories.
@@
 Acknowledgements
-Sam Gershman for guidance and feedback on this article and on how to integrate cognitive science and machine learning methodology.
-We thank Maria Eckstein, Tyler Bonnen, Jay McClelland, Mike Mozer, Rachel Calcott, Nick Blauch, and Shuze Liu for thoughtful comments and discussions.
-We thank Felix Hill for many inspirational conversations on generalization, cognition, and more.
+We first thank Sam Gershman for his extensive guidance and feedback on this article and on how to integrate cognitive science and machine learning methodologies.
+We also thank Maria Eckstein, Tyler Bonnen, Jay McClelland, Mike Mozer, Rachel Calcott, Nick Blauch, Aran Nayebi, Ruben Cohen, Greta Tuckute, Fenil Doshi and Shuze Liu for thoughtful comments and discussions.
+We are also grateful for this work being supported by a gift from the Chan Zuckerberg Initiative Foundation to establish the Kempner Institute for the Study of Natural and Artificial Intelligence.
+Finally, we thank Felix Hill for many inspirational conversations on generalization, cognition, and more.

### v4

@@
 We first review a growing body of research spanning neuroscience, cognitive science, and AI that suggests that incorporating a broader range of naturalistic experimental paradigms, and models that accommodate them, may be necessary to resolve some aspects of natural intelligence and ensure that our theories generalize.
-First, we review cases from cognitive science and neuroscience where naturalistic paradigms elicit distinct behaviors or engage different processes.
-We then discuss recent progress in AI that shows that learning from naturalistic data yields qualitatively different patterns of behavior and generalization, and discuss how these findings impact the conclusions we draw from cognitive modeling, and can help yield new hypotheses for the roots of cognitive and neural phenomena.
+We review cases from cognitive science and neuroscience where naturalistic paradigms elicit distinct behaviors or engage different processes.
+We then discuss recent progress in AI that shows that learning from naturalistic data yields qualitatively different patterns of behavior and generalization, and examine how these findings impact the conclusions we draw from cognitive modeling, and can help yield new hypotheses for the roots of cognitive and neural phenomena.
 We then suggest that integrating recent progress in AI and cognitive science will enable us to engage with more naturalistic phenomena without giving up experimental control or the pursuit of theoretically grounded understanding.
-We offer practical guidance on how methodological practices can contribute to cumulative progress in naturalistic computational cognitive science, and illustrate a path towards building computational models that solve the real problems of natural cognition—together with a reductive understanding of the processes and principles by which they do so.
-Short Abstract
-How should cognitive science pursue generalizable theories?
-We argue that progress in Artificial Intelligence (AI) offers opportunities to embrace naturalistic experiments and computational models that can accommodate them.
-We first review literature suggesting that building generalizable theories may require a broader range of naturalistic experimental paradigms.
-We then suggest that progress in AI enables incorporating naturalistic paradigms and models without losing experimental control or theoretically-grounded understanding.
-We argue that we are entering a new era in which cognitive science can build computational models that solve the real problems of natural cognition, and derive a reductive understanding of those solutions.
-Keywords: Artificial intelligence, benchmarks, cognitive science, engineering, generalizability, naturalistic, task-performing models, theory
+We offer practical guidance on how methodological practices can contribute to cumulative progress in naturalistic computational cognitive science, and illustrate a path towards building computational models that solve the real problems of natural cognition, together with a reductive understanding of the processes and principles by which they do so.
+Keywords: Artificial intelligence, benchmarks, cognitive science, ecological validity, engineering, generalizability, naturalistic, task-performing models, theory
 1 Introduction
-Cognitive scientists build models to make our theories concrete — which offers testable predictions, eliminates ambiguities (Guest and Martin,, 2021), and can reveal unexpected properties of cognition  (McClelland,, 2009).
+Cognitive scientists build models to make our theories concrete—which offers testable predictions, eliminates ambiguities (Guest and Martin, 2021), and can reveal unexpected properties of cognition (McClelland, 2009).
 Cognitive models can range from simplified models of a high-level process (e.g.
-Frank and Goodman,, 2012; Cohen et al.,, 1990), to task-performing models of a particular domain (Newell,, 1973), or beyond (Newell,, 1994).
-Many cognitive paradigms—connectionism (McClelland et al.,, 1986), bayesian inference (Tenenbaum and Griffiths,, 2001; Tenenbaum et al.,, 2011), cognitive architectures (e.g.
-Ritter et al.,, 2019; Laird,, 2019)—have sought generalizable models that can explain a broad range of behavior from a simpler set of principles.
-However, despite progress, there is still substantial debate on the generalizability of our theories (Yarkoni,, 2022; Eckstein et al.,, 2021), whether our models and theories are adequately constrained (e.g.
-Jones and Love,, 2011; Rahnev and Denison,, 2018), and calls to accommodate a broader scope of naturalistic phenomena (Nastase et al.,, 2020; Wise et al.,, 2023; Cisek and Green,, 2024).
+Frank and Goodman, 2012; Cohen et al., 1990), to task-performing models of a particular domain (Newell, 1973), or beyond (Newell, 1994).
+Many cognitive paradigms—connectionism (McClelland et al., 1986), Bayesian inference (Tenenbaum and Griffiths, 2001; Tenenbaum et al., 2011), cognitive architectures (e.g.
+Ritter et al., 2019; Laird, 2019)—have sought generalizable models that can explain a broad range of behavior from a simpler set of principles.
+However, despite progress, there is still substantial debate on the generalizability of our theories (Yarkoni, 2022; Eckstein et al., 2021), whether our models and theories are adequately constrained (e.g.
+Jones and Love, 2011; Rahnev and Denison, 2018), and calls to accommodate a broader scope of naturalistic phenomena (Nastase et al., 2020; Wise et al., 2023; Cisek and Green, 2024).
 Our goal in this paper is to motivate and outline a path that we believe will begin to address many of these challenges.
-The path we propose is motivated by recent practical and conceptual developments in AI.
-In the last ten years, AI has progressed substantially in building models that can perform a wide range of tasks in various domains.
-We now have vision models that can classify and segment real-world images without external supervision (He et al.,, 2020; Caron et al.,, 2021), reinforcement learning models that can learn new tasks in a domain with human-like efficiency (Team et al.,, 2023), and language models that can produce human-like language and solve many language-specified tasks (Radford et al.,, 2019; Wei et al.,, 2022).
-Surprisingly, such models produce internal representations that capture many features of brain representations (Yamins et al.,, 2014; Schrimpf et al.,, 2021).
-These findings have led to discussion of how such complex models can be explanatory (Cao and Yamins, 2024a, ; Kanwisher et al.,, 2023), and how they should alter our theories (Hasson et al.,, 2020; Perconti and Plebe,, 2020; Piantadosi,, 2023).
-Yet they have also led to debate over how we test our models and interpret the results (Bowers et al.,, 2023; DiCarlo et al.,, 2023; Dentella et al.,, 2024).
-Thus, it remains an open question if and how the progress in AI can contribute to addressing the challenges of cognitive science.
+The path we propose is inspired by recent practical and conceptual developments in AI.
+We now have vision models that can classify and segment real-world images without external supervision (He et al., 2020; Caron et al., 2021), reinforcement learning models that can learn new tasks in a domain with human-like efficiency (Adaptive Agent Team et al., 2023), and language models that can solve many language-specified tasks (Radford et al., 2019; Wei et al., 2022).
+Surprisingly, such models produce internal representations that capture many features of brain representations (Yamins et al., 2014; Schrimpf et al., 2021).
+This has prompted debate on whether such complex models can be explanatory (Cao and Yamins, 2024a; Kanwisher et al., 2023), what they imply for our theories (Hasson et al., 2020; Perconti and Plebe, 2020; Piantadosi, 2023), and a broader question of how we should test and interpret our theories generally (Bowers et al., 2023; DiCarlo et al., 2023; Dentella et al., 2024).
+Thus, whether and how this progress can serve cognitive science remains open.
 Figure 1: Naturalistic computational cognitive science: the what, why, and the how.
 The first section (§2) provides an overview of “naturalistic computational cognitive science”.
-In (§34), we motivate using naturalistic experimental paradigms and learning-based approached for cognitive research.
-The remaining sections (§56) focuses on how to achieve these goals; how to develop models for naturalistic settings, and how to use naturalistic experiments and models as part of explanatory cognitive theories.
-(Figures reproduced from Saxe et al.,, 2019; Geirhos et al.,, 2018; Doshi and Konkle,, 2023.)
-In this work, we attempt to weave a thread of arguments that synthesizes these literatures, spanning from cognitive science’s motivation and theory development to the practicalities of model engineering and reproducibility.
+In (§34), we motivate using naturalistic experimental paradigms and learning-based approaches for cognitive research.
+The remaining sections (§56) focus on how to achieve these goals; how to develop models for naturalistic settings, and how to use naturalistic experiments and models as part of explanatory cognitive theories.
+(Figures reproduced from Saxe et al., 2019; Geirhos et al., 2018; Doshi and Konkle, 2023.)
+In this work, we attempt to weave a thread of arguments that synthesizes these literatures, ranging from cognitive science’s motivation and theory development to the practicalities of model engineering and reproducibility.
 We articulate a research strategy that we believe will be helpful in enabling theory-driven cognitive science to achieve a deeper understanding of the full range of natural intelligence.
-We present an overview of this framework and our arguments in Figure 1.
+We present an overview of the paper and our arguments in Figure 1.
 Our paper proceeds as follows.
-First, in §2, we unpack what we mean by “naturalistic” computational cognitive science, by providing a pragmatic outline of the features that we believe deserve more emphasis in our experimental paradigms and models.
-We argue that researchers should expand their task paradigms to more closely approximate the breadth of the settings in which their theoretical constructs should be expected to generalize—first, by adding relevant variables that may interact with the theoretical constructs in question, and second by incorporating more variability within existing parameters.
-Likewise, researchers should expand their models to more faithfully approximate the computational task that the natural system is solving.
-In the next two sections, we motivate these goals by arguing that increasing naturalism is necessary for developing generalizable understanding.
-First, in §3, we review strands of evidence spanning neuroscience, cognitive science, and AI, showing that naturalistic experimental paradigms can yield different behavior, and alter how tasks engage computational mechanisms or neural systems.
-For example, Francis et al., (2017) found that moral judgments on a trolley problem were dramatically different in a virtual reality presentation compared to text vignettes.
-Other work has found that processes ranging from visual processing (Amme et al.,, 2024), learning (Rosenberg et al.,, 2021; Collins,, 2024), memory (Helbing et al.,, 2020; Hasson et al.,, 2015) and more are changed when naturalistic elements are introduced.
-These works suggest that experiments that span a broader range of naturalistic settings can be essential for drawing correct inferences about the system as a whole.
-Next, in §4, we highlight the benefits of models that learn from naturalistic data—which can yield qualitatively different patterns of generalization than models trained in simplified settings.
-These differences can change the inferences we make from our models, and yield new insights about the origins of cognitive and neural phenomena.
-Together, these two sections of our paper argue that pursuing naturalistic experiments, and models that accommodate them, may be necessary for achieving a complete understanding of cognition.
-In the remainder of the paper, we turn to the practice of naturalistic computational cognitive science, and its contributions to building generalizable cognitive theories.
-First, in §5, we outline key lessons from AI in developing generalizable task-performing models for increasingly naturalistic data conditions: the benefits of being able to frictionlessly reproducing research artifacts, the importance of generalizability-focused model development, and the value of benchmarking for concentrating research effort on important questions.
-Finally, in §6, we outline how complicated experiments and opaque models can contribute to generalizable theories of cognition—uniting task-performing predictive models with explanatory reductive understanding, using established techniques like rational analysis (Anderson,, 1990) and new approaches to interpreting complex models (Geiger et al.,, 2021).
-We close in §7 by discussing the broader context and implications of naturalistic computational cognitive science.
-1.1 What this article is not
-We first address some potential misconceptions we anticipate about our arguments.
-First, we do not intend to argue that simple experimental paradigms are useless.
-Rather, we argue that incorporating more naturalistic settings alongside simpler ones is essential to developing more generalizable cognitive theories.
-We also do not intend to promote a particular class of models, aside from our emphasis on learning.
-We believe that our arguments are compatible with many learning-based cognitive models, e.g.
-Bayesian cognitive models (Lee and Wagenmakers,, 2013), reinforcement learning (Collins,, 2019), predictive representations (Carvalho et al.,, 2024), program synthesis models (Rule et al.,, 2024), etc.
-Finally, we do not advocate for black-box foundation models trained directly to reproduce experimental data as the goal of naturalistic computational cognitive science.
-While this can be a useful tool (e.g.
-for experimental design), we advocate for theory-driven development of predictive models that capture human behavior as an important tool in the quest to develop and validate generalizable cognitive theories.
-In summary, this paper is not intended to reject the current experimental, modeling, and theoretical approaches of cognitive science.
-Rather, we advocate for extending from simpler experiments to more naturalistic ones.
-We then advocate for building task-performing models that can predict human behavior across these paradigms, and linking these models to simpler reductive models and theories of that behavior.
-We argue that this approach will enable building more generalizable theories that explicitly link cognitive capacities to the full scope of natural behavior to which those capacities contribute.
-Figure 2: Overview of naturalistic computational cognitive science.
-Given a novel computational model, researchers generate predictions to test their model in a simplified setting.
-If this model successfully predicts human behavior, researchers see if their model continues to generalize human predictions in increasingly naturalisic conditions.
-Figure 3: Examples of increasingly naturalistic settings that we can now study in a theory-driven manner.
-All of these are settings where tasks and stimuli can now be parametrically generated—i.e., thanks to “generative AI”, we can now automate the generation of photorealistic synthetic data and virtual worlds; thanks to virtual and augmented reality, we can now scan and parametrically manipulate real environments.
-Thanks to progress in AI, we can now build task-performing models that can operate in these stimuli and task settings.
+In §2, we specify what naturalistic computational cognitive science means in practice.
+We argue that researchers should broaden their paradigms beyond the settings in which their theoretical constructs are typically tested—by both adding variables that interact with those constructs and by sampling more broadly within existing parameters.
+Likewise, we should build models that more faithfully approximate the computational task that the natural system is solving.
+The next two sections address why.
+We argue that increasing naturalism is necessary—not merely beneficial—for developing a generalizable theory of cognition.
+In §3, we review evidence from neuroscience, cognitive science, and AI showing that naturalistic paradigms yield qualitatively different behavior and engage different computational mechanisms—processes spanning visual processing (Amme et al., 2024), learning (Rosenberg et al., 2021; Collins, 2024), memory (Helbing et al., 2020; Hasson et al., 2015), and moral reasoning (Francis et al., 2017) all shift when naturalistic elements are introduced.
+These findings suggest that naturalistic experiments are essential for building a complete theory of the system.
+In §4, we go further: training on naturalistic data does not merely reveal different mechanisms but can also shape them, producing qualitatively different generalization patterns than training in simplified settings.
+These differences can alter the inferences we draw from our models and shed new light on the origins of cognitive and neural phenomena.
+Together, these two sections establish that pursuing naturalistic experiments—and models that accommodate them—is necessary for a complete understanding of cognition (Figure 2).
+The remaining sections address the how.
+In §5, we distill a key lesson from AI for developing task-performing models in naturalistic settings: building generalizable models requires iterative and collective effort.
+We propose dynamic meta-benchmarks—living collections that researchers continually expand—as the coordination mechanism, and outline a development strategy for fast, frictionless iteration.
+In §6, we show how naturalistic experiments and complex models can anchor generalizable cognitive theories—uniting task-performing predictive models with reductive understanding of underlying mechanisms, using tools such as rational analysis (Anderson, 1990) and new approaches to interpreting complex models (Geiger et al., 2021).
+We close in §7 with the broader context and implications of naturalistic computational cognitive science.
 2 What is “naturalistic” computational cognitive science?
+Figure 2: Two examples of naturalistic computational cognitive science.
+Given a computational theory, researchers generate initial predictions to test their theory in a simplified setting.
+Theories are validated by their ability to predict human behavior in increasingly naturalistic conditions, and refined where they fail to predict correctly.
+(Top) Progression of theories of sequential task generalization.
+(Bottom) Examples of progression in understanding PRC involvement in difficult visual judgments..
+Figures adapted from Momennejad et al.
+(2017); Tomov et al.
+(2021); Carvalho et al.
+(2025); Bonnen et al.
+(2021, 2026).
 Naturalistic computational cognitive science is a research strategy for theory-driven cognitive science that aims to predict and understand human behavior across increasingly naturalistic settings.
 This approach gradually increases the ecological validity of experimental designs while ensuring models can accommodate both simplified and more natural conditions.
+Of course, our emphasis on naturalism does not imply that simple experimental paradigms are useless—they remain valuable, and the goal is to extend them with naturalistic settings rather than replace them.
 In machine learning (§5), analogous strategies have enabled models that can work with real-world data, albeit imperfectly.
+Aside from our focus on learning-based methods, this approach is model-class agnostic and compatible with many learning-based frameworks, including Bayesian models (Lee and Wagenmakers, 2013), RL (Collins, 2019), predictive representations (Carvalho et al., 2024), and program synthesis (Rule et al., 2024).
+The main criterion we advocate for is that models are theory-driven—not black-box foundation models trained to reproduce experimental data (Binz et al., 2024), though these may be a useful tool for experimental design.
 As in other approaches within computational cognitive science, this approach starts with formulating models and designing controlled experiments to isolate core phenomena.
 Like prior work, we place an emphasis on model generalization, i.e.
-predicting human behavior on data the model hasn’t been exposed to—including data from new experimental conditions or tasks (Busemeyer and Wang,, 2000).
+predicting human behavior on data the model hasn’t been exposed to—including data from new experimental conditions or tasks (Busemeyer and Wang, 2000).
 Therefore,
@@
     Models should predict human behavior on held-out examples that neither the model nor humans have encountered.
-One key facet of naturalistic computational cognitive science is that we place emphasis on having models—in particular, the same model—predict human behavior across increasingly naturalistic conditions that cover a broader space of settings in which the theoretical construct would be expected to generalize.
-We see two main paths toward increasing naturalism in experimental paradigms:
+One key facet of naturalistic computational cognitive science is that we place emphasis on having models—in particular, the same model—predict human behavior across increasingly naturalistic conditions that cover the space of settings in which the theoretical construct is expected to generalize.
+We see two main paths toward increasing naturalism in experimental paradigms (and provide examples in Figure 3):
   1. 1.
@@
 For example, in a linguistic judgment task, we might test judgments across both spoken and written utterances.
-We provides examples in Figure 3.
 We describe a paradigm as more naturalistic if it incorporates a broader scope of the range of natural variability over which the theoretical constructs would be expected to generalize.
 We describe a model as more naturalistic if it is capable of generalizing its predictions across a broader range of contexts.
-Note that increasing naturalism does not always mean faithfully approximating all the joint statistics of the natural distribution—increasing the scope of variation along naturalistic dimensions can also enable moving outside the natural data distribution for controlled experiments (see §6.1).
+Note that increasing naturalism does not always mean faithfully approximating all the joint statistics of the natural distribution—increasing the scope of variation along naturalistic dimensions can also enable moving outside the natural data distribution for controlled experiments testing a hypothesis (see §6.1).
 Generally, identifying the parameters for increasing naturalism in a particular experiment is not trivial—it is task- and theory-specific.
 To illustrate “naturalistic” more explicitly, we therefore offer some examples of dimensions along which naturalism can be increased.
+Figure 3: Examples of increasingly naturalistic settings that we can now study in a theory-driven manner.
+All of these are settings where tasks and stimuli can now be parametrically generated—i.e., thanks to “generative AI”, we can now automate the generation of photorealistic synthetic data and virtual worlds; thanks to virtual and augmented reality, we can now scan and parametrically manipulate real environments.
+Thanks to progress in AI, we can now build task-performing models that can operate in these stimuli and task settings.
 Task paradigm
@@
 Environment
-  • Large state spaces that reflect the complexity of real-world environments (Wise et al.,, 2023).
-  • Incorporating other social agents that learn with the agent (Hu et al.,, 1998) and that the organism must interact with (Vélez and Gweon,, 2021).
-  • A continually changing environment (Abel et al.,, 2024).
+  • Large state spaces that reflect the complexity of real-world environments (Wise et al., 2023).
+  • Incorporating other social agents that learn with the agent (Hu et al., 1998) and that the organism must interact with (Vélez and Gweon, 2021).
+  • A continually changing environment (Abel et al., 2024).
 Model architecture
-  • Architectures that operate over sensory inputs like natural images (Yamins et al.,, 2014), speech (Kell et al.,, 2018), or natural language (Schrimpf et al.,, 2021) rather than simplified stimuli (e.g.
+  • Architectures that operate over sensory inputs like natural images (Yamins et al., 2014), speech (Kell et al., 2018), or natural language (Schrimpf et al., 2021) rather than simplified stimuli (e.g.
 low-dimensional or discrete inputs containing only task-relevant features).
-  • Architectures with naturalistic action spaces—for example, modeling motor control and embodiment rather than abstract, symbolic actions (Merel et al.,, 2019).
+  • Architectures with naturalistic action spaces—for example, modeling motor control and embodiment rather than abstract, symbolic actions (Merel et al., 2019).
 Learning algorithm
-  • Unsupervised (Higgins et al.,, 2021), self-supervised (Konkle and Alvarez,, 2022), or intrinsic (Chentanez et al.,, 2004; Oudeyer et al.,, 2013) learning algorithms.
-  • Social (Henrich,, 2016) or cultural (Cook et al.,, 2024) learning algorithms.
-  • Prospective algorithms that aim to model how the tasks evolve in continually changing environments (Seligman et al.,, 2013; De Silva et al.,, 2023).
-In summary, naturalistic computational cognitive science seeks to (1) explain real-world intelligent behavior by developing models that operate over naturalistic inputs, produce naturalistic outputs, and are optimized according to the actual constraints and affordances of the environment, and (2) develop cognitive theories that unite these models with reductive understanding.
-In the coming sections, we will first motivate how naturalistic experimental settings can help us develop more complete accounts of intelligence, before returning in more detail to how these goals can be achieved.
-3 The benefits of naturalistic experimental settings
-Figure 4: Overview of the benefits of increasingly naturalisic experimental conditions.
-In this section, we outline the benefits of naturalistic experimental settings for cognitive science.
-Specifically, we argue that naturalistic experimental settings can engage mechanisms that are qualitatively different than those engaged in simpler settings — and thereby alter the scientific inferences that we draw.
-We review examples where naturalistic experimental settings have changed our understanding of a system, or have helped to disentangle competing models that simpler settings could not.
-As an orienting conceptual example, imagine that we are doctors studying heart function, and we chose to only study it while the participants are lying down and resting.
-We could certainly learn interesting things thereby; the resting setting offers high test-retest reliability, and a highly-controlled environment to study processes like how respiration affects heart rate.
-However, this control and reliability comes precisely from removing the many factors of natural variation that affect the heart, and indeed for which heart function evolved — such as rapid adaptation to movement, stress, etc.
-Studying heart function across a broader range of naturalistic settings would be necessary to understand the system in full, and how the physiological (and psychological) processes involved interact.
-For example, if we experimentally manipulated breathing across a range of different naturalistic activities (sitting, speaking, running, etc.) we would more effectively determine its effect on heart function, and how that effect varies across contexts.
-By studying phenomena across varied naturalistic settings we elaborate our understanding of the system, and can build more complete theories of its function.
-3.1 Behavior can be different in naturalistic experiments
-Many cognitive analyses begin by studying a task designed to isolate some cognitive capability.
-However, in many cases task-relevant behavior can be altered by seemingly-orthogonal features of the task—thus raising the risk that our theories will be overfit to a restricted setting unless we also consider a broader range of variation.
-In this section, we review some examples illustrating changes in behavior or capabilities in naturalistic settings.
+  • Unsupervised (Higgins et al., 2021), self-supervised (Konkle and Alvarez, 2022), or intrinsic (Chentanez et al., 2004; Oudeyer et al., 2013) learning algorithms.
+  • Social (Henrich, 2016) or cultural (Cook et al., 2024) learning algorithms.
+  • Prospective algorithms that aim to model how the tasks evolve in continually changing environments (Seligman et al., 2013; De Silva et al., 2023).
+These four dimensions are not a definition of naturalism.
+They are examples of directions along which a theory of cognition must hold up to count as general.
+Generalizability means extending along axes like these—and often theories are not extended along any.
+3 The importance of naturalistic experimental settings for building generalizable theories
+Figure 4: Overview of the benefits of increasingly naturalistic experimental conditions.
+As an orienting conceptual example, let us imagine that we are doctors studying heart function, and have chosen to study participants lying at rest.
+The setting offers high test-retest reliability and a controlled view of processes like how respiration affects heart rate.
+But that control comes precisely from removing the variation the heart evolved to handle — rapid adaptation to movement, stress, and the like.
+Understanding the system in full, and how its physiological and psychological processes interact, requires studying the heart across more naturalistic settings.
+For instance, manipulating breathing across activities like sitting, speaking, and running would pin down its effect on heart function more sharply than rest alone, and reveal how that effect varies with context.
+Varied naturalistic settings are how we elaborate the system and build complete theories of its function.
+3.1 Naturalistic experimental paradigms can expose computational challenges that engage mechanisms differently
+Here, we highlight examples where the challenges posed by increasing naturalism can engage computational mechanisms differently—which can help to disentangle underlying mechanisms, and achieve more generalizable theoretical understanding.
+Example 1: mechanisms for credit assignment are only exposed when actions have delayed consequences.
+In naturalistic settings, people learn how actions lead to consequences hundreds of timesteps away.
+Consider navigating a university campus: you learn that heading north leads to the music school, then discover a jazz cafe just off that path.
+Credit assignment (Sutton and Barto, 2018) is what lets you recognize the northward route serves both.
+Prior work modeled this with successor features (Gershman, 2018; Tomov et al., 2021), but only studied human behavior in worlds of around 13 states—small enough that models can succeed without any explicit long-horizon mechanism.
+In larger worlds, successor features predict fast reaction times but cannot handle nearby-object transfer.
+In contrast, model-based planning handles this transfer but predicts slow deliberation.
+Humans display both.
+To explain this, Carvalho et al.
+(2025) proposed multitask preplay: people mentally rehearse potential detours in advance and cache the results for fast retrieval later on (Figure 2, top).
+Testing multitask preplay required a large environment: in smaller worlds, planning is cheap enough that cached retrieval looks no different from deliberation.
+Implementing cached retrieval with a brain-plausible learning rule then made the credit-assignment problem concrete—the agent must credit a northward choice (taken to reach the music school) for the cafe discovered along the way.
+Such cross-task transfer is a known failure mode of temporal-difference learning: approximation errors inflate the value of actions that should not be taken.
+Multitask preplay required additional machinery to mitigate this, pointing to mechanisms the brain may use for the same job.
+Because the model was developed in a naturalistic setting, it continued to predict human behavior in an even more naturalistic one—a large, partially observable 2D world with many objects.
+Example 2: seemingly-similar tasks that engage brain mechanisms differently can be reconciled by naturalistic models.
+A long-running debate asks whether perceptual “oddity” tasks (identifying the object that differs from the others in a set) recruit the Medial Temporal Lobe (MTL).
+Murray et al.
+(2007) and others argue that MTL is recruited to discriminate “complex” stimulus sets that visual cortex cannot resolve alone, so MTL lesions selectively impair performance on those stimuli; others find no such interaction and attribute it to methodological artifacts (Buffalo et al., 1998; Suzuki, 2009).
+Bonnen et al.
+(2021); Bonnen and Eldridge (2023) resolved the impasse by passing all the raw stimuli from prior experiments through pretrained convolutional vision models and testing whether the oddity could be picked out in representation space.
+When the vision model alone could distinguish the oddity, MTL was dispensable; when it could not, MTL was required.
+Some “complex” stimulus sets from prior experiments fell on each side of this split.
+“Complexity” had been underspecified: differences in operationalization that conceptual descriptions abstracted away drove the seemingly-conflicting findings.
+This example illustrates the value of building models that directly perform the same naturalistic tasks as the subjects, and testing these models on benchmarks that incorporate stimuli from many experimental paradigms—resolving otherwise intractable theoretical disagreements.
+Example 3: Different learning mechanisms only yield distinct predictions under working memory load.
+One challenge of overly-simplified tasks is potential aliasing of different solutions, where many different computational approaches yield the same behavior.
+Collins (2024) presents one example where a behavior ostensibly implemented as a standard reinforcement learning algorithm may instead be produced from simpler value-free associative learning combined with working memory.
+[^1] Whether this alternative counts as an alternative implementation of RL is orthogonal to our argument; our point is that the models yield distinct predictions in only some experimental regimes.
+Collins notes that “Disentangling multiple processes requires considering more complex tasks to elicit differentiable behavior.” The more complex tasks in question simply increase the set of stimuli within learning blocks to a handful of objects, rather than only two or three—i.e., expanding the range of set sizes more clearly disentangles the learning processes.
+Thus, by restricting to minimal paradigms, we enable a system to use many solutions and prevent ourselves from discriminating between them.
+By contrast, by imposing a broad range of evaluations on the system (cf.
+Nau et al., 2024)—especially naturalistic evaluations that increase demands along different axes of task difficulty—the increased constraints actually make it easier to map the model on to natural intelligence.
+More constraints on the mapping make it less ambiguous, as is highlighted in the contravariance principle of Cao and Yamins (2024b).
+3.2 Behavior can be qualitatively different in naturalistic experiments
+Cognitive analyses typically isolate a capacity with a task designed to elicit it.
+But behavior often shifts with seemingly-orthogonal aspects of that task, and theories tuned to one setting risk overfitting to it.
+Three cases illustrate the risk concretely—putatively well-isolated capacities (memory, learning, moral choice) behave qualitatively differently under naturalistic demands.
+Example 1: memory performance differs between naturalistic search and explicit memorization.
+When participants are explicitly instructed to memorize objects in a 3D home environment, their subsequent recall accuracy was significantly lower than when they incidentally encountered the same objects during visual search tasks (Helbing et al., 2020).
+This effect also extends to spatial memory—participants were able to place objects closer to their original positions in the search condition.
+These findings demonstrate that tasks focused explicitly on memory may not capture all aspects of how memory operates during naturalistic behavior.
+Thus, computational models of memory may need to be revised to account for these differences in naturalistic settings.
+Example 2: mice can learn $1000\times$ faster during natural behavior, compared to two-alternative forced choice (2AFC) tasks.
+Recent research has shown that learning efficiency can dramatically differ between simplified laboratory tasks and more naturalistic settings (Rosenberg et al., 2021).
+When mice are tested in traditional 2AFC tasks, they typically require 10,000 training trials over 3-6 weeks to reach asymptotic performance of only about 67% accuracy.
+However, when allowed to freely explore a labyrinth (similar to natural burrow systems (Small, 1901)), mice demonstrate remarkably accelerated learning (Rosenberg et al., 2021); they learned to navigate a complex maze with 63 T-junctions in just a few hours, optimizing their path to a reward after only about 10 reward experiences.
+The time it took them to learn in the naturalistic setting was approximately $1000\times$ faster than in 2AFC paradigms.
+While it may not be meaningful to directly compare “learning rate” as a parameter between the settings, this clear and dramatic difference in learning efficiency demonstrates how traditional simplified experimental paradigms may fundamentally underestimate an organism’s learning capabilities, and suggests that learning mechanisms may operate differently when they are engaged through more naturalistic behavior.
 Example 3: moral behavior can change when moving from textual vignettes to virtual reality.
 Recent research shows that behavioral choices can change when moving from text to virtual reality, by driving a shift from moral judgment (i.e.
-deciding whether things are moral) to moral action (Francis et al.,, 2017, see Figure 4 A).
+deciding whether things are moral) to moral action (Francis et al., 2017, see Figure 4 A).
 When participants were presented the classic trolley dilemma in text form, only 40% said they would pull a switch to divert the trolley, thus killing one person to save five others.
@@
 These findings demonstrate that people’s stated moral preferences in hypothetical scenarios can differ substantially from their actions in more naturalistic settings.
-Example 2: memory performance differs between naturalistic search and explicit memorization.
-When participants are explicitly instructed to memorize objects in a 3D home environment, their subsequent recall accuracy was significantly lower than when they incidentally encounter the same objects during visual search tasks (Helbing et al.,, 2020).
-This effect also extends to spatial memory—participants were able to place objects closer to their original positions in the search condition.
-These findings demonstrate that tasks focused explicitly on memory may not capture all aspects of how memory operates during naturalistic behavior.
-Thus, computational models of memory may need to be revised to account for these differences in naturalistic settings.
-Example 3: mice can learn $1000\times$1000 × faster during natural behavior, compared to two-alternative forced choice (2AFC) tasks.
-Recent research has shown that learning rates can dramatically differ between simplified laboratory tasks and more naturalistic settings (Rosenberg et al.,, 2021).
-When mice are tested in traditional 2AFC tasks, they typically require 10,000 training trials over 3-6 weeks to reach asymptotic performance of only about 67% accuracy.
-However, when allowed to freely explore a labyrinth (similar to natural burrow systems (Small,, 1901)), mice demonstrate remarkably accelerated learning (Rosenberg et al.,, 2021).
-In this setting, mice learned to navigate a complex maze with 63 T-junctions in just a few hours, optimizing their path to a reward after only about 10 reward experiences.
-The learning rate in the naturalistic setting was approximately $1000\times$1000 × times faster than in 2AFC paradigms.
-This dramatic difference in learning speed demonstrates how traditional simplified experimental paradigms may fundamentally underestimate an organism’s learning capabilities, and suggests that learning mechanisms may operate differently when they are engaged through more naturalistic behavior.
 3.3 Neural systems can operate differently under naturalistic conditions
@@
 integrating sensory inputs for body control.
-Cisek and Green, (2024) argues that as we increased naturalism from head-fixed to head-free and body-free settings for studying monkeys, we expanded our theory of superior colliculus from controlling saccadic eye movement to generally integrating multimodal cues to guide bodily action-selection (Figure 4 B).
+Cisek and Green (2024) argues that as we increased naturalism from head-fixed to head-free and body-free settings for studying monkeys, we expanded our theory of superior colliculus from controlling saccadic eye movement to generally integrating multimodal cues to guide bodily action-selection (Figure 4 B).
 That is, as we increased the naturalism of the experimental conditions, we arrived at a more complete model of superior colliculus.
-This illustrates the more general point, that overly focusing on computation as an input-output mapping neglects the fundamental fact that natural intelligence evolved not for single responses, but for closed-loop control in an environment (Cisek,, 1999).
+This illustrates the more general point, that overly focusing on computation as an input-output mapping neglects the fundamental fact that natural intelligence evolved not for single responses, but for closed-loop control in an environment (Cisek, 1999).
 This perspectives yields a different interpretation of the system’s representations and processes.
-More generally, many researchers argue that cognition cannot be understood completely outside its embodiment and environment (e.g., Newen et al.,, 2018).
+More generally, many researchers argue that cognition cannot be understood completely outside its embodiment and environment (e.g., Newen et al., 2018).
 Example 2: visual processing differs during natural viewing compared to passive viewing paradigms.
-One early visual response in the brain (the P100/M100) occurs about 100ms after a person’s eyes fixate on a new location (Vinje and Gallant,, 2000).
+One early visual response in the brain (the P100/M100) occurs about 100ms after a person’s eyes fixate on a new location (Vinje and Gallant, 2000).
 However, this finding comes primarily from experiments where participants passively view images while fixating.
 In contrast, natural vision involves frequent active eye movements to sample information throughout a scene.
-When Amme et al., (2024) examined brain responses during more natural active viewing, they found that the P100/M100 response actually begins when the eye movement (saccade) begins, not when it ends at a new fixation point.
+When Amme et al.
+(2024) examined brain responses during more natural active viewing, they found that the P100/M100 response actually begins when the eye movement (saccade) begins, not when it ends at a new fixation point.
 Furthermore, the neural patterns during active viewing were anti-correlated to those seen during passive viewing, suggesting fundamentally different processing mechanisms.
@@
 Example 3: model-based systems are used for learning moral judgments, but model-free systems are used for learning to avoid harming others.
-Despite moral reasoning being strongly associated with model-based learning in the prefrontal cortex, neural evidence shows a surprising shift when people actually learn to avoid harming others (Lockwood et al.,, 2020).
-Moral reasoning commonly engages model-based systems in the lateral prefrontal cortex (LFPC) (Spitzer et al.,, 2007; Carlson and Crockett,, 2018) with LPFC disruption leading to reduced norm compliance and enforcement (Knoch et al.,, 2006; Ruff et al.,, 2013).
+Despite moral reasoning being strongly associated with model-based learning in the prefrontal cortex, neural evidence shows a surprising shift when people actually learn to avoid harming others (Lockwood et al., 2020).
+Moral reasoning commonly engages model-based systems in the lateral prefrontal cortex (LPFC) (Spitzer et al., 2007; Carlson and Crockett, 2018) with LPFC disruption leading to reduced norm compliance and enforcement (Knoch et al., 2006; Ruff et al., 2013).
 However, when participants perform tasks requiring them to learn moral behavior (rather than just reason about it), researchers instead observed neural activity consistent with model-free learning.
-As in §3.1, this illustrates that increasing naturalism of experiments (learning to take moral actions rather than simply reasoning about them), neural systems can operate in unexpected ways.
-3.3 Naturalistic experimental paradigms can expose computational challenges that engage mechanisms differently
-Here, we highlight examples where the challenges posed by increasing naturalism can engage computational mechanisms differently—which can help to disentangle underlying mechanisms, and achieve more generalizable understanding.
-Example 1: Different learning mechanisms only yield distinct predictions under working memory load.
-One challenge of overly-simplified tasks is potential aliasing of different solutions, where many different computational approaches yield the same behavior.
-Collins, (2024) presents one example where a behavior ostensibly implemented as a standard reinforcement learning algorithm may instead be produced from simpler value-free associative learning combined with working memory.
-[^1] Whether this alternative counts as an alternative implementation of RL is orthogonal to our argument; our point is that the models yield distinct predictions in only some experimental regimes.
-They note that “Disentangling multiple processes requires considering more complex tasks to elicit differentiable behavior.” The more complex tasks in question simply increase the set of stimuli within learning blocks to a handful of objects, rather than only two or three—i.e., expanding the range of set sizes more clearly disentangles the learning processes.
-Thus, by restricting to minimal paradigms, we enable a system to use many solutions and prevent ourselves from discriminating between them.
-By contrast, by imposing a broad range of evaluations on the system (cf.
-Nau et al.,, 2024)—especially naturalistic evaluations that increase demands along different axes of task difficulty—the increased constraints actually make it easier to map the model on to natural intelligence (more constraints on the mapping make it less ambiguous), as is highlighted in the contravariance principle of Cao and Yamins, 2024b .
-Example 2: fear conditioning.
-Traditional fear conditioning research offers fundamental insights into how humans learn associations between stimuli and aversive outcomes (Sehlmeyer et al.,, 2009; Maren,, 2001; Britton et al.,, 2014).
-Through carefully controlled laboratory studies, we have developed a rich mechanistic understanding of basic fear learning processes.
-These paradigms have proven especially valuable for studying certain types of naturally occurring fear learning, such as immediate aversive responses to foods (Riley and Tuck,, 1985) or traumatic events (Engelhard et al.,, 2008).
-However, real-world fear learning often involves additional temporal dynamics that complement these well-studied processes (Figure 4 C).
-For example, a student may develop a fear of exams after receiving a single poor grade weeks after taking a test, without requiring repeated negative experiences (Mobbs et al.,, 2021).
-This naturalistic learning involves delayed and sparse feedback.
-Thus, it relies on a computational mechanism for credit assignment (Sutton and Barto,, 2018)—a mechanism for crediting choices (e.g.
-studying or sleep) as responsible for future events (e.g.
-a poor grade).
-Naturalistic settings enable studying important functions such as mechanisms for credit assignment.
-Increased ecological validity can thereby increase the applicability of our findings; indeed, the gap between lab-based fear conditioning and real-world fear has been a barrier to treating pathological fears (Beckers et al.,, 2013; Scheveneels et al.,, 2016; Krypotos et al.,, 2018).
-Example 3: seemingly-similar tasks that engage brain mechanisms differently can be reconciled by naturalistic models.
-There has been a long debate about whether perceptual processes like “oddity” tasks (identifying an object that differs from the others in a set) recruit the Medial Temporal Lobe (MTL).
-Some researchers (e.g Murray et al.,, 2007) suggest that MTL is recruited to discriminate among “complex” stimulus sets that visual cortex cannot discriminate on its own (thus MTL lesions selectively impair performance on those stimuli); other researchers have found no such interaction and suggested it stems from methodological issues (Buffalo et al.,, 1998; Suzuki,, 2009).
-Recently, stimulus-computable models have shed new light on this debate.
-Specifically, Bonnen et al., (2021); Bonnen and Eldridge, (2023) used pretrained convolutional vision models to directly process all the raw stimuli used in prior experiments, and evaluated whether the “oddity” stimulus could be distinguished from the rest of the set in the representation space of these models.
-By doing so, Bonnen et al.
-showed that purportedly “complex” stimuli that could be discriminated by the pretrained vision network alone did not require the MTL, while the complex stimuli that could not be distinguished by the vision model were precisely the ones that required MTL.
-This shows that the way that prior works had operationalized “complexity” was underspecified; differences in operationalization that got abstracted away in conceptual descriptions of the experiments likely drove the seemingly-conflicting findings.
-This example illustrates the value of building models that directly perform the same naturalistic tasks as the subjects, and testing these models on benchmarks that incorporate stimuli from many experimental paradigms.
-This naturalistic approach can elucidate important differences among experimental paradigms, and clarify conceptual understanding.
+As in §3.2, this illustrates that when we increase the naturalism of experiments (learning to take moral actions rather than simply reasoning about them), neural systems can operate in unexpected ways.
 4 The (surprising) benefits of learning with naturalistic data
-Figure 5: Overview of the benefits of learning with naturalistic data. (Figures reproduced from Doshi and Konkle,, 2023; Manning et al.,, 2020; Team et al.,, 2023.)
-A standard lesson in science is that we must simplify an experimental setting to better arrive at a causal conclusion.
-However, we have reviewed examples illustrating that more is different (Anderson,, 1972)—some properties of intelligence may only emerge in more “complicated” naturalistic settings.
+Figure 5: Overview of the benefits of learning with naturalistic data. (Figures reproduced from Doshi and Konkle, 2023; Manning et al., 2020; Adaptive Agent Team et al., 2023.)
+A standard lesson in science is that we must simplify an experimental setting to better arrive at a causal conclusion (e.g.
+Rust and Movshon, 2005).
+However, we have reviewed examples illustrating that more is different (Anderson, 1972)—some properties of intelligence may only emerge in more “complicated” naturalistic settings.
 Here, we detail examples where naturalistic data itself plays an important role in producing the empirical phenomena of intelligence.
@@
 While we do not yet know the full implications for natural intelligence, these results open interesting questions and challenge prior assumptions.
-In particular, cognitive science should consider the limitations placed on our models when we fail to consider naturalistic data.
-(While we draw case studies from deep learning, we believe that similar factors could likewise affect generalization under other learning paradigms, such as program synthesis, or even evolution—and thus should be of interest even to those who do not work with deep learning models.)
-4.1 Learning enables models that can operate on natural data
-One surprising benefit of learning is that it can enable a system to operate over natural data.
-Until AlexNet (Krizhevsky et al.,, 2012), the prevailing wisdom was that one had to design useful representations of natural data for models.
-However, AlexNet and subsequent work showed that a model could learn useful representations for naturalistic data simply through learning to make predictions about those data.
-Surprisingly, these models learn representations that have notable correspondence to the internal representations in visual cortex (Yamins et al.,, 2014; Khaligh-Razavi and Kriegeskorte,, 2014).
-Here, we review some of this work and how learning-based methods can enable more flexible cognitive models that operate on natural data.
-Learning-based vision models can operate on natural data and predict neural data (surprisingly) well.
-A rich history of work reveals how neural networks can learn internal representations from natural images that strongly correspond to neural data.
-A seminal example comes from Olshausen and Field, (1996), who showed that V1-like Gabor filters could emerge from a convolutional network optimized to transmit input while minimizing overall activation (essentially, a sparse autoencoder).
-This breakthrough demonstrated that V1-like receptive fields could arise from optimization for natural environmental statistics.
-Subsequently, Lee et al., (2007) demonstrated that V2-like convolutional filters could emerge from unsupervised optimization of sparse deep belief networks.
-Later, Yamins and DiCarlo, (2016) showed that deep convolutional networks trained for object classification developed intermediate layers that matched neural representations in V4 and IT better than any previous models.
-Strikingly, task-optimized networks trained to self-organize represent images in a latent space whose geometry corresponds closely to brain topography (Doshi and Konkle,, 2023, Figure 5 A).
-Thus, learning-based approaches not only enable effective visual processing but also can provide surprisingly accurate predictions of brain organization and function.
-Integrating learning-based models with program synthesis enables flexible Bayesian models of human concept learning.
-Recently, Ellis, (2024) showed how language models can be combined with program synthesis to create more generalizable Bayesian models of human reasoning.
-Language models enable using natural language as a flexible hypothesis space with a prior derived from human data.
-By translating natural language hypotheses to executable programs for likelihood computation, such models can capture human learning dynamics in online concept acquisition tasks while continuing to predict human behavior robustly in new domains.
-This overcomes traditional limitations of hand-designed symbolic hypothesis spaces and pure program synthesis approaches, which can fail to transfer due to incomplete symbolic vocabularies—a manifestation of the frame problem (Shanahan,, 2004) in hypothesis space design.Integrating learning-based components with program synthesis thus enables Bayesian models that can flexibly adapt their hypothesis space while maintaining interpretable symbolic computation—a goal of many cognitive modelers.
+In particular, cognitive science should consider the limitations placed on our models—and correspondingly, on the theories we instantiate in those models—when we fail to consider naturalistic data.
 4.1 Learning from naturalistic data can improve generalization
-Studying how systems generalize is fundamental to cognitive science and AI.
-When studying generalization in simplified settings, a key assumption is that the simplifications do not alter the generalization problem.
-Here, we review studies from AI and neuroscience showing that variability in experience (even aloong seemingly-orthogonal axes) can alter learning and generalization—thus implying that studying models within more restricted settings may mislead us about the naturalistic computational problem.
+Studying how systems generalize is fundamental to building theories in cognitive science and AI.
+In order to study generalization in simplified settings, we need to make the key assumption that the simplifications do not alter the generalization problem.
+Here, we review studies from AI and neuroscience showing that variability in experience (even along seemingly-orthogonal axes) can alter learning and generalization—thus implying that studying models within more restricted settings may mislead us about the naturalistic computational problem.
 Example 1: Compositional generalization of image classifiers vs.
@@
 Naturalistic tasks can fundamentally change what models learn and how they generalize.
-Hill et al., (2020) trained two models on a vision-language grounding task, and tested their compositional generalization to held-out instructions.
-Both models were trained on the same language examples, and tested on the same held-out examples.
-However, one model was trained as an agent interacting with a simulated environment, while the other was simply trained trained to classify screenshots from that environment based on language instructions.
-The authors found that the interacting agent exhibited perfect compositional generalization to novel language utterances, while the image classifier generalized substantially worse.
-The authors also explored other factors, including generalization benefits of more naturalistic 3D (rather than 2D) environments, or egocentric embodiments.
-Their results illustrate how richer naturalistic settings can enhance the generalizability of learned solution.
+Hill et al.
+(2020) trained two models on a vision-language grounding task, and tested their compositional generalization to held-out instructions.
+Both models were trained on the same language set, and tested on the same held-out examples.
+However, one model was trained as an agent interacting with a simulated environment, while the other was simply trained to classify screenshots from that environment based on language instructions.
+Surprisingly, the interacting agent exhibited perfect compositional generalization to novel language utterances, while the image classifier generalized substantially worse.
+The authors also found generalization benefits of other naturalistic factors, including 3D (rather than 2D) environments, or egocentric embodiments.
+These results illustrate how richer naturalistic settings can enhance the generalizability of learned solutions.
 Thus, if we are interested in understanding generalization, we may need to build models that learn from appropriately naturalistic data.
 Example 2: Syntactic generalization is enhanced by variability in other structures’ fillers Structural generalization is of deep interest in linguistics.
-Recently, Misra and Mahowald, (2024) performed controlled experiments in which challenging linguistic constructions are systematically held out from language model training data, and showed that models trained on naturalistic language generalize to held out constructions by composing pieces of simpler constructions.
+Recently, Misra and Mahowald (2024) performed controlled experiments in which challenging linguistic constructions are systematically held out from language model training data, and showed that models trained on naturalistic language generalize to held out constructions by composing pieces of simpler constructions.
 Critically, this generalization depended on sufficient variability in the semantic fillers observed in the structures seen in training—thus, incorporating naturalistic variation in model training data can impact our theoretical inferences about generalization.
@@
 Analogously, neuroscientists have found that animals raised in more complex environments can be more skilled than those raised in simpler environments.
-Rats raised in enriched environments — with social interaction and/or more space or toys — show greater exploration, and improved learning and memory on novel tasks (Simpson and Kelly,, 2011).
+Rats raised in enriched environments — with social interaction and/or more space or toys — show greater exploration, and improved learning and memory on novel tasks (Simpson and Kelly, 2011).
 Thus, for natural as well as artificial intelligence, greater naturalistic variation in experience can alter the system’s learning in ways that impact our experimental and theoretical inferences.
@@
 One early success of deep learning came in computer vision.
-Soon after AlexNet (Krizhevsky et al.,, 2012) achieved strong results on the ImageNet dataset (Deng et al.,, 2009), researchers showed that AlexNet’s feature could be repurposed to novel tasks like scene recognition or medical diagnosis (Donahue et al.,, 2014; Sharif Razavian et al.,, 2014; Litjens et al.,, 2017).
+Soon after AlexNet (Krizhevsky et al., 2012) achieved strong results on the ImageNet dataset (Deng et al., 2009), researchers showed that AlexNet’s features could be repurposed to novel tasks like scene recognition or medical diagnosis (Donahue et al., 2014; Sharif Razavian et al., 2014; Litjens et al., 2017).
 This result was striking because features trained via supervision on fixed object classes nevertheless enabled transfer to novel objects or even different types of tasks.
-This pattern has continuend, and more recently researchers showed that the representations learned by these models could transfer from Imagenet 3D household navigation and dexterous tasks, sometimes achieving better performance than using “ground-truth” state (Parisi et al.,, 2022; Yuan et al.,, 2022).
-Beyond their useful representations, vision models trained with naturalistic data also develop mechanisms they were not trained for.
-For example, scene-oriented CNNs develop mechanisms for object detection (Zhou et al.,, 2014) and vision transformers develop mechanisms for segmentation (Caron et al.,, 2021)—both without an explicit training signal.
+This pattern has continued, and more recently researchers showed that the representations learned by these models could transfer from ImageNet to 3D household navigation and dexterous tasks, sometimes achieving better performance than using “ground-truth” state (Parisi et al., 2022; Yuan et al., 2022).
+Beyond their useful representations, vision models trained with naturalistic data also develop mechanisms they were not explicitly trained for.
+For example, scene-oriented CNNs develop mechanisms for object detection (Zhou et al., 2014) and vision transformers develop mechanisms for segmentation (Caron et al., 2021)—both without an explicit training signal.
 Example 2: Reinforcement learning.
-Likewise in reinforcement learning, researchers have found that agents trained on many (billions) of tasks exhibit strong “out-of-distribution” generalization on unknown tasks (Team et al.,, 2021), learning novel tasks as efficiently as humans (Team et al.,, 2023).
-Some algorithms can even generalize to collaborating with humans without any human data (Strouse et al.,, 2021).
-Recently, reinforcement learning algorithms for learning a modern variant of the successor representation (Gershman,, 2018; Carvalho et al.,, 2024) have been shown to develop skills for exploration and behavior without explicit training signals (Liu et al.,, 2024).
+Likewise in reinforcement learning, researchers have found that agents trained on many (billions) of tasks exhibit strong “out-of-distribution” generalization on unknown tasks (Open Ended Learning Team et al., 2021), learning novel tasks as efficiently as humans (Adaptive Agent Team et al., 2023).
+Some algorithms can even generalize to collaborating with humans without any human data (Strouse et al., 2021).
+Recently, reinforcement learning algorithms for learning a modern variant of the successor representation (Gershman, 2018; Carvalho et al., 2024) have been shown to develop skills for exploration and behavior without explicit training signals (Liu et al., 2024).
 Example 3: Large language models.
@@
 These models are only trained to predict the next “token” (or word) to appear in a sequence.
-However, the distribution of internet language effectively includes a broad mixture of many tasks (Radford et al.,, 2019).
-When trained on these vast naturalistic data, language models develop mechanisms for disparate tasks such as modular arithithmic, solving word problems, etc.
-(Wei et al.,, 2022)—and even for adapting to new tasks from examples (Brown et al.,, 2020).
-Moreover, the representations learned by these models can transfer to superficially dissimilar downstream tasks (Lu et al.,, 2022), including playing video games (Reid et al.,, 2022).
+However, the distribution of internet language effectively includes a broad mixture of many tasks (Radford et al., 2019).
+When trained on these vast naturalistic data, language models develop mechanisms for disparate tasks such as modular arithmetic, solving word problems, etc.
+(Wei et al., 2022)—and even for adapting to new tasks from examples (Brown et al., 2020).
+Moreover, the representations learned by these models can transfer to superficially dissimilar downstream tasks (Lu et al., 2022), including playing video games (Reid et al., 2022).
 These examples illustrate how learning from a broad naturalistic distribution can induce many abilities—and can transfer to downstream tasks.
@@
 Example 1: language models and learnability.
-Recent progress in language modeling has suggested challenges to prior assumptions about language learnability and innateness.
-We briefly sketch these issues; see Piantadosi, (2023) and Futrell and Mahowald, (2025) for thorough accounts.
-Classical approaches to language processing focused on simple models and ideas (e.g.
-Chomsky,, 2014), but in doing so imposed strong assumptions.
-Under certain strong assumptions, Gold, (1967) proved that even relatively simple languages cannot be learned from input alone.
-This proof contributed to arguments that for innate universals such as recursion, and separation of syntax from semantics (Chomsky,, 1957, 1965, e.g.)—ideas that were subsequently influential in arguments about the nature of cognition more broadly (Fodor et al.,, 1975; Fodor and Pylyshyn,, 1988; Quilty-Dunn et al.,, 2023).
+Recent language modeling works raise challenges to prior assumptions about language innateness.
+[^2] We briefly sketch these issues; see Piantadosi (2023) and Futrell and Mahowald (2025) for thorough accounts.
+Classical approaches to language focused on simple models (e.g.
+Chomsky, 2014) under strong assumptions.
+Using certain assumptions, Gold (1967) proved that even relatively simple languages cannot be learned from input alone.
+This contributed to arguments that language’s precursors must be innate (Chomsky, 1957, 1965, e.g.)—which influenced cognitive science writ large (Fodor and others, 1975; Fodor and Pylyshyn, 1988; Quilty-Dunn et al., 2023).
 However, these theories did not attempt to model language processing in all its naturalistic detail.
-By contrast, when deep-learning models learn from large amounts of naturalistic data, these models acquire aspects of both syntax (Manning et al.,, 2020) and semantics (Li et al., 2021a, )—and account for many behavioral and neuroscientific phenomena (Schrimpf et al.,, 2021).
-Although these models often learn from inhuman quantities of language (Wilcox et al.,, 2024), even models trained on a human-like quantity can learn complex syntactic features that were previously considered evidence for innate knowledge (e.g.
-Wilcox et al.,, 2023).
-Likewise, while some “impossible” languages that (adult) humans struggle to learn had been considered evidence for universal grammar, Kallini et al., (2024) show that language models can more easily learn real languages than impossible ones.
-Finally, controlled experiments—like those from Misra and Mahowald, (2024), reviewed above—begin to illustrate which features of experience contribute to syntactic generalization: particularly the importance of naturalistic variation.
+By contrast, deep-learning models that learn from plentiful naturalistic data acquire aspects of syntax (Manning et al., 2020) and semantics (Li et al., 2021a)—and capture behavioral and neuroscientific phenomena (Schrimpf et al., 2021).
+Although these models often learn from inhuman quantities of language (Wilcox et al., 2024), even human-like quantities allow learning complex syntactic features that were previously considered evidence for innate structure (e.g.
+Wilcox et al., 2023).
+Furthermore, controlled experiments—like Misra and Mahowald (2024), reviewed above—show how naturalistic variation contributes to this syntactic generalization.
 Thus, studying learnability from naturalistic data has helped to reshape our understanding of the necessary and sufficient features for acquiring linguistic capabilities.
 Example 2: naturalistic data can lead to brain-like functional specialization.
-Given the social importance of face perception, it might seem likely that face recognition is innately encoded in the brain.
-Indeed, the discovery of the Fusiform Face Area (FFA)—a visual region specialized for face perception (Kanwisher and Yovel,, 2006)—supports this hypothesis.
-However, recent computational work shows that this kind of specialization can emerge from a domain-general computer-vision model purely through training on a naturalistic distribution of object and face recognition tasks (Dobs et al.,, 2022).
-As the authors state: “It may be that the only inductive bias humans need to develop their face system is the already well-established early preference of infants to look at faces.” Of course, this does not necessarily imply that the FFA is not innately specified.
-However, it does illustrate how training computational models over naturalistic data can help us generate hypotheses about the causes of our neural organization.
-5 Building generalizable models
-Figure 6: Overview of key strategies cognitive science can adopt to develop generalizable models. (Figures reproduced from Geirhos et al.,, 2018; Bowers et al.,, 2023)
-ML research focuses on developing generalizable models.
-While we focus on their utility for working with naturalistic tasks, we believe cognitive science can broadly benefit from improved practices for developing generalizable models.
-Here, we highlight three key practices of empirical ML research that we believe computational cognitive science can adopt.
-The first is frictionless reproducibility: developing research artifacts that can be reused and repurposed with minimal effort (§5.1).
-The second is a focus on developing generalizable models that can learn successfully across many datasets (§5.2).
-We argue that these two offer the foundation for the third: collectively collaborating on important research problems through dynamic benchmarks (§5.3).
-Before we continue, we acknowledge that modern empirical ML is a young field which, like other fields, has faced challenges in rigor and reproducibility (Melis et al.,, 2017; Lucic et al.,, 2018; Riquelme et al.,, 2018; Henderson et al.,, 2018; Recht et al.,, 2019; Agarwal et al.,, 2021).
-Despite these challenges, we believe that some of the practices of ML are valuable for improving rigor and reproducibility, as we highlight below.
-5.1 Frictionless reproducibility: a superpower
-[…]
-In recent years, data-centric sciences are experiencing a profound transition to what David Donoho calls “frictionless reproducibility” (Donoho,, 2024; cf.
-Recht,, 2024).
-Donoho identifies three key pillars to frictionless reproducibility: data sharing, code sharing, and competitive challenges.
-We focus on the first two in this section, and competitive challenges in §5.3.
-Critically, with frictionless reproducibility, the research artifacts enable future researchers to exactly re-execute the complete workflow with minimal effort.
-Thus, good ideas are spread, adopted, and improved with little friction.
-Donoho argues that ML is ostensibly the most successful at frictionless reproducibility.
-Below we review some strategies that may aid cognitive science.
+Deep neural networks trained on natural images have long predicted neural responses from V1 through IT (Olshausen and Field, 1996; Lee et al., 2007; Yamins and DiCarlo, 2016; Doshi and Konkle, 2023), establishing them as candidate models of visual processing.
+Recently, researchers have found that training with naturalistic data can yield brain-like functional specialization—something once thought to require innate hard-wiring.
+Given the social importance of faces, it might seem likely that face recognition is innately encoded in the brain.
+Indeed, the discovery of the Fusiform Face Area (FFA)—a visual region specialized for face perception (Kanwisher and Yovel, 2006)—supports this view.
+However, Dobs et al.
+(2022) showed that this specialization can emerge from a domain-general vision model trained on a naturalistic distribution of object and face recognition tasks.
+As the authors state: “It may be that the only inductive bias humans need to develop their face system is the already well-established early preference of infants to look at faces.” Of course, these results do not necessarily imply that the brain’s functional organization is not innately specified.
+However, they illustrate how training computational models on naturalistic data can generate new hypotheses about the origins of neural organization.
+Example 3: naturalistic multi-view signals can yield human-level 3D shape perception.
+While large vision models trained on classification tasks capture some features of human-like vision, they show subhuman performance in areas like 3D object perception (Bonnen et al., 2024; cf.
+Malhotra et al., 2023).
+However, recently Bonnen et al.
+(2026) suggest that this may be due to these models lacking another key feature of human visual experience: the experience of seeing a scene from multiple angles (due to movement or stereo vision).
+Indeed, the authors show that models that are trained to predict view positions from multiple views of the same scene can match human-level 3D shape perception accuracy—thus showing how naturalistic learning signals beyond classification or image-only self-supervision might contribute to human 3D perceptual abilities.
+5 Building generalizable models for naturalistic settings
+through systematic and collective effort
+How do we build generalizable models that predict and explain human behavior in naturalistic settings?
+ML has made substantial progress in building models that exhibit aspects of intelligence in settings closer and closer to the real world.
+We argue that this trajectory holds lessons for cognitive scientists who aim to model human-like intelligence in such settings.
+The first lesson is that progress requires systematic iteration from simpler settings to more naturalistic ones.
+In principle, a single researcher can systematically cover the space of experiments needed to build generalizable models.
+In practice, the design space is vast.
+This makes individual progress difficult.
+The second lesson is that collective effort enables fast progress over a vast design space.
+Initially, ML researchers worked on disparate, self-defined problems, leading to diffuse progress (Recht, 2024).
+Progress accelerated when the field adopted shared benchmarks.
+A widely adopted benchmark is a sociological object: it certifies, often through peer review, that experts judge a problem worth solving.
+With the right infrastructure, this allows large populations of researchers to make fast progress together.
+We propose that cognitive science adopt dynamic meta-benchmarks: living collections of benchmarks that researchers continually contribute to, where each benchmark tests a specific facet of a cognitive capacity (§5.1, see Figure 6).
+Consider object recognition.
+One facet might test whether classification relies on shape or texture; another whether shape processing uses local or global features; a third how classification interacts with scene context.
+Meta-benchmarks grow in two ways: researchers can increase the naturalism of existing facets (e.g., expanding shape-versus-texture stimuli from simple shapes to shapes derived from naturalistic stimuli; Geirhos et al., 2020), or add new facets exposing untested dimensions such as whether objects are classified by local- or global-shape features (Baker et al., 2020).
+In either way, benchmarks can cumulatively incorporate experimental tests of new hypotheses, without relinquishing the demand to explain prior results.
+A model that predicts human behavior across the full meta-benchmark provides evidence for a generalizable account of human capacities within that domain; failures on particular benchmarks identify where an account breaks down.
+Adopting meta-benchmarks raises an engineering problem: how can we develop models that can operate across many benchmarks, stimulus sets, and task sets?
+The key is to develop models with an eye towards generalization by ensuring both that models generalize to new data, and that their development procedure generalizes as well (§5.2).
+That is, the procedure used to develop a model for one benchmark should allow development for other benchmarks with minimal modification.
+This is essential for efficiently validating models across a full meta-benchmark.
+But how do we adopt development strategies that promote reuse across many benchmarks, stimulus sets, and task sets?
+One key strategy is frictionless reproducibility  5.3; Donoho, 2024), where researchers develop research artifacts that can be reused and repurposed with minimal effort.
+This includes standardized evaluation protocols and common interfaces that allow researchers to easily run any model against any benchmark.
+We argue that together, these pieces will enable cognitive scientists to develop generalizable models that capture the full range of natural behavior for a cognitive capacity in question.
+Figure 6: How cognitive scientists can systematically build generalizable models of cognition that explain human behavior on naturalistic data.
+Our key proposal is leveraging dynamic meta-benchmarks5.1) where (a) individual benchmarks aggregate human data from experiments studying facets of a cognitive capacity.
+The benchmarks evolve as researchers add new facets—for example, extending object classification benchmarks to test whether classification relies on shape versus texture, then making stimuli (at least partially) reflect real-world data distributions.
+(b) Models are developed with generalization as the primary goal (§5.2): the same model should predict human behavior across all benchmark facets.
+(c) To facilitate frictionless reproducibility5.3), models are evaluated through a standardized interface.
+(d) This makes it easy for researchers to comprehensively evaluate their hypothesis (i.e.
+model) against a large swath of human behavioral data capturing different facets of a cognitive capacity.
+The best hypothesis is then the one that captures human behavior over the most human data.
+(Images reproduced from Geirhos et al., 2018)
+5.1 Benchmarks: catalysts for progress and innovation
+Since at least Newell (1994), cognitive science has sought to develop unified, task-performing models that explain cognition as a whole (Kriegeskorte and Douglas, 2018; Almaatouq et al., 2024).
+While cognitive scientists are aware that relying on a single paradigm may reduce the generalizability of conclusions (Holzmeister et al., 2024; Yarkoni, 2022), the field currently lacks infrastructure for this kind of multi-paradigm evaluation.
+The result is fragmentation and Newell’s critique—that piece-wise investigation of brain components would not yield a cohesive understanding (Newell, 1973)—remains relevant forty years later (Schrimpf et al., 2020; DiCarlo et al., 2023); findings from different experimental paradigms are difficult to build upon and resist integration, even within a domain (Almaatouq et al., 2024).
+We discuss how two concepts from ML—benchmarks and leaderboards—can enable collective, systematic effort within cognitive science.
 Recommendations for cognitive science
 R1.
-Use standardized evaluation protocols where models are compared via a common interface.
-Machine learning’s foundation of standardized evaluation protocols traces back to 1959, when Highleyman and Kamentsky, (1959) created the first alphanumeric pattern-recognition dataset and the first ever “train-test” split.
-This strategy continued with the famous MNIST dataset (LeCun et al.,, 1998).
-MNIST simplified data-sharing and evaluation by pre-processing data, fixing train and test distributions, and sharing the data online.
-This standardization enabled fair comparisons between diverse approaches, from boosted stumps (Kégl and Busa-Fekete,, 2009) to support vector machines (Cristianini and Scholkopf,, 2002) or nearest neighbors (Khan,, 2017).
-Standardized evaluation protocols serve a critical function: mitigating personal bias in model assessment.
-The evolution to Caltech 101 (Fei-Fei et al.,, 2004) and ImageNet (Deng et al.,, 2009) demonstrates this principle, as increasingly naturalistic datasets enabled fair comparisons between diverse model families.
-This enabled AlexNet (Krizhevsky et al.,, 2012) to validate deep learning in an era when many researchers doubted its potential.
-Thus, we recommend that cognitive scientists establish clear hold-out data for both model development and data analysis.
-This is especially critical for confirming exploratory results (Poldrack et al.,, 2017).
-Combined with standardized evaluation protocols, this strategy enables fair comparison between diverse theoretical frameworks ranging from program synthesis to deep neural networks to reinforcement learning.
+Embrace the leaderboard principle.
+A common concern with benchmarks is that researchers will “overfit” to test sets, which theoretically should underestimate the true test error (Hastie et al., 2009).
+Yet despite widespread iteration on benchmarks in machine learning the field has produced models with strong performance on held-out data.
+This apparent paradox is partially explained by a phenomenon known as the “leaderboard principle” (Hardt and Recht, 2022).
+Researchers typically only publish and build upon models that demonstrate substantial improvements over prior results, rather than minor variations.
+This selective pressure towards meaningful advances effectively constrains the degree of adaptation to test sets, as researchers focus on substantial improvements rather than exhaustively exploring the test set’s peculiarities.
+Nevertheless, a risk of overfitting to data remains.
+We argue that cognitive science can avoid this through the creation of dynamic phenomena-oriented benchmarks, where researchers can contribute new challenges that identify a missing component of a cognitive domain of interest.
+We detail this below.
 R2.
-Make data accessible with a single line of code.
-Fully processed research data should be programmatically accessible with a single line of code, requiring no permissions.
-While uploading data to platforms like the Open Science Framework (OSF) is valuable, we additionally recommend platforms like Hugging Face that offer version control, free hosting, authentication-free access, access with one line of code, and uploading data with a few lines of code.
-This removes friction from reusing and updating datasets.
-Model Citations Age GitHub Stars Single file examples
-ResNet
-He et al., (2016)
-247,921 8 years 2,300 model
-Contrastive Learning
-He et al., (2020)
-13,480 4 years 4,800 model, training
-Object Discovery
-Locatello et al., (2020)
-790 4 years 394 model
-RL Library
-Huang et al., (2022)
-252 2 years 5,600 model 1, model 2
-RLHF Method
-Rafailov et al., (2024)
-1,730 1 year 2,200 train
-Multi-agent RL
-Rutherford et al., (2023)
-24 0 years 433 model
-Table 1: Comparison of ML libraries that adopted the single file philosophy.
-R3.
-Adopt the “single file” philosophy.
-Ideally, key workflows (e.g.
-model or evaluation definitions) should be defined in a single file.
-This has emerged as a powerful force for reproducibility in ML.
-While this approach contradicts traditional software engineering principles of modularity and abstraction, it offers distinct advantages for research code: (1) complete understanding of an important component (e.g.
-a modell) requires reading only one file, (2) minimal abstraction layers and dependencies, and (3) easy adaptation of prior methods or their components, by simply copying this single file into another codebase.
-Removing friction from future reuse also benefits us—as Table 1 shows, single-file methods get many citations.
+Develop dynamic benchmarks to iteratively capture important phenomena.
+ML benchmarks are not collected to test specific hypotheses (Hardt and Recht, 2022) and so may seem unscientific.
+However, cognitive science can embrace scientific benchmarks that allow testing specific hypotheses, while situating those tests within the context of the broader phenomena of the cognitive domain.
+We use the story of “Brain-Score” as an illustrative example of initial steps in this direction.
+Brain-Score (Schrimpf et al., 2020) began as a benchmark collecting behavioral and neural data from humans and monkeys on object recognition tasks.
+One common critique of Brain-Score is that it does not test hypotheses about human object recognition.
+For example, Bowers et al.
+(2023) noted that matching human predictions is insufficient to identify a good model of human object recognition.
+The authors reference 9 experiments from the psychophysics literature which presented stimuli generated to test specific hypotheses—for example, testing use of local shape features (Baker et al., 2020), shape vs.
+texture  (Geirhos et al., 2018), and contours  (Puebla and Bowers, 2022)—and showed that deep neural networks failed to capture human behavior across these.
+In response, DiCarlo et al.
+(2023) added many of these stimulus sets to Brain-Score, so that future models can be evaluated against all of them.
+This story illustrates what an exciting strand of computational cognitive science research could look like: one where researchers develop models and experiments that easily integrate with prior findings on a phenomenon thanks to phenomena-oriented benchmarks, generalizable models (§5.2), and frictionless reproducibility (§5.3).
+We are not the first to identify that cognitive science is fragmented and to suggest benchmarks that aggregate results from the literature as a solution.
+Indeed, researchers spanning memory, timing, confidence and more have already begun to do so (MacWhinney and Snow, 1985; Oberauer et al., 2018; Aydoğan et al., 2024; Rahnev et al., 2020; Frank et al., 2017).
+Here, we both encourage more subdisciplines within cognitive science to follow suit, and emphasize the importance of frictionless reproducibility and building generalizable models.
+While prior work has begun developing benchmarks, documentation for downloading and using the data is sometimes lacking; download access is not always authentication-free; and maybe most importantly, if one wants to add new data, there is not always an easy way to re-run all previous models and ensure that findings are commensurate.
+More critically, while prior benchmarks have focused on aggregating experimental results from “simplified” experiments, we advocate for dynamic benchmarks whose development is oriented around explaining data from increasingly naturalistic experiments (i.e.
+experiments with either parametrically generated naturalistic stimuli or, eventually, stimuli taken from real-world conditions).
 5.2 Developing models with an eye towards generalizability
-Task-performing computational models have been a goal of cognitive science since at least  Newell, (1973).
-This sentiment persists today (Kriegeskorte and Douglas,, 2018; Almaatouq et al.,, 2024) and is exemplified in conferences that explicitly seek to integrate cognitive science and neuroscience with AI  (Naselaris et al.,, 2018).
-Cognitive scientists are well aware that relying on a single paradigm to study an effect may reduce the generalizability of conclusions (Holzmeister et al.,, 2024; Yarkoni,, 2022).
 Ideally, the models we develop generalize beyond the setting they were designed for.
@@
 Second, that the behavioral (and potentially neural) predictions made by our models capture data on these new tasks.
-Clearly, the first is a requirement for the second.
+Clearly, the first requirement is a prerequisite for the second.
 Here, we detail suggestions from ML for how cognitive science can develop more generalizable models.
 Recommendations for cognitive science
+R3.
+Iterate between (1) model-design (2) task design where the other is fixed.
+One crucial lesson from ML research is the importance of separating model development from task design.
+Patterson et al.
+(2023) strongly caution against simultaneously developing both the problem setting (datasets or tasks) and the solution method, as researchers may inadvertently design evaluations that favor their proposed solution.
+Historical examples illustrate this risk.
+When developing Q-learning algorithms, researchers modified the classic pendulum control problem by adding episode cutoffs and random state resets (Machado et al., 2018).
+These modifications made Q-learning more tractable by implicitly improving exploration, but thereby obscured fundamental limitations in Q-learning’s exploration compared to policy gradient methods.
+Maintaining a strict separation between task design and model development enforces a clear delineation between hypothesis formation and testing, which is important for reducing experimenter bias.
+This is especially important because researchers tend to find “positive” results (Scheel et al., 2021; Sarafoglou et al., 2022), a bias exacerbated by an increasingly competitive academic landscape (Lee, 2014; Reithmeier et al., 2019; Woolston, 2021).
+Much of ML’s progress can be attributed to researchers evaluating their models on independently developed datasets and tasks—a practice that both simplifies the research process and promotes scientific rigor.
+Thus, we recommend that cognitive science also develop models on tasks/stimuli developed by others.
+This may seem to contradict the scientific method, where researchers craft experiments to test their specific hypotheses.
+The data from prior work was for a different hypothesis!
+However, this is precisely where dynamic, phenomena-oriented benchmarks can offer a solution.
+First, where researchers hypothesize a failure of existing theories or models, they can develop a new experimental paradigm that illustrates it, and then commit this paradigm to a dynamic benchmark.
+Once the paradigm has been publicly committed, they can move on to proposing and testing new models to explain it—which will simultaneously be tested for their continued ability to explain the other phenomena included in the benchmark.
+In this way, researchers can pursue hypothesis-driven experiment- and model-development, while maintaining scientific rigor by avoiding the temptation to overfit tasks and models to each other.
+We detail the practical approaches that enable this beneficial cycle further in the next section.
 R4.
 Develop models with simplified tasks, but ensure they work with increasingly naturalistic tasks.
-The frame problem (Shanahan,, 2004) and the “simulation is doomed to succeed” critique (Grim et al.,, 2013) highlight a fundamental challenge in model development: the evaluation tasks we design often fail to capture the full complexity of the phenomenon we care about.
-mchine learning addresses this by developing models that must succeed in both simplified and increasingly complex settings.
-The history of RL shows one such progression, focused on increasing naturalism: initial environments were grid-worlds (Sutton and Barto,, 2018), later 2D video games (Bellemare et al.,, 2013), then virtual 3D homes (Szot et al.,, 2021; Li et al., 2021b, ), and now open-world environments (Fan et al.,, 2022; Matthews et al.,, 2024; Hughes et al.,, 2024).
-Despite this increased emphasis on naturalistic settings, there is still emphasis on the value of simpler settings for studying specific aspects of models (Osband et al.,, 2019; Obando Ceron et al.,, 2024).
-Indeed, the strategy is generally to first develop your model in simplified settings, but then to verify that your model continues to “work” with more naturalistic settings.
-Collectively, this has allowed RL researchers to advance both theoretical (Dadashi et al.,, 2019; Grimm et al.,, 2020; Lyle et al.,, 2022) and empirical (Obando Ceron et al.,, 2024; Farebrother et al.,, 2024) understanding of how their models work, but also develop performant models for many“complex” settings, including beating world experts in Go (Silver et al.,, 2017) and controlling real robots (Levine et al.,, 2016; Cheng et al.,, 2024).
+The frame problem (Shanahan, 2004) and the “simulation is doomed to succeed” critique (Grim et al., 2013) flag a recurring failure: the evaluation tasks we design rarely capture the full complexity of the phenomenon we care about.
+ML’s response is to demand that models succeed in both simplified and increasingly complex settings.
+RL exemplifies the trajectory: initial environments were grid-worlds (Sutton and Barto, 2018), later 2D video games (Bellemare et al., 2013), then virtual 3D homes (Szot et al., 2021; Li et al., 2021b), and now open-world environments (Matthews et al., 2024; Hughes et al., 2024).
+Even with this push toward naturalism, researchers still rely on simpler settings to study specific aspects of model behavior and generalization (Osband et al., 2019; Obando Ceron et al., 2024).
+Indeed, the standard strategy is to first develop a model in simplified settings, then verify that it continues to work in more naturalistic ones.
+This has yielded both theoretical (Grimm et al., 2020; Lyle et al., 2022) and empirical (Obando Ceron et al., 2024; Farebrother et al., 2024) understanding of how RL models work, alongside performant models for complex settings—beating world experts in Go (Silver et al., 2017) and controlling real robots (Levine et al., 2016; Cheng et al., 2024).
 R5.
-Develop models that work on many task.
-The ultimate goal of ML is generalization (Highleyman and Kamentsky,, 1959; Hardt and Recht,, 2022; Recht,, 2024).
+Develop models that work on many tasks.
+The ultimate goal of ML is generalization (Highleyman and Kamentsky, 1959; Hardt and Recht, 2022; Recht, 2024).
 Initially, this meant models that generalize to new data.
 Now, it means both models and learning procedures that generalize to new settings.
-Thus, the practice of evaluating models across diverse datasets to showcase their generality has become foundational in ML.
-Early examples include “dropout” (Srivastava et al.,, 2014), which demonstrated improvements across 6 image and speech datasets, and MoCo (He et al.,, 2020), which validated its approach on 9 vision datasets.
-Another notable example is the seminal “Deep Q-Network” paper (Mnih et al.,, 2015), which showed that a single neural network architecture with fixed hyperparameters [^2] E.g., the number of layers or the learning rate.
+Thus, evaluating models across diverse datasets has become foundational in ML.
+Early examples include “dropout” (Srivastava et al., 2014), which demonstrated improvements across 6 image and speech datasets, and MoCo (He et al., 2020), which validated its approach on 9 vision datasets.
+Another notable example is the seminal “Deep Q-Network” paper (Mnih et al., 2015), which showed that a single neural network architecture with fixed hyperparameters [^3] E.g., the number of layers or the learning rate.
 could master 50 different Atari games.
-The field has since evolved to organize environments by the cognitive capabilities they test (e.g., exploration, generalization, manipulation) (Patterson et al.,, 2023), with researchers developing models on some task sets and validating on entirely different ones to ensure robust generalization (Machado et al.,, 2018).
+The field has since evolved to organize environments by the cognitive capabilities they test (e.g., exploration, generalization, manipulation) (Patterson et al., 2023), with researchers developing models on some task sets and validating on entirely different ones to ensure robust generalization (Machado et al., 2018).
 We recommend cognitive scientists also test the models that support their theories more broadly.
-Importantly, we should evaluate models on different tasks from those we use to develop them (Machado et al.,, 2018).
+Importantly, we should evaluate models on different tasks from those we use to develop them (Machado et al., 2018).
 This further helps prevent our model (and thereby theory) from “overfitting” to a particular paradigm.
+5.3 Frictionless reproducibility: a superpower
+Cognitive science and psychology face substantial friction reusing prior data and code (Nosek et al., 2015; Hardwicke et al., 2018): data are rarely standardized, analyses are often underreported, and shared code frequently contains errors (Poldrack et al., 2017).
+In a large-scale replication analysis, $38\%$ of code was not usable and only $31\%$ of workflows reproduced (Hardwicke et al., 2018).
+Meanwhile, data-centric sciences are undergoing a profound transition—what Donoho (2024) calls frictionless reproducibility—in which good data- and code-sharing practices combine with competitive challenges to let researchers re-execute prior workflows with minimal effort, accelerating progress dramatically.
+ML, Donoho argues, is ostensibly the most successful at frictionless reproducibility.
+Below we review some strategies that cognitive science can adopt.
+§5.1 addresses competitive challenges via dynamic meta-benchmarks; here we focus on data and code sharing.
+Recommendations for cognitive science
 R6.
-Iterate between (1) model-design (2) task design where the other is fixed.
-One crucial lesson from ML research is the importance of separating model development from task design.
-Patterson et al., (2023) strongly caution against simultaneously developing both the problem setting (datasets or tasks) and the solution method, as researchers may inadvertently design evaluations that favor their proposed solution.
-Historical examples illustrate this risk.
-When developing Q-learning algorithms, researchers modified the classic pendulum control problem by adding episode cutoffs and random state resets (Machado et al.,, 2018).
-These modifications made Q-learning more tractable by implicitly improving exploration, but thereby obscured fundamental limitations in Q-learning’s exploration compared to policy gradient methods.
-Maintaining a strict separation between task design and model development enforces a clear delineation between hypothesis formation and testing, which is important for reducing experimenter bias.
-This is especially important because researchers tend to find “positive” results (Scheel et al.,, 2021; Sarafoglou et al.,, 2022), a bias exacerbated by an increasingly competitive academic landscape (Lee,, 2014; Reithmeier et al.,, 2019; Woolston,, 2021).
-Much of ML’s progress can be attributed to researchers evaluating their models on independently developed datasets and tasks—a practice that both simplifies the research process and promotes scientific rigor.
-Thus, we recommend that cognitive science also develop models on tasks/stimuli developed by others.
-This may seem to contradict science, where researchers craft experiments to test their specific hypotheses.
-The data from prior work was for a different hypothesis!
-We recommend that the field develop phenomena-oriented benchmarks that aggregate stimuli and tasks designed to test specific theoretical predictions for broad phenomena.
-This approach would allow researchers to leverage existing experimental paradigms while maintaining scientific rigor through independent validation.
-We detail this further in the next section.
-5.3 Benchmarks: catalysts for progress and innovation
-Another challenge for cognitive science is fragmentation.
-Newell’s seminal critique (Newell,, 1973)—that piece-wise investigation of brain components would not yield a cohesive understanding—remains relevant forty years later (Schrimpf et al.,, 2020; DiCarlo et al.,, 2023).
-Current experimental paradigms often produce disconnected, incommensurate findings that resist integration across studies, even within a domain (Almaatouq et al.,, 2024).
-This fragmentation manifests not only in methodology but also in how researchers distribute their attention across different problems, thereby diluting collective progress.
-The history of machine learning offers an instructive parallel.
-Initially, researchers worked on disparate, self-defined problems, leading to diffuse progress (Recht,, 2024).
-The introduction of shared benchmarks drove a crucial shift: researchers began focusing on standardized problems that others had identified as important.
-Putting aside whether this is good or bad, it has been effective.
-Indeed, this is the third pillar of frictionless reproducibility which Donoho, (2024) argues has been a major catalyst for progress in ML: competitive challenges.
-We argue that cognitive science can similarly stand to benefit from benchmarks.
-Recommendations for cognitive science
+Use standardized evaluation protocols where models are compared via a common interface.
+Machine learning’s foundation of standardized evaluation protocols traces back to 1959, when Highleyman and Kamentsky (1959) created the first alphanumeric pattern-recognition dataset and the first ever “train-test” split.
+This strategy continued with the famous MNIST dataset (LeCun et al., 1998).
+MNIST simplified data-sharing and evaluation by pre-processing data, fixing train and test distributions, and sharing the data online.
+This standardization enabled fair comparisons between diverse approaches, from boosted stumps (Kégl and Busa-Fekete, 2009) to support vector machines (Cristianini and Scholkopf, 2002) or nearest neighbors (Khan, 2017).
+Standardized evaluation protocols serve a critical function: mitigating personal bias in model assessment.
+The evolution to Caltech 101 (Fei-Fei et al., 2004) and ImageNet (Deng et al., 2009) demonstrates this principle, as increasingly naturalistic datasets enabled fair comparisons between diverse model families.
+This enabled AlexNet (Krizhevsky et al., 2012) to validate deep learning in an era when many researchers doubted its potential.
+Thus, we recommend that cognitive scientists establish clear hold-out data for both model development and data analysis.
+This is especially critical for confirming exploratory results (Poldrack et al., 2017).
+Combined with standardized evaluation protocols, this strategy enables fair comparison between diverse theoretical frameworks ranging from program synthesis to deep neural networks to reinforcement learning.
 R7.
-Embrace the leaderboard principle.
-A common concern with benchmarks is that researchers will “overfit” to test sets, which theoretically should underestimate the true test error (Hastie et al.,, 2009).
-Yet despite widespread iteration on benchmarks in machine learning the field has produced models with strong performance on held-out data.
-This apparent paradox is partially explained by a phenomenon known as the “leaderboard principle” (Hardt and Recht,, 2022).
-Researchers typically only publish and build upon models that demonstrate substantial improvements over prior results, rather than minor variations.
-This selective pressure towards meaningful advances effectively constrains the degree of adaptation to test sets, as researchers focus on substantial improvements rather than exhaustively exploring the test set’s peculiarities.
-Nevertheless, a risk of overfitting to data remains.
-We argue that cognitive science can avoid this through the creation of dynamic phenomena-oriented benchmarks, where researchers can contribute new challenges that identify a missing component of a phenomenon of interest.
-We detail this below.
+Make data accessible with a single line of code.
+Fully processed research data should be programmatically accessible with a single line of code, requiring no permissions.
+While uploading data to platforms like the Open Science Framework (OSF) is valuable, we additionally recommend platforms like Hugging Face that offer version control, free hosting, authentication-free access, access with one line of code, and uploading data with a few lines of code.
+This removes friction from reusing and updating datasets.
 R8.
-Develop dynamic benchmarks to iteratively capture important phenomena.
-ML benchmarks are not collected to test specific hypotheses (Hardt and Recht,, 2022) and so may seem unscientific.
-However, cognitive science can embrace “scientific” benchmarks oriented toward testing specific hypotheses.
-We use the story of “Brain-Score” as an illustrative example of initial steps in this direction.
-Brain-Score (Schrimpf et al.,, 2020) began as a benchmark collecting behavioral and neural data from humans and monkeys on object recognition tasks.
-One common critique of Brain-Score is that the behavioral analysis essentially measures whether models and humans found the same stimuli easy or challenging to identify.
-However, it did not test hypotheses about how humans identify objects, nor show that models identified objects in this same way.
-For example, Bowers et al., (2023) noted that matching human predictions is insufficient to identify a good model of human object recognition.
-The authors reference 9 experiments from the psychophysics literature which presented stimuli generated to test specific hypotheses, and showed that deep neural networks failed to capture human behavior across these.
-Examples include stimuli testing the hypothesis that object recognition was based on local shape features (Baker et al.,, 2020), stimuli testing the hypothesis that human object recognition is defined by shape more than by texture (Geirhos et al.,, 2018), and stimuli testing the hypothesis that pairs of objects were judged to be similar based on shape contours (Puebla and Bowers,, 2022).
-In response, DiCarlo et al., (2023) added many of these stimulus sets to Brain-Score so that when future models are tested on whether they capture human behavior across all of these stimulus sets.
-This story illustrates what an exciting strand of computational cognitive science research could look like: one where researchers develop models and experiments that easily integrate with prior findings on a phenomenon thanks to phenomena-oriented benchmarks, frictionless reproducibility (§5.1), and generalizable models (§5.2).
-By combining these, we achieve a robust foundation for building theories that span the full space of natural behavior associated with a phenomenon.
-We are not the first to identify that cognitive science is fragmented and to suggest benchmarks that aggregate results from the literature as a solution.
-Indeed, researchers spanning memory, timing, confidence and more have already begun to do so (MacWhinney and Snow,, 1985; Oberauer et al.,, 2018; Aydoğan et al.,, 2024; Rahnev et al.,, 2020; Frank et al.,, 2017).
-Here, we both encourage more subdisciplines within cognitive science to follow suite but also emphasize the importance of frictionless reproducibility and building generalizable models.
-While prior work has begun the development of benchmarks, documentation for downloading and using the data is sometimes lacking; download access is not always authentication-free; and maybe most importantly, if one wants to add new data, there isn’t an easy way to re-run all previous models and ensure that findings are commensurate.
-Critically, while prior benchmarks have focused on aggregating experimental results from “simplified” experiments, we advocate for dynamic benchmarks whose development is oriented around explaining data from increasingly naturalistic experiments (i.e.
-experiments with either parametrically generated naturalistic stimuli or, eventually, stimuli taken from real-world conditions).
+Code should be easy to “fork” and iterate on.
+There should be minimal friction for new researchers to take your existing code and try their ideas on it.
+One particularly promising strategy is the “single file” philosophy.
+Ideally, key workflows (e.g.
+model or evaluation definitions) should be defined in a single file.
+This has emerged as a powerful force for reproducibility in ML.
+While this approach contradicts traditional software engineering principles of modularity and abstraction, it offers distinct advantages for research code: (1) complete understanding of an important component (e.g.
+a model) requires reading only one file, (2) minimal abstraction layers and dependencies, and (3) easy adaptation of prior methods or their components, by simply copying this single file into another codebase.
+Removing friction from future reuse may also benefit us, as more people can build on and cite our work.
 6 Building from naturalistic experiments to cognitive theories
-Figure 7: Overview of how we can develop theories with potentially opaque models and unnatural manipulations of natural data. (Panel C bottom figure is reproduced from Saxe et al., 2019.)
+Figure 7: Overview of how we can develop theories with potentially opaque models and unnatural manipulations of natural data: building task-performing models that can reproduce behavior across naturalistic and unnatural stimuli, while tightly coupling them to reductive explanations.
+(Panel C top figure is reproduced from Saxe et al.
+2019; panel A images reproduced from Geirhos et al.
+(2018).
 A primary goal of cognitive science is understanding.
-By instantiating our theories in a model, we are forced to make our theories more precise by concretizing the ambiguous details of the mapping from a verbal hypothesis to its implementation (Guest and Martin,, 2021).
-From this perspective, building computational models that perform the task in as naturalistic a way as possible, across as wide a variety of settings as possible, imposes much stronger constraints on our theories than an abstracted model at a higher-level (cf.
-Cao and Yamins, 2024b, ).
+By instantiating our theories in a model, we are forced to make our theories precise by concretizing the ambiguous details of the mapping from a verbal hypothesis to its implementation (Guest and Martin, 2021).
+From this perspective, building computational models that perform the task in as naturalistic a way as possible, across as wide a variety of settings as possible, imposes much stronger constraints on our theories than an abstracted model at a higher level (cf.
+Cao and Yamins, 2024b).
 These added constraints are important; otherwise, high-level theories that do not directly engage with the details of the implementation can be so unconstrained as to lack explanatory value (cf.
-Erev and Greiner,, 2015; Jones and Love,, 2011; Rahnev and Denison,, 2018; Andrews,, 2021), can be intractable for real problems (Van Rooij,, 2008), or can elide important details (see above).
+Erev and Greiner, 2015; Jones and Love, 2011; Rahnev and Denison, 2018; Andrews, 2021), can be intractable for real problems (Van Rooij, 2008), or can elide important details (see above).
 Yet moving towards naturalistic settings, and the complex models they require, alters the way that we need to approach experimentation and theory building.
-Here, we discuss the consequences for experiment design (§6.1-6.2), the role and interpretation of models (6.3), and theory building (§6.4).
+Here, we discuss the consequences for experiment design (§6.1-6.2), the role and interpretation of models (§6.3), and theory building (§6.4).
 6.1 Performing controlled experiments by parametrically manipulating naturalistic data (in unnatural ways)
@@
 One approach to deconfounding these variables is to use a narrow distribution of highly-simplified artificial stimuli (light green) that isolate the variables of interest.
-However, we advocate for (additionally) creating broader distributions of stimuli that preserve more of the breadth of the natural data distribution, while deconfounding or manipulating features parametrically in order to test hypotheses—as in the unnatural combinations of natural shapes and textures created by Geirhos et al., (2018).
-Developing and testing models on truly naturalistic data introduces new issues that are absent in simple artificial settings.
+However, we advocate for (additionally) creating broader distributions of stimuli that preserve more of the breadth of the natural data distribution, while deconfounding or manipulating features parametrically in order to test hypotheses—as in the unnatural combinations of natural shapes and textures created by Geirhos et al.
+(2018).
+Developing and testing models on truly naturalistic data introduces new issues compared to simple artificial settings.
 Naturalistic data distributions may be too “easy”—e.g.
-they may only rarely include edge-cases that test key capabilities (Zhang et al.,, 2023)—thus, testing average performance over naturalistic data may disguise important failures.
+they may only rarely include edge-cases that test key capabilities (Zhang et al., 2023)—thus, testing average performance over naturalistic data may disguise important failures.
 Relatedly, in naturalistic data, features may be confounded, which can prevent readily disentangling the solutions a model uses.
-For example, models trained on ImageNet (Deng et al.,, 2009) rely on features like texture more than humans do (Geirhos et al.,, 2018), but still perform similarly on ImageNet.
-These issues limit our ability to achieve deep understanding from experimenting with only the distribution of naturalistic data (Rust and Movshon,, 2005).
-However, we believe that these challenges are surmountable by parametrically manipulating rich, naturalistic data (Fig.
-8)—preserving or enhancing its richness, while more carefully manipulating theoretically-motivated constructs (cf.
-Jain et al.,, 2024).
-For example, Geirhos et al., (2018) engineered datasets that combine natural shapes and textures in unnatural ways.
-To do so, the authors exploited the contemporary ML technique of iterative style transfer Gatys et al., (2016)—combining the features of different images at different spatial scales—thus illustrating how progress in AI models unlocks new experimental approaches.
-These conflicting shape-texture datasets helped to inspire similar experiments on humans (Jagadeesh and Gardner, 2022b, ; Jagadeesh and Gardner, 2022a, ); surprisingly, those studies found that human visual cortex likewise appears to use textural rather than shape-focused representations, and shape-driven behavior is thus mediated by downstream readout processes.
-Thus, manipulating features to test models can lead to new insights into brain function.
-A growing range of studies similarly systematically manipulate natural stimuli, while preserving as much of the natural variation as possible—whether introducing new parameters like stimulus modality manipulation (e.g.
+For example, models trained on ImageNet (Deng et al., 2009) rely on features like texture more than humans do (Geirhos et al., 2018), but still perform similarly on ImageNet.
+These issues limit our ability to achieve deep understanding when experimenting with only the naturalistic distribution (Rust and Movshon, 2005).
+However, these challenges are surmountable by parametrically manipulating rich, naturalistic data (Fig.
+8)—preserving or enhancing its richness, while still carefully manipulating theoretically-motivated constructs (cf.
+Jain et al., 2024).
+For example, Geirhos et al.
+(2018) engineered datasets that combine natural shapes and textures in unnatural ways.
+To do so, the authors exploited the contemporary ML technique of iterative style transfer (Gatys et al., 2016)—combining features of different images at different spatial scales—thus illustrating how progress in AI models unlocks new experimental approaches.
+These conflicting shape-texture datasets helped to inspire similar experiments on humans (Jagadeesh and Gardner, 2022a, b); surprisingly, those studies found that human visual cortex likewise appears to use textural rather than shape-focused representations.
+Thus, shape-driven behavior is mediated by downstream readout processes.
+Thus, manipulating naturalistic features to test models can lead to new insights into brain function.
+Many studies similarly systematically manipulate natural stimuli, while preserving as much natural variation as possible—whether introducing new parameters like modality manipulation (e.g.
 listening vs.
-reading; Deniz et al.,, 2019), inserting challenging syntactic structures in natural stories (Futrell et al.,, 2021), or systematically selecting stimuli that drive particular responses while preserving variety (e.g.
-Tuckute et al.,, 2024; Hosseini et al.,, 2023).
-Theory-driven, controlled experiments can be performed in more complex, naturalistic settings.
-Another direction for parametrically manipulating naturalistic data comes from using video games to study sequential decision-making and higher-level cognition (Allen et al.,, 2024).
-While games are artificial, different parameters such as the size of a world, the number of agents, the length of tasks, can be parametrically manipulated to reflect values more inline with those found in the real-world.
-As such, simulated environments used in ML offer valuable settings for cognitive experimentation.
-For example, Craftax (Matthews et al.,, 2024), a 2D Minecraft environment, offers parametrically-generated open-world environments containing complex goal hierarchies, which can be used for studying topics like exploration or planning.
-Other environments offer opportunities to study long-horizon generalization (Nikulin et al.,, 2024), coordination (Rutherford et al.,, 2024), physical reasoning (Matthews et al.,, 2025), etc.
-Because these environments are parametrically generated it is easy to create highly-variable-yet-controlled manipulations, e.g.
-many environments that systematically vary planning difficulty.
-Critically, a new tool called NiceWebRL [^3] https://github.com/KempnerInstitute/nicewebrl, enables easily running web-browser human experiments using the environemnts above and the many other ML environments written using Jax (Bradbury et al.,, 2018).
-Thus, there is a growing ecosystem of tools for building complex naturalistic experimental paradigms, and evaluating humans and models with them.
+reading; Deniz et al., 2019), inserting challenging syntactic structures in natural stories (Futrell et al., 2021), or systematically selecting stimuli that drive particular responses while preserving variety (e.g.
+Tuckute et al., 2024; Hosseini et al., 2023).
+Theory-driven, controlled experimental manipulations can be performed even in more complex, naturalistic settings.
+Another direction for parametrically manipulating naturalistic data comes from using video games to study sequential decision-making and higher-level cognition (Allen et al., 2024).
+Games combine rich environments with the ability to parametrically manipulate features to reflect naturalistic values, while introducing experimental manipulation.
+For example Carvalho et al.
+(2025) tested human goal preplay using Craftax (Matthews et al., 2024)—a 2D Minecraft environment that offers parametrically-generated open-world environments containing complex goal hierarchies.
+By sampling complex maps, while constraining possible paths and goal locations, the authors created varied naturalistic experiments that could nevertheless cleanly disentangle different models of goal generalization.
+For researchers who want to build on games, there is a growing ecosystem of tools for building complex naturalistic experimental paradigms, and evaluating humans and models with them— such as NiceWebRL [^4] https://github.com/KempnerInstitute/nicewebrl, which enables web-browser human experiments using ML environments.
 Moreover, expanding the naturalism of our experiments enhances our ability to develop generalizable theories, as we argue next.
 6.2 Incorporating naturalistic variation to enhance generalizability
-By testing hypotheses in settings where other aspects of the task and data generating process are sampled broadly from the naturalistic distribution, we address some of the challenges of generalizability that arise from testing on a much narrower distribution than the written hypothesis implies (cf.
-Yarkoni,, 2022).
-In particular, testing our hypotheses across varied settings—“design[ing] with variability in mind” (Yarkoni,, 2022)—increases the likelihood of identifying robust effects that will generalize (cf.
-Baribault et al.,, 2018).
+By testing hypotheses in settings where other aspects of the task and data generating process are sampled broadly from the naturalistic distribution, we address some of the generalizability challenges that arise from testing on a much narrower distribution than the written hypothesis implies (cf.
+Yarkoni, 2022); “design[ing] with variability in mind” (Yarkoni, 2022) increases the likelihood of identifying robust effects that will generalize (cf.
+Baribault et al., 2018).
 In the language of regression, varying as many other data generating factors as possible brings us closer to estimating the “main effect” of an experimentally-manipulated construct, rather than estimating the “simple effect” of the manipulation in the single setting we have tested.
 For example, it is a long-standing principle of experimental design to test a hypothesis with multiple stimuli, and statistically quantify the stimulus effect to estimate generalizability (e.g.
-Clark,, 1973).
+Clark, 1973).
 However, many other experiment features—even exact task formulations such as multi-arm bandit tasks—are often preserved within a paper, or even across an entire literature.
-Just like testing a hypothesis with a single stimulus, this limited variation poses challenges for generalizability (Eckstein et al.,, 2022)—especially because the way that researchers operationalize a theory can substantially affect their conclusions (e.g.
-Holzmeister et al.,, 2024; Schweinsberg et al.,, 2021; Strickland and Suben,, 2012).
+Just like testing a hypothesis with a single stimulus, this limited variation poses challenges for generalizability (Eckstein et al., 2022)—especially because the way that researchers operationalize a theory can substantially affect their conclusions (e.g.
+Holzmeister et al., 2024; Schweinsberg et al., 2021; Strickland and Suben, 2012).
 While naturalistic variation cannot resolve this issue in full, we believe that incorporating a broader range of the settings, stimuli, and tasks within which we expect a theory to hold, will help us to develop more generalizable theories.
 Incorporating variation can also help us to revise the constructs underlying our theories.
-For example, Eisenberg et al., (2019) study the putative construct of self-regulation by simultaneously testing many existing self-regulation measures.
+For example, Eisenberg et al.
+(2019) study the putative construct of self-regulation by simultaneously testing many existing self-regulation measures.
 The authors find that these measures do not form a unitary construct.
@@
 By contrast, by substantially varying the tasks we use to measure a theoretical component like self-regulation, we can more fully disentangle the underlying constructs—and thereby improve the extent to which our theories generalize.
+Likewise, Peterson et al.
+(2021) collect a large dataset of risky choice judgments, and use it to explore how human choices interpolate between different prior decision-making theories depending on the problem context.
 6.3 Interpreting complex models to yield explanations
-If we perform experiments in naturalistic settings, and develop complex models that can predict human behavior across them, how does this support explaining the cognitive phenomena of interest?
+How can complex experiments in naturalistic settings, and complex models that can predict behavior across them, support explaining the cognitive phenomena of interest?
 Here, we illustrate some paths to deriving explanations from complex models.
 [^5] N.B., not all complex systems admit abstract explanations.
-Where natural intelligence incorporates chaotic (Freeman,, 1995) or critical (O’Byrne and Jerbi,, 2022) dynamics, they may alter the explanations that we can seek (cf.
-Kellert,, 1993).
+For example, where natural intelligence incorporates chaotic (Freeman, 1995) or critical (O’Byrne and Jerbi, 2022) dynamics, they may alter the explanations that we can seek (cf.
+Kellert, 1993).
 A benefit of seeking predictive models as a route to explanation is that the models may be useful even if simple explanations do not follow.
@@
 One way to explain model behavior is as a consequence of the properties of the data from which it learns.
-This idea has a rich history in cognitive science andd AI (McClelland and Jenkins,, 1991; Elman,, 1996; Rogers and McClelland,, 2004), in which researchers attempted to identify minimal properties of naturalistic data that give rise to some behavior(s).
-For example, Chan et al., (2022) explore how the bursty, long-tailed nature of natural data can give rise to in-context learning, and Prystawski et al., (2024) examine how local dependencies can yield certain kinds of sequential reasoning.
-These works illustrate how rich properties of natural data can be distilled down to minimal elements that drive a behavior—and can thus offer a candidate explanation for the origin of those capabilities.
-Moreover, explanations of behavior via properties of the data offer a route toward normative explanations of that behavior as a rational solution to constraints—precisely the perspective taken by rational analysis (Anderson,, 1991).
+This idea has a rich history in cognitive science and AI (McClelland and Jenkins, 1991; Elman, 1996; Rogers and McClelland, 2004), in which researchers attempted to identify minimal properties of naturalistic data that give rise to some behavior(s).
+More recently, these techniques have been applied to understanding AI models.
+For example, Chan et al.
+(2022) explore how the bursty, long-tailed nature of natural data can give rise to in-context learning, and Prystawski et al.
+(2024) examine how local dependencies can yield certain kinds of sequential reasoning.
+Likewise, Hermann et al.
+(2020) demonstrate how properties of data created by augmentation strategies can drive the texture bias investigated with parametric manipulations by Geirhos et al.
+(2018), as discussed above.
+These works illustrate how rich properties of natural data can be distilled down to minimal elements that drive a behavior or computational strategy—and can thus offer a candidate explanation for the origin of those capabilities or biases.
+Moreover, explanations of behavior via properties of the data point towards normative explanations of that behavior as a rational solution to constraints—precisely the perspective taken by rational analysis (Anderson, 1991; Lieder and Griffiths, 2020).
+For example, (Lampinen et al., 2024) compare language models and humans on various logical reasoning tasks—from natural language inference to syllogism validity judgments (Evans et al., 1983) and the Wason selection task (Wason, 1968)—instantiated in both synthetic and naturalistic settings.
+The authors find that models and humans show common patterns of content-entangled reasoning and graded uncertainty.
+By studying these common patterns, the authors argue that prior interpretations of particular phenomena (e.g., social explanations of the Wason selection task; Cosmides, 1989) were overfit to the experiments explained.
+Instead, the authors argue for a unifying (resource) rational account of these superficially-different biases as stemming from the underlying statistical properties of natural data for both humans and models: content effects arise from optimizing prediction over natural data where certain conclusions tend to hold much more often than others.
+This example shows how applying rational analysis across naturalistic settings can generate unifying accounts of seemingly-disparate phenomena.
 Mechanistic explanation.
-Understanding that a behavior arises from data properties does not explain how that behavior is implemented.
-Fortunately, the internal workings of task-performing computational models can generally be inspected and intervened upon.
-For example, a variety of works have analyzed the mechanisms that implement language model behaviors, either in simplified settings (e.g.
-Nanda et al.,, 2023; Zhong et al.,, 2024), or in large models trained on natural data (Geiger et al.,, 2021; Wu et al.,, 2024).
-A particularly promising approach for cognitive science is illustrated by Distributed Alignment Search (Geiger et al.,, 2024), which attempts to map postulated abstract causal models onto low-level features of a complex model that play corresponding causal roles—thus enabling linking between abstract models or hypotheses about the computational solution, and evidence about concrete implementational mechanisms.
-Morever, mechanistic interventions can even be combined with data studies, to examine the causal role of different mechanisms over the course of learning (Singh et al.,, 2024).
-Appropriate regularization of models can also produce representations that yield greater interpretability—e.g.
-Miller et al., (2024) train networks to predict animal behavior while restricting their representations, and used these to identify non-optimal features influencing the decisions—thus allowing bottom-up discovery of possible implementations.
+Understanding that a behavior arises from data properties does not explain its implementation.
+Fortunately, the internal workings of computational models can generally be inspected and intervened upon.
+For example, many works have analyzed the mechanisms that implement language model behaviors, either in simplified settings (e.g.
+Nanda et al., 2023; Zhong et al., 2024), or in large models trained on natural data (Geiger et al., 2021; Wu et al., 2024).
+A particularly promising approach is illustrated by Distributed Alignment Search (Geiger et al., 2024), which maps postulated abstract causal models onto low-level features of a complex model that play corresponding causal roles—thus enabling linking between abstract models or computational hypotheses to evidence about concrete implementational mechanisms.
+Moreover, mechanistic interventions can even be combined with data studies, to examine the causal role of different mechanisms over learning (Singh et al., 2024).
+Appropriate regularization of models can also yield more interpretable representations—e.g.
+Miller et al.
+(2024) train networks with regularized representations to predict animal behavior, and used them to identify non-optimal features influencing the decisions—thus allowing bottom-up discovery of possible implementations.
 Thus, various methods for mechanistic study of models can offer powerful routes to experimentally determining sufficient implementations of cognitive processes.
@@
 From data properties or mechanisms, we can progress to formal theories.
-Saxe et al., (2019) offers an illustrative example.
-Prior works observed empirically that pseudo-naturalistic data elicit various aspects of human-like semantic development and representation in neural networks (Rogers and McClelland,, 2004), such as progressive differentiation.
+Saxe et al.
+(2019) offers an illustrative example.
+Prior works observed empirically that pseudo-naturalistic data elicit various aspects of human-like semantic development and representation in neural networks (Rogers and McClelland, 2004), such as progressive differentiation.
 However, these works had not explained why these parallels emerge.
-Saxe et al., (2019) bridged this gap by formally deriving how the data properties combine with gradient-based learning to yield these cognitive phenomena.
+Saxe et al.
+(2019) bridged this gap by formally deriving how the data properties combine with gradient-based learning to yield these cognitive phenomena.
 This illustrates how preliminary explanations in terms of data can be developed into more rigorous analytic theories.
 Thus, building naturalistic models can be a stepping stone toward more rigorous, normative theories of intelligence as a rational solution to a particular problem setting—much like prior work in cognitive science (e.g.
-Anderson,, 1991; Oaksford and Chater,, 2009).
-6.4 Theories that combine task-performing models with reductive explanation.
+Anderson, 1991; Oaksford and Chater, 2009).
+Bayesian or other reductive cognitive models.
+Another analytic approach is to interpret complex models using the very tools of cognitive modeling, for example by building computational-level Bayesian models (Griffiths et al., 2024; Ku et al., 2025).
+Indeed, various recent works have taken the approach of reducing from more complex task-performing models to simpler models based on prior cognitive work.
+For example, Peterson et al.
+(2021, discussed above) initially fit an unconstrained context-dependent neural model of the behavior, but reduced it to a similarly-predictive model that effectively does context-dependent interpolation between prior decision-making theories—thus providing a unifying but relatively interpretable account of a broader range of phenomena (cf.
+Fintz et al., 2022).
+As an example of the Bayesian approach, Zhu and Griffiths (2024) show how language models produce some human-like (Zhu et al., 2020) patterns and biases in their probabilistic judgments.
+The authors link these patterns of behavior to a computational-level Bayesian model—and formally relate that model to the probabilistic inference performed by the language models.
+These examples, like those in the prior section, show how it is possible to build explicit links from complex task-performing models to simpler, more classic computational cognitive models.
+6.4 Theories that couple task-performing models with reductive explanations.
 Above, we have argued for the value of embracing the richness of naturalistic tasks, even if it leads us to build more complex task-performing models that are hard to interpret, but that perhaps make more accurate predictions of behavior.
@@
 These two components serve different purposes.
-The first helps to demonstrate that the models are real candidate models of the phenomena in question, rather than models that simply could not scale to the real problem or that are overfit to a particular instantiation of the task.
-This component is necessary because otherwise the problem space is under-constrained.
+The first component—a task-performing model—helps to demonstrate that the posited theory offers a real candidate account of the full range of natural phenomena in question, rather than relying on approaches that simply could not scale to the real problem or that are overfit to a particular instantiation of the task.
+Conversely, this component increases the probability that our model accounts for all the possible constraints the real stimuli may impose on human learning and task behavior.
+As outlined above, this component is necessary because otherwise the problem space is under-constrained.
 Moreover, it provides a test-bed for experiments that would be impossible in reality, which is useful both scientifically and practically.
-The second component allows linking these predictions to the explanatory understanding that cognitive scientists typically seek, and that may help us to generalize our insights beyond current paradigms.
-This perspective aligns with past arguments on how deep learning can contribute to understanding in cognitive neuroscience (Saxe et al.,, 2021; Kanwisher et al.,, 2023; Cao and Yamins, 2024b, ; Doerig et al.,, 2023).
-However, we emphasize that combining these two components makes explicit links between the actual problem the system solves and the theorized constructs underlying that solution.
+From a scientific perspective, building task-performing models enables testing the model on other phenomena within the same task space, which can enable more unifying accounts of how phenomena relate.
+It also enables exploring hypotheses that would be impossible to test in reality, such as the effect of removing a particular syntactic construction from a learner’s linguistic experience.
+Practically, being able to simulate responses to real stimuli is also useful for the many applications of cognitive science to practical problems in domains such as education.
+The second component—reductive explanations—allows linking these complex predictive models to the explanatory understanding that cognitive scientists typically seek, and that may help us to generalize our insights beyond current paradigms.
+As outlined above, this component can build on the explanatory approaches of prior Bayesian, connectionist, analytic, or normative theories in cognitive science.
+This part of our perspective aligns with past arguments on how deep learning can contribute to understanding in cognitive neuroscience (Saxe et al., 2021; Kanwisher et al., 2023; Cao and Yamins, 2024b; Doerig et al., 2023).
+However, we emphasize that synthesizing these two components makes explicit, testable links between the actual problem the system solves and the theorized constructs underlying that solution.
 It also ensures that our theories make testable predictions beyond the narrow settings of a particular task paradigm.
 7 Discussion
-In this paper, we have tried to outline a direction of research that we call “naturalistic computational cognitive science,” which aims to build generalizable models of cognition that scale to naturalistic tasks, while still offering routes to explanatory and theoretical understanding.
+In this paper, we have tried to outline a direction of research that we call “naturalistic computational cognitive science,” that aims to build generalizable models and theories of cognition that scale to naturalistic tasks, while still affording explanatory understanding.
 This perspective synthesizes a growing literature on the importance of naturalistic experiments and generalizable models, and grows out of a long-standing focus in cognitive science on explaining a broad range of phenomena.
@@
 The quest for building general models of natural intelligence.
-A long history of cognitive science research has sought frameworks with the generality to explain a wide scope of cognitive phenomena.
+Many cognitive scientists have sought frameworks with the generality to explain a wide scope of cognitive phenomena.
 Researchers in cognitive architecture (e.g.
-Ritter et al.,, 2019; Laird,, 2019), connectionist (e.g.
-McClelland et al.,, 1986, 2003), and Bayesian (e.g.
-Tenenbaum and Griffiths,, 2001; Griffiths et al.,, 2010; Tenenbaum et al.,, 2011), paradigms have tried to identify underlying principles or mechanisms that explain many phenomena.
+Ritter et al., 2019; Laird, 2019), connectionist (e.g.
+McClelland et al., 1986, 2003), and Bayesian (e.g.
+Tenenbaum and Griffiths, 2001; Griffiths et al., 2010; Tenenbaum et al., 2011), paradigms have tried to identify underlying principles or mechanisms that explain many phenomena.
 However, while these modeling frameworks themselves are general, typical instantiations of the frameworks have built specialized models focused on individual tasks.
-Indeed, one core critiques of many modeling paradigms is their failure to accurately capture patterns of human generalization (e.g.
-Fodor et al.,, 1975; McClelland and Patterson,, 2002; Lake et al.,, 2017).
+Indeed, one core critique of many modeling paradigms is their failure to accurately capture patterns of human generalization (e.g.
+Fodor and others, 1975; McClelland and Patterson, 2002; Lake et al., 2017).
 In pursuit of generality, some recent studies have built upon foundation models from AI that can perform many tasks.
 Some works have applied cognitive paradigms to study the behaviors of these models (e.g.
-Binz and Schulz,, 2023; Lampinen et al.,, 2024; Buschoff et al.,, 2025).
+Binz and Schulz, 2023; Lampinen et al., 2024; Buschoff et al., 2025).
 Others have finetuned these models using cognitive data, to create generalizable models that can predict behavior on new experimental paradigms in areas like vision (e.g.
-Muttenthaler et al., 2024b, ; Muttenthaler et al., 2024a, ; Fu et al.,, 2024), or the broader space of cognitive tasks that can be presented in language (Binz and Schulz,, 2024; Binz et al.,, 2024).
+Muttenthaler et al., 2024, 2025; Fu et al., 2024), or the broader space of cognitive tasks that can be presented in language (Binz and Schulz, 2024; Binz et al., 2024).
 We see the arguments that we have laid out in this paper as being broadly compatible with many of these approaches and frameworks.
 However, we offer a stronger emphasis on the role of theory (cf.
-Frank and Goodman,, 2025), and the virtuous cycle between expanding the scope and naturalism of our experimental designs, and expanding the generalizability of our models and theories.
+Frank and Goodman, 2025), and the virtuous cycle between expanding the scope and naturalism of our experimental designs, and expanding the generalizability of our models and theories.
 Towards experiments that involve naturalistic behavior Our perspective aligns with various works that have likewise highlighted the value of considering complex naturalistic behaviors and the environments in which they occur.
 The strands of cognitive research focused on embodiment and enactivism have advocated that the environment, and organism-environment interactions, form an important component of cognition (e.g.
-Clark,, 1998; Varela et al.,, 2017).
+Clark, 1998; Varela et al., 2017).
 More generally, naturalistic experiments are frequently advocated in neuroscience (e.g.
-Mobbs et al.,, 2021; Cisek and Green,, 2024), where even prominent visual processing signals can fundamentally differ in active, naturalistic paradigms (e.g.
-Amme et al.,, 2024).
-Similarly, Gao and Ganguli, (2015); Gao et al., (2017) highlight that more complex behavioral paradigms are needed to unlock the full space of neural dynamics, and Nastase et al., (2020) argue that isolating only a few variables in our experiments impairs our ability to understand the full scope of natural behavior.
-Likewise, Wise et al., (2023) advocate for the need to explore more naturalistic environments and behaviors in reinforcement learning, to grapple with the complexity of behavior in environments closer to the real world.
-We concur with these works, and emphasize that advances in technology both enable richer experimental paradigms, and the models that can perform them.
+Mobbs et al., 2021; Cisek and Green, 2024), where even prominent visual processing signals can fundamentally differ in active, naturalistic paradigms (e.g.
+Amme et al., 2024).
+Similarly, Gao and Ganguli (2015); Gao et al.
+(2017) highlight that more complex behavioral paradigms are needed to unlock the full space of neural dynamics, and Nastase et al.
+(2020) argue that isolating only a few variables in our experiments impairs our ability to understand the full scope of natural behavior.
+Likewise, Wise et al.
+(2023) advocate for the need to explore more naturalistic environments and behaviors in reinforcement learning, to grapple with the complexity of behavior in environments closer to the real world.
+We concur with these works, and emphasize that advances in technology both enable richer experimental paradigms and provide the models that can perform them.
 We differ from many of these works in more strongly emphasizing the role of models that generalize across task paradigms.
-Linking between learning models and the brain A large recent literature has explored surprising alignments between the representations learned by task-optimized deep learning models and neural representations (Khaligh-Razavi and Kriegeskorte,, 2014; Yamins et al.,, 2014; Yamins and DiCarlo,, 2016; Schrimpf et al.,, 2021; Sucholutsky et al.,, 2023).
-These findings have stimulated substantial interest and debate, ranging from critical argument that failures to capture particular phenomena doom these models (e.g.
-Bowers et al.,, 2023), to suggestions that deep learning models upend the theoretical assumptions of the field (Perconti and Plebe,, 2020; Hasson et al.,, 2020).
-Other authors have written frameworks that attempt to integrate deep learning models within the existing scientific paradigms (Doerig et al.,, 2023).
-In this vein, several papers have proposed frameworks for understanding the role of deep learning in cognitive neuroscience (e.g Richards et al.,, 2019; Storrs and Kriegeskorte,, 2019; Cichy and Kaiser,, 2019)—and how these models can contribute to deriving explanatory and theoretical understanding (Cao and Yamins, 2024a, ; Cao and Yamins, 2024b, ; Saxe et al.,, 2021).
-In this context, Feather et al., (2025) argue for the importance of benchmarks that test alignment of both behavior and representations—and highlight how dataset limitations like lack of stimulus variability can limit our ability to discriminate between models.
+Linking between learning models and the brain A large recent literature has explored surprising alignments between the representations learned by task-optimized deep learning models and neural representations (Khaligh-Razavi and Kriegeskorte, 2014; Yamins et al., 2014; Yamins and DiCarlo, 2016; Schrimpf et al., 2021; Sucholutsky et al., 2025).
+These findings have stimulated substantial interest and debate, ranging from critical arguments that failures to capture particular phenomena doom these models (e.g.
+Bowers et al., 2023), to suggestions that deep learning models upend the theoretical assumptions of the field (Perconti and Plebe, 2020; Hasson et al., 2020).
+Other authors have written frameworks that attempt to integrate deep learning models within the existing scientific paradigms (Doerig et al., 2023).
+In this vein, several papers have proposed frameworks for understanding the role of deep learning in cognitive neuroscience (e.g., Richards et al., 2019; Storrs and Kriegeskorte, 2019; Cichy and Kaiser, 2019)—and how these models can contribute to deriving explanatory and theoretical understanding (Cao and Yamins, 2024a, b; Saxe et al., 2021).
+In this context, Feather et al.
+(2025) argue for the importance of benchmarks that test alignment of both behavior and representations—and highlight how dataset limitations like lack of stimulus variability can limit our ability to discriminate between models.
 As above, our perspective aligns with many of these works, but we place greater emphasis on the complementary role of increasing naturalism of experimental paradigms, and learning from the engineering paradigms of AI, rather than simply adopting its models as artifacts.
+Underdetermination in science Our arguments touch on broader issues in the philosophy of science.
+In particular, our scientific theories are always underdetermined by the finite experiments that we have performed (Stanford, 2023)—our conclusions about what an experiment (dis)proves hinge upon a collection of background assumptions, e.g.
+about the methods involved, the systems being experimented upon, etc.
+We see this paper as highlighting several common background assumptions underlying computational cognitive science that we believe are suspect and can be detrimental to building effective theories—in particular, the assumptions that simplified experimental settings can allow building theories of a system’s behavior that will generalize consistently to more naturalistic ones, and that modeling experiments over simplified stimuli can accurately characterize the computational challenges of learning and generalizing over naturalistic stimuli.
+In the earlier sections of the paper, we provided illustrative examples of how these assumptions can break down; in the later sections we advocated for approaches that can avoid making these assumptions.
+While our proposed methods (like any empirical approach) still rely on other background assumptions—e.g., about the fidelity of the links we can make between naturalistic models and the system being modeled—we hope that by avoiding background assumptions that have been empirically found to be problematic, we can make some progress towards more robust and general cognitive theories.
 7.1 Looking forward: challenges & opportunities for naturalistic computational cognitive science
@@
 While researchers can draw on domain expertise to identify potentially-relevant parameters, there may be unexpected influences from other factors that are not typically manipulated–e.g.
-the differences between moral thoughts and moral actions when shifting from vignettes to virtual reality (Francis et al.,, 2017).
+the differences between moral thoughts and moral actions when shifting from vignettes to virtual reality (Francis et al., 2017).
 Thus, one important direction for naturalistic research may be to identify more general overarching naturalistic factors—such as perception, embodiment, or action—that tend to influence phenomena across many domains.
@@
 While this has led to fruitful understanding, it leaves open the question of what exactly is learned during each stage.
-More complete “learning”-based accounts would specify precisely what an “evolutionary” data distribution would be, and the model behaviors (evolved inductive biases) that are expected from it; a succession of “developmental” data distributions and the model behaviors that would result from them; and finally predict both the behaviors that people begin tasks with and how they will with task exposure.
+More complete “learning”-based accounts would specify precisely what an “evolutionary” data distribution would be, and the model behaviors (evolved inductive biases) that are expected from it; a succession of “developmental” data distributions and the model behaviors that would result from them; and finally predict both the behaviors that people begin tasks with and how they change with task exposure.
 Fortunately, work in this direction has already begun.
-For example, researchers have begun to explicitly map different timescales of learning to different levels of optimization in a meta-learning process (Wang,, 2021)—and indeed, some works have highlighted how meta-learning can give rise to human-like inductive biases, e.g.
-for linguistic (McCoy et al.,, 2020; McCoy and Griffiths,, 2025) or compositional (Lake and Baroni,, 2023) generalizations.
+For example, researchers have begun to explicitly map different timescales of learning to different levels of optimization in a meta-learning process (Wang, 2021)—and indeed, some works have highlighted how meta-learning can give rise to human-like inductive biases, e.g.
+for linguistic (McCoy et al., 2020; McCoy and Griffiths, 2025) or compositional (Lake and Baroni, 2023) generalizations.
 Thus, a key opportunity for naturalistic computational cognitive science is to develop learning-based account of cognition that begin to specify how learning from different data-distributions within different segments of learning can produce the changing cognitive phenomena we observe throughout ontogeny.
@@
 While we acknowledge this challenge, we see it as an opportunity for larger-scale collaborative efforts—which are increasingly adopted in computer science research on training large-scale models (e.g.
-Le Scao et al.,, 2023).
+Le Scao et al., 2023).
 By pooling expertise and resources, collaborations make progress at larger scales than individual research groups could.
 Likewise, as discussed above, collaborations on shared benchmarks (e.g.
-Schrimpf et al.,, 2020) can allow broader and deeper assessments than any individual researcher could.
+Schrimpf et al., 2020) can allow broader and deeper assessments than any individual researcher could.
 Furthermore, the growing overlap between cognitive modeling and artificial intelligence brings opportunities for mutually-beneficial collaborations between the two fields.
@@
 We first thank Sam Gershman for his extensive guidance and feedback on this article and on how to integrate cognitive science and machine learning methodologies.
-We also thank Maria Eckstein, Tyler Bonnen, Jay McClelland, Mike Mozer, Rachel Calcott, Nick Blauch, Aran Nayebi, Ruben Cohen, Greta Tuckute, Fenil Doshi and Shuze Liu for thoughtful comments and discussions.
-We are also grateful for this work being supported by a gift from the Chan Zuckerberg Initiative Foundation to establish the Kempner Institute for the Study of Natural and Artificial Intelligence.
+We also thank George Alvarez, Maria Eckstein, Tyler Bonnen, Jay McClelland, Mike Mozer, Rachel Calcott, Nick Blauch, Aran Nayebi, Ruben Cohen, Greta Tuckute, Fenil Doshi, and Shuze Liu for thoughtful comments and discussions.
 Finally, we thank Felix Hill for many inspirational conversations on generalization, cognition, and more.
+Funding statement
+This work was supported by a gift from the Chan Zuckerberg Initiative Foundation to establish the Kempner Institute for the Study of Natural and Artificial Intelligence.
+Conflicts of interest statement
+The second author is employed by Anthropic.